### v1

# Feature-aware Hypergraph Generation via Next-Scale Prediction

Abstract
Hypergraphs generalize traditional graphs by allowing hyperedges to connect multiple nodes, making them well-suited for modeling complex structures with higher-order relationships, such as 3D meshes, molecular systems, and electronic circuits. While topology is central to hypergraph structure, many real-world applications also require node and hyperedge features. Existing hypergraph generation methods focus solely on topology, often overlooking feature modeling. In this work, we introduce FAHNES (feature-aware hypergraph generation via next-scale prediction), a hierarchical approach that jointly generates hypergraph topology and features. FAHNES builds a multi-scale representation through node coarsening, then learns to reconstruct finer levels via localized expansion and refinement, guided by a new node budget mechanism that controls cluster splitting. We evaluate FAHNES on synthetic hypergraphs, 3D meshes, and molecular datasets. FAHNES achieves competitive results in reconstructing topology and features, establishing a foundation for future research in featured hypergraph generative modeling.
1 Introduction
Figure 1: Examples of generated featured hypergraphs by a sequential disjoint generation baseline and our model (FAHNES).
Hypergraphs extend traditional graphs by allowing edges—called hyperedges—to connect more than two nodes, enabling a natural representation of complex, multi-way interactions. This expressiveness makes hypergraphs well-suited for modeling systems in molecular chemistry, electronic circuit design, and 3D geometry. As a result, they have been applied in diverse areas such as drug discovery [1], contagion modeling [2], recommendation systems [3], molecular biology [4], and urban planning [5].
Despite their wide applicability, generative models for hypergraphs remain relatively underexplored. Existing approaches primarily focus on generating topology alone [6], neglecting the generation of node and hyperedge features. As shown in Figure 1, the sequential generation of topology followed by features is ineffective. Joint generation is particularly challenging for hypergraphs due to their variable-size hyperedges and higher structural complexity. Moreover, existing methods for featured graph generation use flat, non-hierarchical architectures, which scale poorly [7, 8, 9].
In this work, we introduce a feature-aware hypergraph generation via next-scale prediction (FAHNES) model, a novel hierarchical approach for jointly generating hypergraph topology and features. Inspired by multi-scale generative strategies in image synthesis [10], FAHNES leverages a coarsening-expansion strategy [6, 11]: during training, it merges nodes based on connectivity to build a hierarchical representation, then learns to progressively recover finer-scale hypergraphs. A key contribution is the use of a budget mechanism that tracks the number of permissible expansions per cluster. This enables global structural constraints to influence local generation, helping to mitigate the limited receptive fields of deep models. Additionally, we adapt flow-matching optimal transport (OT) coupling techniques [12, 13] to learn straighter and more consistent flows during generation. Our main contributions are:
  • We introduce the first hierarchical model for joint generation of hypergraph topology and features targeting specific data distributions, such as 3D meshes and molecules.
  • We propose a novel budget-based mechanism that improves global consistency during generation for hierarchical methods.
  • We generalize minibatch OT-coupling to our hypergraph generative setting.
  • We validate our approach on both synthetic and real-world datasets, including molecular structures and 3D meshes, showing promising results for hierarchical generative methods.
2 Related Work
Graph and hypergraph generation using deep learning. Graph generation has seen significant advances in recent years. Early approaches, such as GraphVAE [14], employed autoencoders to embed graphs into latent spaces for sampling. Subsequent models leveraged recurrent neural networks to sequentially generate adjacency matrices, improving structural fidelity [15]. More recently, diffusion-based methods have enabled permutation-invariant graph generation [16, 7], with extensions incorporating structural priors such as node degrees [17]. Many of these previous methods jointly model node features and topology, but are limited to small graphs due to scalability challenges. For example, diffusion-based models [7, 8, 9] operate over entire graphs by gradually corrupting structures and features, then training models to denoise them. However, their scalability is constrained by the combinatorial nature of edge prediction in large graphs.
To mitigate this, hierarchical methods have been proposed. For example, Bergmeister et al. [11] introduced a scalable graph generation framework based on a coarsen-then-expand approach, merging nodes to form coarse representations and progressively reconstructing finer details. This framework was extended to hypergraphs by Gailhard et al. [6], which allows edges to connect more than two nodes. However, both methods focus exclusively on topology generation, neglecting node and hyperedge features essential for many applications.
Applications of hypergraph generation. Generative models that capture both higher-order structure and node/hyperedge features are critical in numerous domains. In molecular design, regular graph generative models struggle to accurately represent rings and scaffolds [7], which naturally correspond to hyperedges involving multi-atom interactions rather than pairwise bonds. Similarly, 3D shape modeling involves surfaces such as triangles, quads, or general polygons that extend beyond simple pairwise connectivity. Conventional approaches often rely on fixed topology, quantization, or autoregressive sequence modeling with transformers [18, 19], limiting their flexibility and scalability. Hypergraphs provide a more expressive framework by treating faces as hyperedges, enabling the joint generation of topology and features.
In contrast to previous approaches that either target hypergraph topology alone or feature-aware graph generation, we introduce the first method that jointly models both hypergraph structure and features within a unified and scalable framework. FAHNES extends hierarchical generative methods with feature-aware expansions, a novel node budget mechanism, and flow-matching training adapted to our strategy. This combination enables effective generative modeling for featured hypergraphs.
3 Feature-aware Hypergraph Generation via Next-scale Prediction
((a)) Hypergraph
((b)) Clique expansion
((c)) Star expansion
Figure 2: Comparison of a hypergraph and its clique and star expansions.
[…]
Basic definitions. We define a graph $G=(\mathcal{V},\mathcal{E})$ as a pair consisting of a set of vertices $\mathcal{V}$ and a set of edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Graphs may also carry node and edge features, represented by matrices $\mathbf{F}_{\mathcal{V}}\in\mathbb{R}^{|\mathcal{V}|\times m}$ and $\mathbf{F}_{\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|\times l}$, where $m$ and $l$ denote the dimensionality of the features. Each edge $e\in\mathcal{E}$ corresponds to a pair $(u,v)$, indicating a connection between nodes $u$ and $v$. A bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ is a special case where the vertex set is split into two disjoint subsets $\mathcal{V}_{L}$ and $\mathcal{V}_{R}$, and edges exist only between the two parts, i.e., $\mathcal{E}\subseteq\mathcal{V}_{L}\times\mathcal{V}_{R}$. The full set of nodes is $\mathcal{V}=\mathcal{V}_{L}\cup\mathcal{V}_{R}$. In this work, we consider node features for bipartite graphs, denoted by $\mathbf{F}_{L}$ for the left-side nodes and $\mathbf{F}_{R}$ for the right-side nodes.
A hypergraph $H$ extends the concept of a graph and is specified by a pair $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the vertex set and $\mathcal{E}$ comprises hyperedges, with each $e\in\mathcal{E}$ being a subset of $\mathcal{V}$. The key distinction of hypergraphs lies in their ability to connect arbitrary numbers of vertices through a single hyperedge. Similar to graphs, hypergraphs can possess node and edge features, which are denoted $\mathbf{F}_{\mathcal{V}}$ and $\mathbf{F}_{\mathcal{E}}$. We consider two fundamental graph-based representations of hypergraphs: clique and star expansions. The clique expansion transforms a hypergraph $H$ into a graph $C=(\mathcal{V}_{c},\mathcal{E}_{c})$, where $\mathcal{E}_{c}=\{(u,v)\mid\exists~{}e\in\mathcal{E}:u,v\in e\}$. The star expansion converts a hypergraph $H$ into a bipartite structure $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E}_{b})$, where $\mathcal{V}_{L}=\mathcal{V}$, $\mathcal{V}_{R}=\mathcal{E}$, and $\mathcal{E}_{b}=\{(v,e)\mid v\in\mathcal{V}_{L},e\in\mathcal{V}_{R},v\in e\text{ in }H\}$. Left side nodes $\mathcal{V}_{L}$ represent the nodes of the hypergraph, while right side nodes $\mathcal{V}_{R}$ represent hyperedges. We provide a visual example of the clique and star expansions in Figure 2.
Our work aims to develop a model capable of sampling from the underlying distribution of a given featured hypergraph dataset $(H_{1},\dots,H_{N})$, i.e., learning the joint probability of features and topology of this distribution of hypergraphs. For detailed mathematical proofs of all propositions and lemmas presented in this work, please consult the supplementary material.
3.1 Overview
Previous work [6] simplifies hypergraph generation by reducing it to standard graph generation through two projections: i) the clique expansion where every hyperedge becomes a clique that links all its incident nodes, and ii) the star expansion, which produces a bipartite graph by introducing one node per hyperedge and connecting it to the nodes it contains. While conceptually elegant, this approach faces significant challenges: i) it struggles to produce the correct number of nodes, especially in large hypergraphs, and ii) it does not support the generation of node and hyperedge features, limiting its applicability to real-world, featured domains.
To overcome the limitations of prior approaches, we propose two key contributions. First, we introduce a node budget mechanism, in which generation begins with a single super-node carrying the full node budget. At each generation step, this budget is recursively split among child nodes. This approach enables a more localized and structured control over the final node count, compared to previous methods that simply concatenate the desired size to node embeddings. Second, we propose hierarchical feature generation, where each expansion step predicts the mean feature of its future child nodes, conditioned on the parent’s feature. This hierarchical feature generation scheme is thus a generalization of [20] to hypergraphs, where the prediction at current scale conditions the predictions at next scale. To train FAHNES, we adopt the flow-matching framework [21], where the generation process is formulated as learning to reverse a known noising procedure. Specifically, the model is trained to recover the true values for node and hyperedge expansions, edge deletions, budget allocations, and features, all of which are subjected to noise. Our workflow is illustrated in Figure 3.
Figure 3: Our framework adopts a coarsening-expansion strategy. i) During training, input hypergraphs are progressively coarsened by merging nodes and hyperedges, yielding a multiscale representation. Node features are averaged during merging, and budgets are summed. ii) The model is trained to predict which nodes were merged at each scale. iii) In the expansion phase, merged nodes are expanded (shown in dark), inheriting their parent’s features, budget, and connectivity. The model is trained to (a) identify which edges should be removed (dotted lines), (b) predict how the parent’s budget should be split across the children (instead of predicting absolute values), and (c) refine the features of newly expanded nodes (refinement).
3.2 Budgeted Coarsening
During coarsening, we keep track of two quantities for every cluster we create. First, budgets count how many nodes the cluster already contains. Every node and hyperedge begins with a budget of $1$; when several of them merge, their budgets are summed and assigned to the new super-node or super-hyperedge. Secondly, features summarize the attributes of the merged vertices. We use a weighted average: the feature of a cluster is the budget-weighted mean of the features of its members.
Proposition 1.
Setting cluster features as the average of the nodes they contain minimizes the mean squared error (MSE) between the features of the fully expanded hypergraph and those of the original hypergraph.
In what follows we denote left-side (node) budgets and features by $\mathbf{b}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{F}_{L}$, and right-side (hyperedge) budgets and features by $\mathbf{b}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ and $\mathbf{F}_{R}$.
Definition 2 (Bipartite graph coarsening).
Let $H$ be an arbitrary hypergraph, $C=(\mathcal{V}_{c},\mathcal{E}_{c})$ its unfeatured weighted clique expansion, and $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{b}_{L},\mathbf{b}_{R},\mathbf{F}_{L},\mathbf{F}_{R})$ its featured bipartite representation. Given a partitioning $\mathcal{P}_{L}=\{\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(n)}\}$ of the node set $\mathcal{V}_{L}$ such that each part $\mathcal{V}^{(p)}$ induces a connected subgraph in $C$, we construct an intermediate coarsening $\tilde{B}(B,\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\mathcal{V}_{R},\bar{\mathcal{E}},\bar{\mathbf{b}},\bar{\mathbf{F}}_{L},\mathbf{F}_{R})$ by merging each part $\mathcal{V}^{(p)}$ into a single node $v^{(p)}\in\bar{\mathcal{V}}_{L}$, and by defining:
$\left(\bar{\mathbf{b}}_{L}\right)_{v^{(p)}}=\sum_{v\in\mathcal{V}^{(p)}}\left(\mathbf{b}_{L}\right)_{v},\qquad\left(\bar{\mathbf{F}}_{L}\right)_{v^{(p)}}=\frac{1}{\left(\bar{\mathbf{b}}_{L}\right)_{v^{(p)}}}\sum_{v\in\mathcal{V}^{(p)}}\left(\mathbf{b}_{L}\right)_{v}\left(\mathbf{F}_{L}\right)_{v},$ (1)
for every supernode $\mathcal{V}^{(p)}$. An edge $e_{\{p,q\}}\in\bar{\mathcal{E}}$ is added between $v^{(p)}\in\bar{\mathcal{V}}_{L}$ and $v^{(q)}\in\mathcal{V}_{R}$ if there exists an edge $e_{\{i,q\}}\in\mathcal{E}$ in the original bipartite representation between some $v^{(i)}\in\mathcal{V}^{(p)}$ and $v^{(q)}$.
To complete the coarsening process, we define an equivalence relation $v_{1}\sim v_{2}\iff\mathcal{N}(v_{1})=\mathcal{N}(v_{2})$ on $\mathcal{V}_{R}$, where $\mathcal{N}(v)$ denotes the set of neighbors of $v$. This induces a partitioning $\mathcal{P}_{R}=\{\mathcal{V}_{R}^{(1)},\dots,\mathcal{V}_{R}^{(m)}\}$, allowing us to construct the fully coarsened bipartite representation $\bar{B}(\tilde{B},\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\bar{\mathcal{V}}_{R},\bar{\mathcal{E}},\bar{\mathbf{b}},\bar{\mathbf{F}}_{L},\bar{\mathbf{F}}_{R})$ by merging each part $\mathcal{V}_{R}^{(p)}$ into a single node $v_{R}^{(p)}\in\bar{\mathcal{V}}_{R}$, similarly to the construction of $\bar{\mathcal{V}}_{L}$, and by defining:
$\left(\bar{\mathbf{b}}_{R}\right)_{v^{(p)}}=\sum_{v\in\mathcal{V}_{R}^{(p)}}\left(\mathbf{b}_{R}\right)_{v},\qquad\left(\bar{\mathbf{F}}_{R}\right)_{v^{(p)}}=\frac{1}{\left(\bar{\mathbf{b}}_{R}\right)_{v^{(p)}}}\sum_{v\in\mathcal{V}_{R}^{(p)}}\left(\mathbf{b}_{R}\right)_{v}\left(\mathbf{F}_{R}\right)_{v},$ (2)
for every superhyperedge $\mathcal{V}_{R}^{(p)}$.
Remark 3.
Informally, we first pick the node clusters on the clique expansion, merge those nodes on the left side of the bipartite graph, and compute their budgets and features as above. We then merge right-side nodes that appear multiple times, i.e., they connect the same left-side nodes. In practice, for the final reduction step, when only one node and one hyperedge remain, we also replace their features by matrices full of zeros. This design choice enables a fully agnostic initial step during generation.
3.3 Budgeted Expansion and Refinement
During expansion, since our framework only conditions on the total number of nodes in the final hypergraph, we maintain a single node budget vector $\mathbf{b}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ for the left-side (node) partition and discard the right-side (hyperedge) budget. Each cluster is split into multiple child nodes, which initially inherit both the parent’s budget and feature vector. A subsequent refinement step then determines: i) how the parent’s budget should be divided among its children, and ii) how to update the children’s feature vectors.
More specifically, the expansion step duplicates vertices and hyperedges, with child nodes inheriting all of their parent’s connections. The refinement step then i) removes edges that should not persist at the finer resolution, ii) redistributes the integer budget among the child nodes according to a probability-like split vector, and iii) replaces the coarse features with new refined predictions.
Formally, the expansion and refinement steps can be described as follows.
Definition 4 (Bipartite graph expansion).
Given a bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{b},\mathbf{F}_{L},\mathbf{F}_{R})$ and two cluster size vectors $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$, $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$, denoting the expansion size of each node, let $\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b}^{\text{expanded}},\mathbf{F}_{L}^{\text{expanded}},\mathbf{F}_{R}^{\text{expanded}})$ denote the expansion of $B$, whose node sets, budgets, and features are given by:
$\begin{split}&\tilde{\mathcal{V}}_{L}=\mathcal{V}_{L}^{(1)}\cup\cdots\cup\mathcal{V}_{L}^{(|\mathcal{V}_{L}|)},\text{~{}where~{}~{}}\mathcal{V}_{L}^{(p)}=\{v_{L}^{(p,i)}\mid 1\leq i\leq\mathbf{v}_{L}[p]\}\text{~{}~{}for~{}~{}}1\leq p\leq|\mathcal{V}_{L}|,\\ &\tilde{\mathcal{V}}_{R}=\mathcal{V}_{R}^{(1)}\cup\cdots\cup\mathcal{V}_{R}^{(|\mathcal{V}_{R}|)},\text{~{}where~{}~{}}\mathcal{V}_{R}^{(p)}=\{v_{R}^{(p,i)}\mid 1\leq i\leq\mathbf{v}_{R}[p]\}\text{~{}~{}for~{}~{}}1\leq p\leq|\mathcal{V}_{R}|,\\ &\left(\mathbf{b}^{\text{expanded}}\right)_{v_{L}^{(p,i)}}=\mathbf{b}_{\mathcal{V}_{L}^{(p)}}\text{~{}~{}for~{}~{}}1\leq i\leq\mathbf{v}_{L}[p],1\leq p\leq|\mathcal{V}_{L}|,\\ &\left(\mathbf{F}_{L}^{\text{expanded}}\right)_{v_{L}^{(p,i)}}=\left(\mathbf{F}_{L}\right)_{\mathcal{V}_{L}^{(p)}}\text{~{}~{}for~{}~{}}1\leq i\leq\mathbf{v}_{L}[p],1\leq p\leq|\mathcal{V}_{L}|,\\ &\left(\mathbf{F}_{R}^{\text{expanded}}\right)_{v_{R}^{(p,i)}}=\left(\mathbf{F}_{R}\right)_{\mathcal{V}_{R}^{(p)}}\text{~{}~{}for~{}~{}}1\leq i\leq\mathbf{v}_{R}[p],1\leq p\leq|\mathcal{V}_{R}|.\end{split}$ (3)
The edge set $\tilde{\mathcal{E}}$ includes all the cluster interconnecting edges: $\{e_{\{p,i;q,j\}}\mid e_{\{p,q\}}\in\mathcal{E},v_{L}^{(p,i)}\in\mathcal{V}_{L}^{(p)},v_{R}^{(q,j)}\in\mathcal{V}_{R}^{(q)}\}$.
Remark 5.
Expansion thus acts as a clone-and-rewire operation: vertices and hyperedges are duplicated, and each child inherits every incident edge of its parent.
Definition 6 (Bipartite graph refinement).
Given a bipartite graph $\tilde{B}=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b},\mathbf{F}_{L},\mathbf{F}_{R})$, an edge selection vector $\mathbf{e}\in\{0,1\}^{|\mathcal{E}|}$, a budget split vector $\mathbf{f}\in[0,1]^{|\tilde{\mathcal{V}}_{L}|}$, satisfying $\sum_{v\in\mathcal{V}_{L}^{(p)}}\mathbf{f}_{v}=1$ for all cluster $\mathcal{V}_{L}^{(p)}$ in $\mathcal{V}_{L}$, and two feature refinement vectors $\mathbf{F}_{L}^{\text{refine}}$ and $\mathbf{F}_{R}^{\text{refine}}$ with the same dimensions as $\mathbf{F}_{L}$ and $\mathbf{F}_{R}$, let $B(\tilde{B},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\mathcal{E},\lceil\mathbf{b}\odot\mathbf{f}\rfloor,\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$ denote the refinement of $\tilde{B}$, where $\mathcal{E}\subseteq\tilde{\mathcal{E}}$ such that the $i$-th edge $e_{(i)}\in\mathcal{E}$ if and only if $\mathbf{e}[i]=1$.
Remark 7.
Edges are selectively removed based on the binary indicator vector $\mathbf{e}$, and features are updated with new predictions. Node budgets are divided among child nodes according to the split proportions specified by the vector $\mathbf{f}$. Since budgets must be integers, the resulting values are rounded. In the case of a tie (e.g., when an odd number must be split evenly), the child with the lowest index receives the larger share, and the remaining budget is distributed among the others accordingly.
The generation of a hypergraph with $N$ nodes proceeds in three main stages:
  1. 1.
    Initialization. Start from a minimal bipartite graph: $B^{(L)}=(\{1\},\{2\},\{(1,2)\})$, consisting of a single node on each side connected by one edge. The left-side node is assigned the full node budget, i.e., $\mathbf{b}=(N)$. If node and hyperedge features need to be generated, $\mathbf{F}_{L}$ and $\mathbf{F}_{R}$ are initialized as zero matrices.
  2. 2.
[…]
  3. 3.
    Hypergraph reconstruction. Once the final bipartite graph is generated, construct the hypergraph by collapsing each right-side node into a hyperedge connecting its adjacent left-side nodes.
3.4 Probabilistic Modeling
We now present a formalization of our learning framework, generalizing [11, 6]. Let $\{H^{(1)},\ldots,H^{(N)}\}$ denote a set of i.i.d. hypergraph instances. Our aim is to approximate the unknown generative process by learning a distribution $p(H)$. We model the marginal likelihood of each hypergraph $H$ as a sum over the likelihoods of its bipartite representation’s expansion sequences:
$p(H)=p(B)=\sum_{\varpi\in\Pi(B)}p(\varpi),$ (4)
where $\Pi(B)$ denotes the set of valid expansion sequences from a minimal bipartite graph to the full bipartite representation $B$ corresponding to $H$. Each intermediate $B^{(l-1)}$ is generated by refining its predecessor, in accordance with Definitions 4 and 6.
Assuming a Markovian generative structure, the likelihood of a specific expansion sequence $\varpi$ is factorized as:
$p(\varpi)=\underbrace{p(B^{(L)})}_{1}\prod_{l=L}^{1}p(B^{(l-1)}|B^{(l)})=\prod_{l=L}^{1}p(\mathbf{e}^{(l-1)}|\tilde{B}^{(l-1)})p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)}).$ (5)
To simplify the modeling process and avoid learning two separate distributions $p(\mathbf{e}^{(l)}|\tilde{B}^{(l)})$ and $p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)})$, we rearrange terms as follows:
$p(\varpi)=p(\mathbf{e}^{0}|\tilde{B}^{0})p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)})\left[\prod_{l=L-1}^{1}p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)}|\tilde{B}^{(l)})\right],$ (6)
where $p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)},\mathbf{b}^{(L)},\mathbf{F}_{L}^{(L)},\mathbf{F}_{R}^{(L)})=p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)},\mathbf{b}^{(L)},\mathbf{F}_{L}^{(L)},\mathbf{F}_{R}^{(L)}|B^{(L)})$.
We assume that the variables $\mathbf{v}_{L}^{(l)}$, $\mathbf{v}_{R}^{(l)}$, $\mathbf{b}^{(l)}$, $\mathbf{F}_{L}^{(l)}$, and $\mathbf{F}_{R}^{(l)}$ are conditionally independent of $\tilde{B}^{(l)}$ when conditioned on $B^{(l)}$:
$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)},\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)}).$ (7)
This allows us to write the combined likelihood as:
$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)}|\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{e}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)}).$ (8)
3.5 Flow-matching
We use a flow-matching framework [21] with an endpoint parameterization similar to [22], or alternatively a denoising diffusion model [23] with linear interpolation between the prior and target distribution. The objective is to align two probability distributions $p_{0}$ and $p_{1}$ by evolving the points from the initial distribution $p_{0}$ to the target distribution $p_{1}$. This is achieved by defining a flow through a time-dependent vector field $\mathbf{f}(\mathbf{x},t)$, governed by the following ordinary differential equation (ODE) $\frac{\partial\mathbf{x}}{\partial t}=\mathbf{f}(\mathbf{x},t),\mathbf{x}_{0}\sim p_{0}$, where $\mathbf{f}(\mathbf{x},t)$ is learned such that, after integrating the ODE, samples from $p_{1}$ can be generated.
With endpoint parametrization, the model learns to predict the endpoint $\mathbf{x}_{1}$ of the trajectory. The flow $\mathbf{f}(\mathbf{x},t)$ can then be recovered from the predicted endpoint $\hat{\mathbf{x}}_{1}(\mathbf{x},1)$ using the relation:
$\mathbf{f}(\mathbf{x}_{t},t)=\frac{\hat{\mathbf{x}}_{1}(\mathbf{x},1)-\mathbf{x}_{t}}{1-t}.$ (9)
To train the model, the loss function minimizes the squared difference between the predicted endpoint $\hat{\mathbf{x}}_{1}(x_{t},t)$ and the actual endpoint $\mathbf{x}_{1}$:
$\mathcal{L}=\mathbb{E}_{t\sim\mathcal{U}(0,1),\>\mathbf{x}_{0}\sim p_{0},\>\mathbf{x}_{1}\sim p_{1}}\left[\left(\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{1}\right)^{2}\right],\text{~{}~{}where~{}~{}}\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}.$ (10)
For the expansion process and edge selection, we set $p_{0}$ as the Gaussian distribution. For budget splits, as their support is constrained to the simplex, we use flow matching similarly to [22], where $p_{0}$ is a Dirichlet distribution across all clusters. The model predictions are projected on the simplex using a Von Neumann projection [24]. Features are generated using the process introduced in [20], i.e., $p_{0}$ is the Gaussian distribution, and the feature of the parent cluster is given as conditioning during generation using a FiLM layer [25]. More details can be found in the supplementary material.
3.6 Minibatch OT-coupling
Minibatch OT-coupling [12, 13] has shown promising results in image generation by encouraging smoother and more linear generative trajectories, which can improve sampling efficiency and model convergence. Instead of sampling independently from the prior and target distributions, OT-coupling aligns samples within each minibatch via an optimal transport plan. Specifically, samples are first drawn from the prior (typically Gaussian), then their corresponding target samples are drawn, and finally, the prior samples are optimally matched (reindexed) to the targets to minimize transport cost. We extend this idea to our setting by applying OT-coupling to the children of an expanded node, treating these children as a minibatch. For each expansion, we solve a local OT problem to align the initial noised samples of children with their corresponding targets.
A crucial constraint is to avoid swapping samples between non-equivalent nodes—i.e., only nodes that are structurally and feature-wise indistinguishable up to noise can be swapped. Swapping non-equivalent nodes would alter the learned data distribution and degrade model fidelity. To keep computational costs manageable, we restrict OT-coupling to groups of children within a single expanded cluster, satisfying the equivalence condition. Since our expansion step produces two children per cluster, OT-coupling simplifies to evaluating just two permutations per cluster, making it extremely lightweight and easily parallelizable using standard tensor operations.
We summarize our theoretical guarantee as follows:
Proposition 8 (Informal).
Minibatch OT-coupling in our framework preserves the target distribution while improving the straightness and consistency of the learned generative flow.
4 Experiments and Results
In this section, we detail our experimental setup, covering datasets, evaluation metrics, results, and ablation studies. The datasets are divided into featureless and featured hypergraphs. For the featureless hypergraphs, we compare FAHNES against HyperPA [26], a Variational Autoencoder (VAE) [27], a Generative Adversarial Network (GAN) [28], and a standard 2D diffusion model [23] trained on incidence matrix images, where hyperedge membership is represented by white pixels and absence by black pixels. For the featured hypergraphs, we compare FAHNES against the graph-based models DiGress [7], DisCo [9], Cometh [29], and DeFoG [30] in a molecular dataset. For 3D meshes, we also compare FAHNES with a sequential disjoint generation baseline (see Figure 1), where we first generate the hypergraph topology and then the features. Extended numerical results and several visualizations of our generated hypergraphs are available in the supplementary material.
The graph-based approaches DiGress, DisCo, Cometh, and DeFoG are only tested on a simple graph dataset because they cannot generate hypergraphs. Although one might attempt to generate hypergraphs using these existing graph generators over the star or clique expansions, this strategy is impractical for several reasons. First, generating the clique expansion of a hypergraph and then recovering the original hyperedges is an NP-hard problem, as it requires enumerating all cliques. Second, in the bipartite representation of a hypergraph, it is non-trivial to determine which partition represents nodes and which represents hyperedges—a distinction that is critical for correct reconstruction. Empirically, we also find that graph-based generative models often fail to produce valid bipartite structures. For instance, we observed that only 30% of outputs from both models [11, 7] generated valid bipartite graphs.
Datasets. We evaluate our method on five featureless datasets: Stochastic Block Model (SBM) [31], Ego [32], Tree [33], ModelNet40 bookshelf, and ModelNet40 piano [34]. We also evaluate FAHNES on three featured hypergraphs datasets, including two sets of 3D meshes and one molecular dataset: Manifold40 bench, Manifold40 airplane [^1] The full set of numerical results for the Manifold40 datasets is provided in the supplementary material. [35], and QM9 [36, 37] with implicit hydrogens. In the case of bench and airplane, node features are 3D positions, and hyperedges do not have features. For QM9, both nodes and hyperedges have features, which are one-hot vectors coding for atom types and bond types, respectively. Although molecules can be represented as traditional graphs, modeling them with hypergraphs opens the door to representing higher-order interactions between atoms. However, in this work, we treat molecules strictly as graphs, with hyperedges corresponding to simple pairwise bonds. Future work could group functional groups’ atoms together in hyperedges. Further details of the datasets are provided in the supplementary material.
Metrics. For featureless hypergraphs, we follow the evaluation criteria in [6]. These include: i) structural comparison metrics such as Node Num (difference in node counts), Node Deg (Wasserstein distance between node degree distributions), and Edge Size (Wasserstein distance between hyperedge size distributions); ii) topological analysis with Spectral (maximum mean discrepancy between the spectral distributions). In scenarios where datasets enforce structural constraints, we report Valid—the proportion of generated samples satisfying those constraints. For all metrics except Valid, lower values indicate improved performance, while higher values are preferable for Valid.
In the case of 3D meshes, we use the nearest training sample Chamfer distance, which computes the Chamfer distance between point clouds sampled from a generated sample and all training samples and outputs the minimum distance. For QM9, we use the same metrics as in [30]: i) overall validity of molecules with Valid (fraction of molecules that can be sanitized using RDKit [38]); ii) diversity of samples with Unique (fraction of unique SMILES generated); and iii) chemical similarity with the learned distribution with FCD (Fréchet inception distance between the activations of ChemNet on the train samples and generated samples [39]).
Implementation details. We adopt the Local PPGN architecture [11, 6]. During generation, we use inpainting to enforce constraints: i) budget splits are fixed to 1 for unexpanded or terminal clusters, ii) equal splits are enforced for expanded clusters with a budget of 2, iii) clusters of size one cannot expand, and iv) features for non-expanded clusters are copied.
Table 1: Comparison between FAHNES and other baselines for the SBM, Ego, and Tree hypergraphs.
Model SBM Hypergraphs Ego Hypergraphs Tree Hypergraphs
Valid SBM $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Ego $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Tree $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
HyperPA [26] 2.5% 0.075 4.062 0.407 0.273 0% 35.83 2.590 0.423 0.237 0% 2.350 0.315 0.284 0.159
VAE [27] 0% 0.375 1.280 1.059 0.024 0% 47.58 0.803 1.458 0.133 0% 9.700 0.072 0.480 0.124
GAN [28] 0% 1.200 2.106 1.203 0.059 0% 60.35 0.917 1.665 0.230 0% 6.000 0.151 0.459 0.089
Diffusion [23] 0% 0.150 1.717 1.390 0.031 0% 4.475 3.984 2.985 0.190 0% 2.225 1.718 1.922 0.127
HYGENE [6] 65% 0.525 0.321 0.002 0.010 90% 12.55 0.063 0.220 0.004 77.5% 0.000 0.059 0.108 0.012
[…]
FAHNES 81.4% 0.010 0.603 0.005 0.005 100% 0.162 0.171 0.129 0.007 82.8% 0.000 0.043 0.046 0.002
Table 2: Evaluation on the ModelNet40 hypergraphs.
Model ModelNet40 Bookshelf ModelNet40 Piano
Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
HyperPA [26] 8.025 7.562 0.044 0.048 0.825 9.254 0.023 0.067
VAE [27] 47.45 6.190 1.520 0.190 75.35 8.060 1.686 0.396
GAN [28] 0.000 397.2 46.30 0.456 0.000 409.0 86.38 0.697
Diffusion [23] 0.000 20.36 2.346 0.079 0.050 20.90 4.192 0.113
HYGENE [6] 69.73 1.050 0.034 0.068 42.52 6.290 0.027 0.117
[…]
FAHNES 0.875 5.600 0.087 0.019 0.125 1.878 0.028 0.036
Table 3: Evaluation on QM9.
Model QM9
Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
DiGress [7] 99.0% 96.2
DisCo [9] 99.3%
Cometh [29] 99.6% 96.8 0.25
DeFoG [30] 99.3% 96.3 0.12
[…]
FAHNES 77.8% 94.3 3.86
Comparison with the baselines. Tables 1 and 3 show the comparisons in the featureless hypergraphs. We observe that FAHNES significantly improves previous methods in many metrics. This is especially evident in the SBM and Ego hypergraphs, where the percentage of valid hypergraphs increases by 16.4% and 10%, respectively. Similarly, the spectral similarity has been improved in the ModelNet40 and Tree hypergraphs.
Table 4: Ablation studies on the node budget (Budg.) and minibatch OT-coupling (Coup.) for SBM, Ego, and Tree hypergraphs.
Budg. Coup. SBM Hypergraphs Ego Hypergraphs Tree Hypergraphs
Valid SBM $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Ego $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Tree $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
77.7% 0.020 0.889 0.041 0.006 99.0% 2.383 0.239 0.139 0.006 92.2% 0.005 0.066 0.161 0.011
79.9% 0.123 0.815 0.023 0.006 98.5% 1.721 0.223 0.184 0.010 83.8% 0.000 0.045 0.083 0.006
71.6% 0.039 1.413 0.005 0.013 100% 0.459 0.129 0.279 0.010 74.5% 0.005 0.039 0.050 0.003
81.4% 0.010 0.603 0.005 0.005 100% 0.162 0.171 0.129 0.007 82.8% 0.000 0.043 0.046 0.002
Regarding the Manifold40 dataset, FAHNES obtains a Chamfer distance of 0.073 and 0.049 for bench and airplane, respectively, while the sequential baseline reaches 0.143 and 0.117, respectively. This shows the better modeling capacity of FAHNES to jointly generate topology and features in hypergraphs instead of a simple two-step approach. Similarly, Table 3 shows the results in the QM9 dataset, where we observe FAHNES presents promising results, even though our model is not explicitly tailored for the problem of molecule generation. Extended results on featured hypergraphs and visual comparisons of the different approaches are provided in the supplementary material.
Table 5: Ablation studies on the node budget (Budg.) and minibatch OT-coupling (Coup.) for ModelNet40 datasets.
Budg. Coup. ModelNet40 Bookshelf ModelNet40 Piano
Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
8.6 5.951 0.043 0.027 8.125 3.546 0.075 0.026
7.225 4.183 0.109 0.037 2.375 5.529 0.160 0.028
1.525 5.232 0.069 0.013 11.275 4.127 0.072 0.027
0.875 5.600 0.087 0.019 0.125 1.878 0.028 0.036
Ablation studies. Tables 4 and 5 present some ablation studies of FAHNES. More precisely, we analyze the importance of the budget and minibatch OT-coupling. In general terms, we observe that using budgets instead of concatenating the target size to each node embedding has an important effect on the Node Num, Node Deg, and Edge Size metrics. Similarly, the minibatch OT-coupling has a positive effect on the validity of the spectral similarity of the generated hypergraphs.
We also observe additional interesting properties tied to the budget: i) training is more stable, with smoother curves for the metrics on the validation and test sets, and ii) we observe less variance in the quality of generated samples, which translates into a higher correlation between the metrics of the validation set and those of the test set. We hypothesize that the constraints introduced by the node budgets maintain the predictions in the high probability regions.
Limitations. While our node budget mechanism helps mitigate the issue of missing nodes, it does not fully solve it. On molecular generation tasks, FAHNES is less competitive due to the maturity of one-shot models in this domain: in hierarchical settings, structural errors accumulate over sequential predictions. In addition, FAHNES assumes continuous features, making it less effective for categorical data like atom or bond types—scenarios where discrete diffusion is more appropriate. Similarly, specialized techniques such as marginal probabilities [7], designed specifically for molecule generation, cannot be readily integrated into our setting. Furthermore, our framework prioritizes scalability, which limits the use of high-capacity backbones such as graph transformers [40]. Future work could explore discrete diffusion for topology and adopt a mixed discrete-continuous formulation [22] to better handle structured, symbolic domains.
5 Conclusion
We presented FAHNES, the first hierarchical framework for jointly generating hypergraph topology and features. By combining coarse-to-fine structural modeling with feature-aware generation, our method enables scalable synthesis of complex hypergraphs, surpassing the limitations of flat or disjoint approaches. Key innovations include a recursive node budget mechanism and the use of minibatch OT-coupling, which together improve sample quality and training convergence. Experiments across synthetic, molecular, and geometric datasets demonstrate strong performance in both topology and feature generation, often outperforming existing baselines. While FAHNES is not yet competitive with domain-specific pipelines—particularly for molecule generation—it opens promising directions for structured generative modeling.
Appendix A Implementation Details
A.1 Model Architecture
Our method represents the expansion numbers for left and right nodes, along with edge presence, as attributes of the bipartite graph. To model the distribution $p(\mathbf{v}^{(l)}_{L},\mathbf{v}^{(l)}_{R},\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}}\mid\tilde{B}^{(l)})$, we adopt an endpoint-parameterized flow-matching framework [21]. Within this framework, the attributes—namely, the expansion vectors and edge indicators—are corrupted with noise, and a denoising network is trained to reconstruct the original values.
The denoising network is structured as follows:
  1. 1.
    Positional encoding: Node positions within the graph are encoded using SignNet [41]. These encodings are replicated according to the respective expansion numbers.
  2. 2.
    Attribute embedding: Five separate linear layers are used to embed the bipartite graph attributes: left node features, right node features, edge features, node-specific features, and hyperedge-specific features. FiLM conditioning [25] is applied to incorporate contextual information into node and hyperedge features. Node budgets are embedded using sinusoidal positional encodings [42].
  3. 3.
    Feature concatenation:
    • For each left and right node, embeddings are concatenated with positional encodings and the desired reduction fraction. Left nodes also receive the node budget embedding.
    • If node features are present, they are appended to the left nodes. Likewise, hyperedge features are appended to the right nodes when available.
    • For edges, embeddings include edge features, concatenated positional encodings of the incident nodes, and the reduction fraction.
  4. 4.
    Graph processing: The attribute-enriched bipartite graph is processed through a stack of sparse PPGN layers, following the architecture from [11].
  5. 5.
    Output prediction: The final graph representations are passed through three linear projection heads to generate outputs.
    • Left node head: Predicts expansion values, budget splits, and refined node features.
    • Right node head: Predicts hyperedge expansions and refined hyperedge features.
    • Edge head: Predicts edge existence.
A.2 Flow-matching Framework
We employ a flow-matching generative modeling framework [21], with endpoint parameterization following [22], equivalent to denoising diffusion models [23] using linear interpolation between the prior $p_{0}$ and target $p_{1}$ distributions. The goal is to align two distributions by transporting samples from $p_{0}$ to $p_{1}$ through a learned time-dependent vector field $\mathbf{f}(\mathbf{x},t)$, governed by the ODE:
$\frac{\partial\mathbf{x}}{\partial t}=\mathbf{f}(\mathbf{x},t),\quad\mathbf{x}_{0}\sim p_{0}.$ (11)
Here, $\mathbf{f}(\mathbf{x},t)$ is trained such that integrating this ODE produces samples from $p_{1}$.
Endpoint parameterization. Instead of directly modeling the flow field, we learn the terminal point $\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)$ of the trajectory. The flow can be recovered using:
$\mathbf{f}(\mathbf{x}_{t},t)=\frac{\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{t}}{1-t}.$ (12)
Training objective. The model is trained to minimize the expected squared error between the predicted and true endpoints:
$\mathcal{L}=\mathbb{E}_{t\sim\mathcal{U}(0,1),\>\mathbf{x}_{0}\sim p_{0},\>\mathbf{x}_{1}\sim p_{1}}\left[\left\|\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{1}\right\|^{2}\right],$ (13)
where $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$ is a linear interpolation between samples from the prior and target distributions.
Modeling setup. We use different prior distributions depending on the task:
  • Node and edge predictions: Prior samples $p_{0}$ are drawn from a Gaussian distribution. Targets are either $-1$ or $1$, indicating binary decisions (e.g., whether a node is expanded or an edge is retained).
  • Hyperedge expansion: Similar to node and edge predictions, with targets $-1$, 00, or $1$, encoding the number of expansions (none, one, or two).
  • Budget fractions: Prior samples $p_{0}$ are drawn from a Dirichlet distribution with concentration parameter $\alpha=1.5$, linearly mapped to $[-1,1]$ via $12x-$ 1. The target is the budget fraction of each child node, linearly mapped to $[-1,1]$ via $12x-$ 1. If a cluster is not expanded, the corresponding budget becomes $1$. Parent cluster budgets are encoded using sinusoidal positional encodings [42] (dimension 32, base frequency $10^{-4}$).
  • Feature generation: Following [20], we draw prior features from a Gaussian and predict true node features, conditioned on the parent node’s feature using a FiLM layer [25].
Simplex projection via Von Neumann method. When modeling budget fractions, predictions must lie on the probability simplex. To ensure this, we project the model’s outputs using the Von Neumann projection [24], which finds the closest point (in Euclidean distance) on the simplex:
$\Delta^{K}=\left\{\mathbf{x}\in\mathbb{R}^{K}\mid x_{i}\geq 0,\sum_{i=1}^{K}x_{i}=1\right\}.$ (14)
  1. i)
    Sort $\mathbf{z}\in\mathbb{R}^{K}$ into a descending vector $\mathbf{u}$, such that $u_{1}\geq u_{2}\geq\dots\geq u_{K}$.
  2. ii)
    Find the smallest index $\rho\in\{1,\dots,K\}$ such that:
    $u_{\rho}-\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right)>0.$ (15)
  3. iii)
    Compute the threshold:
    $\tau=\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right).$ (16)
  4. iv)
    The projection is then:
    $\mathbf{x}^{*}=\max(\mathbf{z}-\tau,0).$ (17)
A.3 Additional Details
Perturbed expansion. Building on [11, 6], we augment Definitions 4 and 6—which are sufficient for reversing coarsening steps—with additional randomness to enhance generative quality. This modification is especially beneficial in low-data regimes where overfitting is a concern. Specifically, we introduce a probabilistic mechanism that supplements the set of edges $\tilde{\mathcal{E}}$ by randomly adding edges between node pairs on opposite sides of the bipartite graph that are within a fixed distance in $B$. The following definition extends the expansion process (Definition 4) to include this stochastic component.
Definition 9 (Perturbed hypergraph expansion).
Let $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ be a bipartite graph, and let $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ denote the left and right cluster size vectors. For a given radius $r\in\mathbb{N}$ and probability $0\leq p\leq 1$, we construct $\tilde{B}$ as in Definition 4. Additionally, for each pair of distinct nodes $\mathbf{v}_{L}{(p)}\in\tilde{\mathcal{V}_{L}}$ and $\mathbf{v}_{R}{(q)}\in\tilde{\mathcal{V}}_{R}$ that are within a distance of at most $12r+$ 1 in $B$, we independently add an edge $e_{\{p_{i},q_{j}\}}$ to $\tilde{\mathcal{E}}$ with probability $p$.
Spectral conditioning. In line with [43, 11], we incorporate spectral information—specifically, the principal eigenvalues and eigenvectors of the normalized Laplacian—as a form of conditioning during the generative process. This technique has been shown to improve the quality of generated graphs. To generate $B^{(l)}$ from its coarser form $B^{(l+1)}$, we leverage the approximate spectral invariance under coarsening. We compute the $k$ smallest non-zero eigenvalues and their corresponding eigenvectors from the normalized Laplacian matrix $\mathbf{\mathcal{L}}^{(l+1)}$ of $B^{(l+1)}$. These eigenvectors are processed using SignNet [41] to produce node embeddings for $B^{(l+1)}$. These embeddings are then propagated to the expanded nodes of $B^{(l)}$, helping to preserve structural coherence and facilitate cluster identification. The hyperparameter $k$ controls the number of spectral components used.
Minibatch OT-coupling. Minibatch OT-coupling [12, 13] has demonstrated effectiveness in image generation tasks by encouraging models to learn more linear and efficient generative trajectories. Rather than sampling from the prior and target distributions independently, the two distributions are sampled jointly through the following procedure. First, samples are drawn from the prior distribution (typically Gaussian), followed by sampling corresponding targets (e.g., images). Then, the prior samples are reindexed to minimize the total distance between each input and its associated output. This coupling yields faster convergence and smoother generative flows, thereby reducing the number of sampling steps required during inference.
We extend this idea to our setting by treating the children of an expanded node as a minibatch and applying OT-coupling to the prior samples associated with these nodes. Importantly, we avoid mixing prior samples between “non-equivalent nodes”, i.e., nodes for which swapping the targets results in a different target hypergraph, as doing so changes the learned distribution. To prevent this, we restrict OT-coupling to groups of nodes that only differ with respect to their starting noise, i.e., nodes that are equivalent both topologically and in terms of features.
To manage computational complexity, we adopt the reasonable assumption that such groups of equivalent nodes correspond to the children of an expanded cluster. This assumption holds in practice when the joint topology-feature distribution is sufficiently complex. As a result, our method achieves faster convergence and requires fewer sampling steps at inference time, with only a very negligible impact on training efficiency. In our framework, expanded clusters yield two child nodes, so minibatch OT-coupling reduces to evaluating two permutations per cluster—an operation that can be efficiently parallelized using tensor operations. Algorithm 1 summarizes the process.
Algorithm 1 Minibatch OT-coupling for coarsening/expansion strategy
1:Expanded bipartite representation $B$ with clusters $\{V^{(p)}\}$, each $V^{(p)}$ containing 1 or 2 nodes; target samples $\{x_{i}\}$
2:Reindexed noise samples $\{\tilde{z}_{i}\}$
3:Sample noise $\{z_{i}\}$ for each node in $B$
4:for all clusters $V^{(p)}\in B$ do
5:     if $|V^{(p)}|=1$ then
6:         No reassignment needed for singleton cluster
7:     else if $|V^{(p)}|=2$ then
8:         Let $i,j$ be the indices of the two nodes in $V^{(p)}$
9:         Let $x_{i},x_{j}$ be their corresponding targets
10:         Compute normal order cost:
$C_{\text{normal}}\leftarrow\|z_{i}-x_{i}\|^{2}+\|z_{j}-x_{j}\|^{2}$
11:         Compute swapped order cost:
$C_{\text{swap}}\leftarrow\|z_{j}-x_{i}\|^{2}+\|z_{i}-x_{j}\|^{2}$
12:         if $C_{\text{swap}}<C_{\text{normal}}$ then
13:              Swap $z_{i}\leftrightarrow z_{j}$
14:         end if
15:     end if
16:end for
17:return $\{\tilde{z}_{i}\}$ (reassigned noise samples)
Appendix B Proofs
B.1 Averaging Node Features for Clusters’ Features
Proposition 10.
Setting cluster features as the average of the nodes they contain minimizes the Mean Squared Error (MSE) between the fully expanded hypergraph and the original hypergraph.
Proof.
Let $H=(\mathcal{V},\mathcal{E})$ be a hypergraph with node set $\mathcal{V}$ and hyperedge set $\mathcal{E}$, and let $H_{l}=(\mathcal{V}_{l},\mathcal{E}_{l})$ denote the lifted hypergraph obtained by expanding each cluster $\mathcal{C}\subseteq\mathcal{V}$ of size $|C|$-clique. By construction, there exists a bijection $\phi:\mathcal{V}\to\mathcal{V}_{l}$ mapping each original node to its corresponding lifted node. Suppose each node $v\in\mathcal{V}$ is associated with a feature vector $\mathbf{x}_{v}\in\mathbb{R}^{d}$, and each cluster $\mathcal{C}\subseteq\mathcal{V}$ is assigned a cluster feature vector $\mathbf{x}_{\mathcal{C}}\in\mathbb{R}^{d}$, which is inherited by all lifted nodes $\phi(v)$ for $v\in\mathcal{C}$. Define the mean squared error between the original node features and the cluster features in the lifted hypergraph as:
$\mathrm{MSE}=\sum_{v\in\mathcal{V}}\|\mathbf{x}_{v}-\mathbf{x}_{C(v)}\|^{2},$ (18)
where $C(v)$ denotes the cluster containing node $v$.
To minimize the MSE, it suffices to minimize, for each cluster $\mathcal{C}$,
$J_{\mathcal{C}}(\mathbf{x}_{\mathcal{C}})=\sum_{v\in\mathcal{C}}\|\mathbf{x}_{v}-\mathbf{x}_{\mathcal{C}}\|^{2}.$ (19)
Since $J_{\mathcal{C}}$ is a convex quadratic function in $\mathbf{x}_{\mathcal{C}}$, we find its minimum by setting the gradient to zero:
$\nabla_{\mathbf{x}_{\mathcal{C}}}J_{\mathcal{C}}(\mathbf{x}_{C})=\sum_{v\in\mathcal{C}}2(\mathbf{x}_{\mathcal{C}}-\mathbf{x}_{v})=2|\mathcal{C}|\mathbf{x}_{\mathcal{C}}-2\sum_{v\in\mathcal{C}}\mathbf{x}_{v}=\mathbf{0}.$ (20)
Solving for $\mathbf{x}_{\mathcal{C}}$, we obtain
$\mathbf{x}_{\mathcal{C}}=\frac{1}{|\mathcal{C}|}\sum_{v\in\mathcal{C}}\mathbf{x}_{v},$ (21)
which is the arithmetic mean of the feature vectors in the cluster $\mathcal{C}$. ∎
B.2 Minibatch OT-coupling
Let $B$ be a bipartite graph. Let $\mathbf{b}$ be its current budget repartition, $\mathbf{F}_{L}$ its node features, and $\mathbf{F}_{R}$ its hyperedge features. Denote $\mathbf{x}=(\mathbf{v}_{L},\mathbf{v}_{R},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$, i.e., the predictions of the model. In the following, all distributions are conditioned on $\mathbf{b}$, $\mathbf{F}_{L}$ and $\mathbf{F}_{R}$.
Proposition 11 (Marginal preservation).
Under the OT-coupling of Algorithm 1, the joint distribution:
$q(\mathbf{x}_{0},\mathbf{x}_{1})$
has marginals:
$q_{0}(\mathbf{x}_{0}),\quad q_{1}(\mathbf{x}_{1}).$
Definition 12 (Isomorphism of bipartite graphs in our setting).
Let
$B_{1}=(\mathcal{V}_{L}^{1},\mathcal{V}_{R}^{1},\mathcal{E}^{1},\mathbf{b}^{1},\mathbf{F}_{L}^{1},\mathbf{F}_{R}^{1}),\quad B_{2}=(\mathcal{V}_{L}^{2},\mathcal{V}_{R}^{2},\mathcal{E}^{2},\mathbf{b}^{2},\mathbf{F}_{L}^{2},\mathbf{F}_{R}^{2}),$
be bipartite graphs.
We say $B_{1}\cong B_{2}$ (are isomorphic in our setting) if there exist bijections
$\sigma_{L}\colon\mathcal{V}_{L}^{1}\to\mathcal{V}_{L}^{2},\quad\sigma_{R}\colon\mathcal{V}_{R}^{1}\to\mathcal{V}_{R}^{2}$
such that:
  1. 1.
    Edge structure is preserved: $(v,w)\in\mathcal{E}^{1}\iff(\sigma_{L}(v),\sigma_{R}(w))\in\mathcal{E}^{2}$,
  2. 2.
    Budgets are preserved: $\mathbf{b}^{1}(v)=\mathbf{b}^{2}(\sigma_{L}(v))\text{ for all }v\in\mathcal{V}_{L}^{1},\quad\mathbf{b}^{1}(w)=\mathbf{b}^{2}(\sigma_{R}(w))\text{ for all }w\in\mathcal{V}_{R}^{1}$,
  3. 3.
    Node and hyperedge features are preserved: $\mathbf{F}_{L}^{1}(v)=\mathbf{F}_{L}^{2}(\sigma_{L}(v)),\quad\mathbf{F}_{R}^{1}(w)=\mathbf{F}_{R}^{2}(\sigma_{R}(w))$.
Proof.
Let $f$ be an arbitrary test function defined on bipartite graphs. Algorithm 1 swaps noise samples between nodes that are equivalent in the sense that their target graphs (conditioned on the same topology and conditioning features) remain isomorphic after swapping. That is, if $\mathbf{x}_{0}$ and $\mathbf{x}_{0}^{\prime}$ differ only by such a swap, then the resulting graphs are isomorphic: $B(\mathbf{x}_{0})\cong B(\mathbf{x}_{0}^{\prime})$. Let $\sigma$ be the bijective reindexing function corresponding to this isomorphism. Since $f$ is defined on graphs and graphs are invariant under isomorphism, we have:
$f(\mathbf{x}_{0})=f(\sigma(\mathbf{x}_{0})).$
Therefore:
$\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{x}_{1})}[f(\mathbf{x}_{0})]$ $=\mathbb{E}_{q(\mathbf{x}_{1})}\left[\mathbb{E}_{q(\mathbf{x}_{0}\mid\mathbf{x}_{1})}[f(\mathbf{x}_{0})]\right]$ (22)
$=\mathbb{E}_{q(\mathbf{x}_{1})}\left[\mathbb{E}_{q(\mathbf{x}_{0})}[f(\sigma(\mathbf{x}_{0}))]\right]$ (23)
$=\mathbb{E}_{q(\mathbf{x}_{1})}\left[\mathbb{E}_{q(\mathbf{x}_{0})}[f(\mathbf{x}_{0})]\right]$ (24)
$=\mathbb{E}_{q(\mathbf{x}_{0})}[f(\mathbf{x}_{0})].$ (25)
Thus, the marginal distribution over $\mathbf{x}_{0}$ remains unchanged. The same argument applies symmetrically for $\mathbf{x}_{1}$ by using $\sigma^{-1}$, concluding the proof. ∎
Appendix C Complexity Analysis
In this section, we investigate the asymptotic complexity of our proposed algorithm, which extends the methodology introduced by [11] and [6]. To construct a hypergraph comprising $n$ nodes, $m$ hyperedges, and $k$ incidences, the algorithm sequentially produces a series of bipartite graphs $B^{(L)}=(\{1\},\{2\},\{(1,2)\}),B^{(L-1)},\dots,B^{(0)}=B$, where the final graph $B$ corresponds to the bipartite representation of the generated hypergraph. We use $n$, $m$, and $k$ to denote, respectively, the number of nodes, hyperedges, and incidences in the hypergraph, and as the number of left-side nodes, right-side nodes, and edges in the corresponding bipartite graph.
[…]
Next, we assess the computational cost of generating a single expansion step. At level $l=L$, this consists of creating a pair of connected nodes, initializing features as matrices of zeros, initializing budget as the targeted node count, and predicting the expansion vectors $\mathbf{v}_{L}$ and $\mathbf{v}_{R}$—a process with constant complexity $\mathcal{O}(1)$. For levels $0\leq l<L$, given $B^{(l+1)}$ and expansion vectors $\mathbf{v}^{(l+1)}_{L}$, $\mathbf{v}^{(l+1)}_{R}$, the algorithm constructs the expanded bipartite graph $\tilde{B}(B^{(l+1)},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)})$ in $\mathcal{O}(n+m)$ time. It then samples $\mathbf{v}_{L}^{(l)}$, $\mathbf{v}_{R}^{(l)}$, $\mathbf{e}^{(l)}$, $\mathbf{f}$, $\mathbf{F}_{L}^{\text{refine}}$, and $\mathbf{F}_{R}^{\text{refine}}$, and constructs the refined graph $B^{(l)}=B(\tilde{B}^{(l)},\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$. Letting $v^{L}_{\max}$ and $v^{R}_{\max}$ be the maximum cluster sizes, the incidence count in $\tilde{B}^{(l)}$ is bounded by $k_{l}\leq k_{l+1}v^{L}_{\max}v^{R}_{\max}$.
The sampling process queries a denoising model a constant number of times per step. The complexity is thus governed by the architecture. In our case, since bipartite graphs are triangle-free, the Local PPGN model [11] has linear complexity $\mathcal{O}(n+m+k)$. Embedding computation for $B^{(l)}$ similarly costs $\mathcal{O}(n+m+k)$. This includes calculating the top $K$ eigenvalues/eigenvectors of the Laplacian via the method from [44], with complexity $\mathcal{O}\left(K(n_{l+1}+m_{l+1}+k_{l+1})\right)$, and embedding via SignNet, also linear in graph size due to fixed $K$.
The final transformation from the bipartite graph to a hypergraph—by collapsing right-side nodes into hyperedges—has a cost of $\mathcal{O}(m+k)$. Under these assumptions, the total complexity to generate a hypergraph $H$ with $n$ nodes, $m$ hyperedges, and $k$ incidences is $\mathcal{O}(n+m+k)$.
Appendix D Coarsening Sequence Sampling
[…]
  • The set is stochastically rejected with probability $1-\lambda$.
  • If not rejected:
    • Overlap check: If the contraction set overlaps with any previously accepted contraction, it is discarded.
    • Coarsening attempt: Otherwise, we compute tentative coarsened representations $C_{\text{temp}}$ and $B_{\text{temp}}$.
    • Cluster constraint check: If all right-side clusters in $B_{\text{temp}}$ contain at most three nodes, the contraction is accepted.
    • Update step: When a contraction is accepted, we:
        Sum the budgets of the nodes in the contraction set to define the new cluster budget.
        Compute the new cluster’s node features as a weighted average of the original features, using node budgets as weights.
The loop terminates once the number of remaining nodes satisfies the stopping condition:
$|\mathcal{V}_{L}^{(l-1)}|-|\bar{\mathcal{V}}_{L}^{(l)}|>\text{red\_frac}\cdot|\mathcal{V}_{L}^{(l-1)}|,$
i.e., when the number of nodes on the left side (corresponding to the original hypergraph nodes) has been reduced by the sampled fraction. This framework is flexible, allowing a variety of cost functions $c$, contraction families $F$, reduction fraction ranges $[\rho_{\text{min}},\rho_{\text{max}}]$, and randomization parameters $\lambda$.
Practical considerations. To avoid oversampling overly small graphs during training, we follow the heuristic of [11]: when the current graph has fewer than 16 nodes, we fix the reduction fraction to $\rho=\rho_{\text{max}}$. Due to the constraint that no right-side cluster in $B^{(l)}$ may contain more than three nodes, achieving the target reduction fraction is not always feasible. However, we observe empirically that this rarely poses a problem when $\rho_{\text{max}}$ is reasonably small.
During training, we sample a coarsening sequence for each dataset graph, but only retain a randomly selected intermediate graph from the sequence. Thus, our practical implementation of Algorithm 2 is designed to return a single coarsened graph with associated features and budgets, rather than the full sequence $\pi$.
To improve efficiency, we incorporate the caching mechanism introduced in [11]. Once a coarsening sequence is generated, its levels are cached. During training, a random level is selected, returned, and then removed from the cache. A new sequence is generated only when the cache for a particular graph is depleted, avoiding unnecessary recomputation.
Hyperparameters. In all experiments described in Section 4, we use the following settings:
  • Contraction family: The set of all edges in the clique representation, i.e., $F(C)=\mathcal{E}$, for a weighted clique expansion $C=(\mathcal{V},\mathcal{E})$.
  • Cost function: Local Variation Cost [45] with a preserving eigenspace size of $k=8$.
  • Reduction fraction range: $[\rho_{\text{min}},\rho_{\text{max}}]=[0.1,0.3]$.
Algorithm 2 Hypergraph coarsening sequence sampling: Randomized iterative coarsening of a hypergraph. At each step, contraction sets are selected based on cost, while ensuring right-side clusters never merge more than three at a time. Accepted contractions update the hypergraph structure, node budgets, and features.
1:contraction family $F$, cost function $c$, reduction fraction range $[\rho_{\text{min}},\rho_{\text{max}}]$, randomization parameter $\lambda$
2:hypergraph $H$ with $n$ nodes and $m$ hyperedges; node features $\mathbf{F}_{L}$; hyperedge features $\mathbf{F}_{R}$
3:coarsening sequence $\pi=(H^{(0)},\ldots,H^{(L)})\in\Pi_{F}(H)$
4:function HypergraphCoarseningSeq($H$)
5:     $H^{(0)}\leftarrow H$
6:     $B^{(0)}\leftarrow\text{BipartiteRepresentation}(H^{(0)})$
7:     $C^{(0)}\leftarrow\text{WeightedCliqueExpansion}(H^{(0)})$
8:     $\mathbf{b}_{L}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{n}$triangleright Initial node budgets
9:     $\mathbf{b}_{R}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{m}$triangleright Initial hyperedge budgets
10:     $\pi\leftarrow(B^{(0)},\mathbf{b}_{L}^{(0)},\mathbf{F}_{L},\mathbf{F}_{R})$
11:     $l\leftarrow 0$
12:     while $|\mathcal{V}_{L}^{(l)}|>1$ do
13:         $l\leftarrow l+1$
[…]
15:         $f\leftarrow c(\cdot,C^{(l-1)},(\mathcal{P}^{(l-1)},\ldots,\mathcal{P}^{(0)}))$triangleright Cost function
[…]
17:         for $S\in\text{SortedByCost}(F(C^{(l-1)}))$ do
18:              if $\text{Random}()>\lambda$ then
19:                  if $S\cap(\bigcup_{P\in\text{accepted\_contractions}}P)=\emptyset$ then
20:                       $C_{\text{temp}}\leftarrow\text{CoarsenCliqueExpansion}(C^{(l-1)},S)$
21:                       $B_{\text{temp}}\leftarrow\text{CoarsenBipartite}(B^{(l-1)},S)$
[…]
24:                           $C^{(l)}\leftarrow C_{\text{temp}}$, $B^{(l)}\leftarrow B_{\text{temp}}$
25$\triangleright$ Update budgets and features for the new cluster
26:                           Let $\text{Let }S=\{v_{1},\ldots,v_{k}\},\text{ and the new node be }v^{*}$
27:                           $\mathbf{b}_{L}^{(l)}[v^{*}]\leftarrow\sum_{i=1}^{k}\mathbf{b}_{L}^{(l-1)}[v_{i}]$
28:                           $\mathbf{F}_{L}^{(l)}[v^{*}]\leftarrow\frac{1}{\mathbf{b}_{L}^{(l)}[v^{*}]}\sum_{i=1}^{k}\mathbf{b}_{L}^{(l-1)}[v_{i}]\cdot\mathbf{F}_{L}^{(l-1)}[v_{i}]$
29:                       end if
30:                  end if
31:              end if
32:              if $|\mathcal{V}_{L}^{(l-1)}|-|\bar{\mathcal{V}}_{L}^{(l)}|>\text{red\_frac}\cdot|\mathcal{V}_{L}^{(l-1)}|$ then
33:                  break
34:              end if
35:         end for
36:         $\pi\leftarrow\pi\cup\{B^{(l)},\mathbf{b}_{L}^{(l)},\mathbf{F}_{L}^{(l)}\}$
37:     end while
38:     return $\pi$
39:end function
Appendix E Training and Sampling Procedures
In this section, we present the complete training and inference procedures, detailed in Algorithms 4 and 5. Both pipelines rely on node embeddings produced by Algorithm 3.
Algorithm 3 Node embedding computation: Here we describe the way the left and right side node embeddings are computed for a given bipartite representation of a hypergraph. Embeddings are computed for the input bipartite representation and then replicated according to the cluster size vectors.
1:number of spectral features $k$
2:bipartite representation $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$, spectral feature model SignNetθ, cluster size vector $\mathbf{v}_{L}$ and $\mathbf{v}_{R}$
3:node embeddings computed for all nodes in $\mathcal{V}_{L}$ and $\mathcal{V}_{R}$ and replicated according to $\mathbf{v}_{L}$ and $\mathbf{v}_{R}$
4:function Embeddings($B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$, SignNetθ, $\mathbf{v}_{L}$, $\mathbf{v}_{R}$)
5:     if $k=0$ then
6:         $\mathbf{H}=[h^{(1)},\ldots,h^{(|\mathcal{V}|)}]\overset{i.i.d.}{\sim}\mathcal{N}(0,I)$triangleright Sample random embeddings
7:     else
8:         if $k<|\mathcal{V}|$ then
9:              $[\lambda_{1},\ldots,\lambda_{k}],[u_{1},\ldots,u_{k}]\leftarrow\text{EIG}(B)$triangleright Compute $k$ spectral features
10:         else
11:              $[\lambda_{1},\ldots,\lambda_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1}],[u_{1},\ldots,u_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1}]\leftarrow\text{EIG}(B)$triangleright Compute $|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1$ spectral features
12:              $[\lambda_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|},\ldots,\lambda_{k}],[u_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|},\ldots,u_{k}]\leftarrow[0,\ldots,0],[0,\ldots,0]$triangleright Pad with zeros
13:         end if
14:         $\mathbf{H}=[h^{(1)},\ldots,h^{(|\mathcal{V}_{L}|+|\mathcal{V}_{R}|)}]\leftarrow\text{SignNet}_{\theta}([\lambda_{1},\ldots,\lambda_{k}],[u_{1},\ldots,u_{k}],B)$
15:     end if
16:     $\tilde{B}=(\mathcal{V}_{L}^{(1)}\cup\cdots\cup\mathcal{V}_{L}^{(p_{l})},\mathcal{V}_{R}^{(1)}\cup\cdots\cup\mathcal{V}_{R}^{(p_{r})},\tilde{\mathcal{E}})\leftarrow\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R})$triangleright Expand as per Definition 4
17:     set $\tilde{B}$ s.t. for all $p_{L}\in[|\mathcal{V}_{L}|]$ and all $p_{R}\in[|\mathcal{V}_{R}|]$: for all $\mathbf{v}_{L}^{(p_{i})}\in\mathcal{V}_{L}^{(p_{l})}$, $\tilde{\mathbf{H}}[p_{i}]=\mathbf{H}[p_{l}]$ and for all $\mathbf{v}_{R}^{(p_{i})}\in\mathcal{V}_{R}^{(p_{r})}$, $\tilde{\mathbf{H}}[p_{i}]=\mathbf{H}[p_{r}]$triangleright Replicate embeddings
18:     return $\tilde{\mathbf{H}}$
19:end function
Algorithm 4 End-to-end training procedure: This describes the entire training procedure for our model.
1:number of spectral features $k$ for node embeddings
2:dataset $\mathcal{D}=\{H_{1},\ldots,H_{N}\}$, denoising model GNNθ, spectral feature model SignNetθ
3:trained model parameters $\theta$
4:function Train($\mathcal{D}$, GNNθ, SignNetθ)
5:     while not converged do
6:         $H\sim\text{Uniform}(\mathcal{D})$triangleright Sample graph
7:         $(B^{(0)},\ldots,B^{(L)})\leftarrow\text{RndRedSeq}(H)$triangleright Sample coarsening sequence by Algorithm 2
8:         $l\sim\text{Uniform}(\{0,\ldots,L\})$triangleright Sample level
9:         if $l=0$ then
10:              $\mathbf{v}_{L}^{(0)}\leftarrow 1$, $\mathbf{v}_{R}^{(0)}\leftarrow 1$
11:         else
12:              set $\mathbf{v}_{L}^{(l)}$ and $\mathbf{v}_{R}^{(l)}$ such that the node sets of $\tilde{B}(B^{(l)},\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)})$ equals that of $B^{(l-1)}$
13:         end if
14:         if $l=L$ then
15:              $B^{(l+1)}\leftarrow B^{(l)}=(\{1\},\{2\},\{(1,2)\},\mathbf{b}=size(H),\mathbf{F}_{L}=0,\mathbf{F}_{R}=0)$
16:              $\mathbf{v}_{L}^{(l+1)}\leftarrow 1$
17:              $\mathbf{v}_{R}^{(l+1)}\leftarrow 1$
18:              $\mathbf{e}^{(l)}\leftarrow 1$
19:         end if
20:         set $\mathbf{e}^{(l)}$, $\mathbf{f}$, $\mathbf{F}_{L}^{\text{refine}}$ and $\mathbf{F}_{R}^{\text{refine}}$ such that $B(\tilde{B}(B^{(l+1)},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)}),\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})=B^{(L)}$
21:         $\mathbf{H}^{(l)}\leftarrow\text{Embeddings}(B^{(l+1)},\text{SignNet}_{\theta},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)})$triangleright Compute node embeddings
22:         $\hat{\rho}\leftarrow 1-(n^{(l)}/n^{(l-1)})$, with $n^{(l)}$ and $n^{(l-1)}$ being the size of the left side of $B^{(l)}$ and $B^{(l-1)}$
23:         $D_{\theta}\leftarrow\text{GNN}_{\theta}(\cdot,\cdot,\tilde{B}^{(l)},\mathbf{H}^{(l)},n^{(0)},\rho)$, where $n^{(0)}$ is the size of the left side of $B^{(0)}$
24:         take gradient descent step on $\nabla_{\theta}\text{DiffusionLoss}(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)},\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}},D_{\theta})$
25:     end while
26:     return $\theta$
27:end function
Algorithm 5 End-to-end sampling procedure with deterministic expansion size: This describes the sampling procedure. Note that this assumes that the maximum cluster sizes are 2 and 3, which is the case when using edges of the clique representation as the contraction set family for model training.
1:reduction fraction range $[\rho_{\text{min}},\rho_{\text{max}}]$
2:target hypergraph size $N$, denoising model GNNθ, spectral feature model SignNetθ
3:sampled hypergraph $H=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=N$
4:function Sample($N$, GNNθ, SignNetθ)
5:     $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})\leftarrow(\{1\},\{2\},\{(1,2)\},N,0,0)$triangleright Start with a minimal bipartite graph
6:     $\mathbf{v}_{L}\leftarrow[1]$, $\mathbf{v}_{R}\leftarrow[1]$triangleright Initial cluster size vectors
7:     while $|\mathcal{V}_{L}|<N$ do
8:         $\mathbf{H}\leftarrow\text{Embeddings}(B,\text{SignNet}_{\theta},\mathbf{v}_{L},\mathbf{v}_{R})$triangleright Compute node embeddings
9:         $n\leftarrow\|\mathbf{v}_{L}\|_{1}$
10:         $\rho\sim\text{Uniform}([\rho_{\text{min}},\rho_{\text{max}}])$triangleright random reduction fraction
11:         set $n^{+}$ s.t. $n^{+}=\lceil\rho(n+n^{+})\rceil$triangleright number of left side nodes to add
12:         $n^{+}\leftarrow\min(n^{+},N-n)$triangleright ensure not to exceed target size
13:         $\hat{\rho}\leftarrow 1-(n/(n+n^{+}))$triangleright actual reduction fraction
14:         $D_{\theta}\leftarrow\text{GNN}_{\theta}(\cdot,\cdot,\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R}),\mathbf{H},N,\hat{\rho})$
15:         $(\mathbf{v}_{L})_{0},(\mathbf{v}_{R})_{0},(\mathbf{e})_{0},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}}\leftarrow\text{Sample}(D_{\theta})$triangleright Sample features
16:         set $\mathbf{v}_{L}$ s.t. for $i\in[n]$: $\mathbf{v}_{L}[i]=2$ if |$|\{j\in[n]\mid(\mathbf{v}_{L})_{0}[j]\geq(\mathbf{v}_{L})_{0}[i]\}|\geq n^{+}$ and $v[i]=1$ otherwise
17:         set $\mathbf{v}_{R}$ s.t. for $i\in[|(\mathbf{v}_{R})_{0}|]$: $\mathbf{v}_{R}[i]=1$ if $(\mathbf{v}_{R})_{0}<1.66$, $\mathbf{v}_{R}[i]=2$ if $(\mathbf{v}_{R})_{0}<2.33$ and $\mathbf{v}_{R}[i]=3$ otherwise
18:         set $\mathbf{e}$ s.t. for $i\in[|(\mathbf{e})_{0}|]$: $\mathbf{e}[i]=1$ if $(\mathbf{e})_{0}>0.5$ and $\mathbf{e}[i]=0$ otherwise
19:         $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})\leftarrow B(\tilde{B},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$triangleright Refine as per Definition 6
20:     end while
21:     build $H$ from its bipartite representation $B$
22:     return $H$
23:end function
Appendix F Experimental Details
For all experiments, we use embeddings with 32 dimensions for edge selection vectors and node and hyperedge expansion numbers. When they exist, features are embedded with 128 dimensions. SignNet always has 5 layers and a hidden dimension of 128. Positional encodings always have 32 dimensions. We always use 10 layers of Local PPGN. We always use 25 sampling steps. All experiments are run for 24 hours on a single L40S. We use 8 CPU workers.
F.1 Experimental Details: Unfeatured Hypergraphs
Datasets. Our experiments utilize five datasets: three synthetic and two real-world, consistent with those described in [6]:
  • Stochastic Block Model (SBM) hypergraphs [31]: Constructed with 32 nodes split evenly into two groups. Every hyperedge connects three nodes. Hyperedges are sampled with probability 0.05 within groups and 0.001 between groups.
  • Ego hypergraphs [32]: Created by generating an initial hypergraph of 150–200 nodes with 3000 randomly sampled hyperedges (up to 5 nodes each), then extracting an ego-centric subgraph by selecting a node and retaining only hyperedges that include it.
  • Tree-structured hypergraphs [33]: A tree with 32 nodes is generated using networkx, followed by grouping adjacent tree edges into hyperedges. Each hyperedge contains up to 5 nodes.
  • ModelNet40 meshes [34]: Hypergraphs are derived from mesh topologies of selected ModelNet40 categories. To simplify computation, meshes are downsampled to fewer than 1000 vertices by iteratively merging nearby vertices. Duplicate triangles are removed, and the resulting low-poly mesh is converted into a hypergraph. We focus on the bookshelf and piano categories.
All datasets are divided into 128 training, 32 validation, and 40 testing hypergraphs.
Evaluation Metrics. We evaluate generated hypergraphs using the same suite of metrics as [6]:
  • NodeNumDiff: Average absolute difference in node count between generated and reference hypergraphs.
  • NodeDegreeDistrWasserstein: Wasserstein distance between node degree distributions of generated and reference hypergraphs.
  • EdgeSizeDistrWasserstein: Wasserstein distance between hyperedge size distributions.
  • Spectral: Maximum Mean Discrepancy (MMD) between Laplacian spectra.
  • Uniqueness: Fraction of generated hypergraphs that are non-isomorphic to one another.
  • Novelty: Fraction of generated hypergraphs that are non-isomorphic to training samples.
  • CentralityCloseness, CentralityBetweenness, CentralityHarmonic: Wasserstein distances computed between centrality distributions (on edges for $s=1$). For details see [46].
  • ValidEgo: For the hypergraphEgo dataset only, proportion of generated hypergraphs that contain a central node shared by all hyperedges.
  • ValidSBM: For the hypergraphSBM dataset only, proportion of generated graphs that retain the original intra- and inter-group connectivity patterns.
  • ValidTree: For the hypergraphTree dataset only, proportion of generated samples that preserve tree structure.
Baselines. We compare our method against the following baselines:
  • HyperPA [26]: A heuristic approach for hypergraph generation.
  • Image-based models: We design three baseline models—Diffusion, GAN, and VAE—that operate on incidence matrix representations of hypergraphs:
    • Each model is trained to produce binary images where white pixels signify node-hyperedge membership.
    • To normalize input sizes, incidence matrices are permuted randomly and padded with black pixels.
    • Generated images are thresholded to obtain binary incidence matrices.
  • HYGENE [6]: A hierarchical diffusion-based generator using reduction, expansion, and refinement steps.
Specific hyperparameters. We use $\lambda=0.3$, our Local PPGN layers have a dimension of 128, and the hidden dimension for our MLP is 256. We use perturbed hypergraph expansion with a radius of 2 and dropout of 0.5. Our model has 4M parameters.
Detailed numerical results.
Table 6: Detailed numerical results for ablation studies.
SBM Hypergraphs
Node Budget Minibatch OT Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
77.7% 0.020 0.889 0.041 0.006 5.118 0.015 0.002
79.9% 0.123 0.815 0.023 0.006 6.3 0.015 0.003
71.6% 0.039 1.413 0.005 0.013 9.862 0.006 0.004
81.4% 0.010 0.603 0.005 0.005 3.962 0.009 0.002
Ego Hypergraphs
Node Budget Minibatch OT Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
99.0% 2.383 0.239 0.139 0.006 5.720 0.007 4e-6
98.5% 1.721 0.223 0.184 0.010 3.784 6.7e-4 1.8e-5
100% 0.459 0.129 0.279 0.010 3.010 0 0
100% 0.162 0.171 0.129 0.007 6.061 0 0
Tree Hypergraphs
Node Budget Minibatch OT Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
92.2% 0.005 0.066 0.161 0.011 0.406 0.022 0.006
83.8% 0 0.045 0.083 0.006 0.515 0.025 0.027
74.5% 0.005 0.039 0.050 0.003 0.528 0.042 0.026
82.8% 0 0.043 0.046 0.002 0.258 0.040 0.027
ModelNet Bookshelf
Node Budget Minibatch OT NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
8.615 5.951 0.043 0.027 102.547 0.113 0.003
7.225 4.183 0.109 0.037 63.778 0.197 0.003
1.525 5.232 0.069 0.013 79.016 0.087 0.003
0.875 5.600 0.087 0.019 32.723 0.113 0.002
ModelNet Piano
Node Budget Minibatch OT NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
8.125 3.546 0.075 0.026 42.991 0.102 0.002
2.375 5.529 0.160 0.028 50.349 0.097 0.002
11.275 4.127 0.072 0.027 31.312 0.168 0.002
0.125 1.878 0.028 0.036 43.326 0.072 0.001
F.2 3D Meshes
Datasets. Datasets for meshes are taken from Manifold40 [35], which is a reworked version of ModelNet40 [34] to obtain manifold and watertight meshes. Meshes are subsequently coarsened to obtain low-poly versions of 50 vertices and 100 triangles. We use two classes:
  • Airplane comprising 682 training samples, 21 validation samples, and 23 testing samples.
  • Bench comprising 144 training samples, 19 validation samples, and 30 testing samples.
Metrics. We use the same metrics as for hypergraphs, which are: NodeNumDiff, NodeDegree, EdgeSize, and Spectral. To this we add a metric NearChamDist which computes the Chamfer distance between a point cloud sampled from the surface of the generated mesh and equivalent point clouds sampled from the validation/test set meshes.
Baselines. We compare against a simple sequential baseline:
  1. 1.
    Our model (4M parameters) is trained for 1M steps on the topology of meshes without learning to generate the features.
  2. 2.
    A simple Local PPGN model (4M parameters) is trained for 20 epochs as a flow-matching model to learn to generate the 3D positions, with the topology fixed.
  3. 3.
    We use the best checkpoint of the first model to generate the topology, then apply the second model on this topology to generate the 3D positions.
Specific hyperparameters. We use $\lambda=0.1$, our Local PPGN layers have a dimension of 200, and the hidden dimension for our MLP is 300. We use perturbed hypergraph expansion with a radius of 2 and dropout of 0.5. Our model has 6M parameters.
Detailed numerical results.
Table 7: Evaluation on Manifold40 Bench and Airplane datasets.
Method Manifold40 Bench Manifold40 Airplane
Cham Dist $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Centr Harm $\downarrow$ Centr Clos $\downarrow$ Centr Betw $\downarrow$ Cham Dist $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Centr Harm $\downarrow$ Centr Clos $\downarrow$ Centr Betw $\downarrow$
Sequential 0.143 0.367 0.801 0.004 0.007 4.964 0.087 0.006 0.117 0.078 0.332 0.011 0.015 3.131 0.035 0.007
[…]
FAHNES 0.073 0.067 0.581 0.008 0.014 5.28 0.047 0.004 0.049 0 0.304 0.033 0.015 1.837 0.231 0.004
F.3 Molecules
Datasets. We use the QM9 dataset with implicit hydrogens. We generate 20,000 molecules for the metrics.
Metrics. We use the same metrics as DeFoG [30], that is:
  • Valid: Fraction of generated samples that can be sanitized using RDKit.
  • Novel: Fraction of generated samples whose SMILES are not present in the training dataset.
  • Unique: Fraction of generated samples whose SMILES are unique.
  • FCD: Fréchet ChemNet Distance [39] between the generated samples and test set.
Baselines.
Specific hyperparameters. We use $\lambda=0.1$, our Local PPGN layers have a dimension of 128, and the hidden dimension for our MLP is 256. We do not use perturbed hypergraph expansion. Our model has 4M parameters.
Appendix G Validation Curves
Figure 4: Valid SBM with and without budgets.
Figure 5: Valid Ego with and without budgets.
Figure 6: Valid Tree with and without budgets.
Figure 7: NodeNumDiff for Bookshelf with and without budgets.
Appendix H Examples of Coarsening Sequences
Figure 8: Examples of coarsening sequence for various meshes. Broad lines represent 2-edges.
Appendix I Comparison between Training and Generated Samples
Train samples
Generated samples
(i) Stochastic Block Model hypergraphs.
Train samples
Generated samples
(ii) Ego hypergraphs.
Train samples
Generated samples
(iii) Tree hypergraphs.
Train samples
Generated samples
(iv) Bookshelf meshes topology.
Train samples
Generated samples
(v) Piano meshes topology
Train samples
Generated samples (sequential)
Generated samples (joint)
(vi) Bench 3D meshes.
Train samples
Generated samples (sequential)
Generated samples (joint)
(vii) Airplane 3D meshes.
(viii) Generated QM9 molecules.

### v2

@@
-Feature-aware Hypergraph Generation via Next-Scale Prediction
+Feature-aware (Hyper)graph Generation via Next-Scale Prediction
 Abstract
-Hypergraphs generalize traditional graphs by allowing hyperedges to connect multiple nodes, making them well-suited for modeling complex structures with higher-order relationships, such as 3D meshes, molecular systems, and electronic circuits.
-While topology is central to hypergraph structure, many real-world applications also require node and hyperedge features.
-Existing hypergraph generation methods focus solely on topology, often overlooking feature modeling.
-In this work, we introduce FAHNES (feature-aware hypergraph generation via next-scale prediction), a hierarchical approach that jointly generates hypergraph topology and features.
-FAHNES builds a multi-scale representation through node coarsening, then learns to reconstruct finer levels via localized expansion and refinement, guided by a new node budget mechanism that controls cluster splitting.
-We evaluate FAHNES on synthetic hypergraphs, 3D meshes, and molecular datasets.
-FAHNES achieves competitive results in reconstructing topology and features, establishing a foundation for future research in featured hypergraph generative modeling.
+Graph generative models have shown strong results in molecular design but struggle to scale to large, complex structures.
+While hierarchical methods improve scalability, they usually ignore node and edge features, which are critical in real-world applications.
+This issue is amplified in hypergraphs, where hyperedges capture higher-order relationships among multiple nodes.
+Despite their importance in domains such as 3D geometry, molecular systems, and circuit design, existing generative models rarely support both hypergraphs and feature generation at scale.
+In this paper, we introduce FAHNES (feature-aware hypergraph generation via next-scale prediction), a hierarchical framework that jointly generates hypergraph topology and features.
+FAHNES builds multi-scale representations through node coarsening and refines them via localized expansion, guided by a novel node budget mechanism that controls granularity and ensures consistency across scales.
+Experiments on synthetic, 3D mesh and graph point cloud datasets show that FAHNES achieves state-of-the-art performance in jointly generating features and structure, advancing scalable hypergraph and graph generation.
 1 Introduction
+Generating discrete geometric structures such as graphs and hypergraphs is an important challenge in modern machine learning [1].
+These structures are central to applications ranging from molecular design and materials discovery to electronic circuits and 3D shape modeling [2, 3, 4, 5].
+Their ability to capture complex relationships—pairwise in graphs, and multi-way in hypergraphs—makes them an essential tool for representing and synthesizing structured data.
+In many real-world settings, topology alone is not enough: node and edge (or hyperedge) features often carry important semantic or geometric information, such as atom and bond types in molecules, or vertex coordinates in 3D meshes.
 Figure 1: Examples of generated featured hypergraphs by a sequential disjoint generation baseline and our model (FAHNES).
-Hypergraphs extend traditional graphs by allowing edges—called hyperedges—to connect more than two nodes, enabling a natural representation of complex, multi-way interactions.
-This expressiveness makes hypergraphs well-suited for modeling systems in molecular chemistry, electronic circuit design, and 3D geometry.
-As a result, they have been applied in diverse areas such as drug discovery [1], contagion modeling [2], recommendation systems [3], molecular biology [4], and urban planning [5].
-Despite their wide applicability, generative models for hypergraphs remain relatively underexplored.
-Existing approaches primarily focus on generating topology alone [6], neglecting the generation of node and hyperedge features.
-As shown in Figure 1, the sequential generation of topology followed by features is ineffective.
-Joint generation is particularly challenging for hypergraphs due to their variable-size hyperedges and higher structural complexity.
-Moreover, existing methods for featured graph generation use flat, non-hierarchical architectures, which scale poorly [7, 8, 9].
-In this work, we introduce a feature-aware hypergraph generation via next-scale prediction (FAHNES) model, a novel hierarchical approach for jointly generating hypergraph topology and features.
-Inspired by multi-scale generative strategies in image synthesis [10], FAHNES leverages a coarsening-expansion strategy [6, 11]: during training, it merges nodes based on connectivity to build a hierarchical representation, then learns to progressively recover finer-scale hypergraphs.
-A key contribution is the use of a budget mechanism that tracks the number of permissible expansions per cluster.
-This enables global structural constraints to influence local generation, helping to mitigate the limited receptive fields of deep models.
-Additionally, we adapt flow-matching optimal transport (OT) coupling techniques [12, 13] to learn straighter and more consistent flows during generation.
+Despite recent advances, existing methods for featured graph generation struggle to scale.
+Most of these approaches use flat architectures that model the entire structure at once, leading to quadratic computational and memory complexities [6, 7, 8].
+This limits their use to moderately sized graphs.
+To address scalability, hierarchical generation strategies have been proposed for both graphs and hypergraphs [9, 10].
+These methods build the structure in stages, starting from a coarse representation and progressively upsampling and refining it, which allows them to handle much larger topologies.
+However, existing hierarchical methods only focus on unfeatured structures [9, 10].
+Extending them to generate features is challenging because different regions of a graph or hypergraph often grow at uneven rates during the refinement process, making it difficult to maintain consistency across scales.
+Furthermore, sequentially generating topology first and then features—rather than modeling them jointly—is ineffective in complex settings, as illustrated in Figure 1.
+To overcome these limitations, we propose FAHNES  (feature-aware hypergraph generation via next-scale prediction), a hierarchical generative framework that jointly models topology and features for both graphs and hypergraphs [^1] Since hypergraphs are a generalization of graphs, we present FAHNES on hypergraphs.
+The explanation for graphs follows straightforwardly..
+FAHNES builds multi-scale representations [11] through node coarsening and reconstructs fine structures via localized expansion, while directly predicting features alongside structure at each stage.
+A fundamental component of our model is a node budget mechanism, which encodes local growth constraints and helps maintain consistency in regions that expand at different rates.
+In addition, we adapt minibatch flow-matching optimal-transport (OT) coupling [12, 13] to the hierarchical setting, improving stability and coherence during generation.
+This combination enables scalable generation of large, featured graphs or hypergraphs across diverse domains, from 3D meshes to point clouds.
 Our main contributions are:
-  • We introduce the first hierarchical model for joint generation of hypergraph topology and features targeting specific data distributions, such as 3D meshes and molecules.
-  • We propose a novel budget-based mechanism that improves global consistency during generation for hierarchical methods.
-  • We generalize minibatch OT-coupling to our hypergraph generative setting.
-  • We validate our approach on both synthetic and real-world datasets, including molecular structures and 3D meshes, showing promising results for hierarchical generative methods.
+  • We introduce the first scalable hierarchical model for the joint generation of hypergraph topology and features, targeting complex data such as 3D meshes and point clouds.
+  • We propose a novel budget-based mechanism that enhances global structural coherence in hierarchical generation.
+  • We extend minibatch OT-coupling to hierarchical graph or hypergraph generation, improving generative stability and quality.
+  • We validate FAHNES on both synthetic and real-world datasets, demonstrating strong performance in jointly generating topology and features.
 2 Related Work
@@
 Many of these previous methods jointly model node features and topology, but are limited to small graphs due to scalability challenges.
-For example, diffusion-based models [7, 8, 9] operate over entire graphs by gradually corrupting structures and features, then training models to denoise them.
-However, their scalability is constrained by the combinatorial nature of edge prediction in large graphs.
+Diffusion-based models [6, 7, 8] operate over complete graphs by gradually corrupting structures and features, then training models to denoise them.
+However, their scalability is constrained by the combinatorial number of possible edges in graphs.
 To mitigate this, hierarchical methods have been proposed.
-For example, Bergmeister et al.
 [9] introduced a scalable graph generation framework based on a coarsen-then-expand approach, merging nodes to form coarse representations and progressively reconstructing finer details.
-This framework was extended to hypergraphs by Gailhard et al.
-[6], which allows edges to connect more than two nodes.
+This framework was extended to hypergraphs by [10], which allows edges to connect more than two nodes.
 However, both methods focus exclusively on topology generation, neglecting node and hyperedge features essential for many applications.
@@
 Generative models that capture both higher-order structure and node/hyperedge features are critical in numerous domains.
-In molecular design, regular graph generative models struggle to accurately represent rings and scaffolds [7], which naturally correspond to hyperedges involving multi-atom interactions rather than pairwise bonds.
+In molecular design, regular graph generative models struggle to accurately represent rings and scaffolds [6, 18], which naturally correspond to hyperedges involving multi-atom interactions rather than pairwise bonds.
 Similarly, 3D shape modeling involves surfaces such as triangles, quads, or general polygons that extend beyond simple pairwise connectivity.
 Conventional approaches often rely on fixed topology, quantization, or autoregressive sequence modeling with transformers [19, 20], limiting their flexibility and scalability.
-Hypergraphs provide a more expressive framework by treating faces as hyperedges, enabling the joint generation of topology and features.
-In contrast to previous approaches that either target hypergraph topology alone or feature-aware graph generation, we introduce the first method that jointly models both hypergraph structure and features within a unified and scalable framework.
-FAHNES extends hierarchical generative methods with feature-aware expansions, a novel node budget mechanism, and flow-matching training adapted to our strategy.
-This combination enables effective generative modeling for featured hypergraphs.
-3 Feature-aware Hypergraph Generation via Next-scale Prediction
-((a)) Hypergraph
-((b)) Clique expansion
-((c)) Star expansion
-Figure 2: Comparison of a hypergraph and its clique and star expansions.
-[…]
+Hypergraphs provide a more general framework by treating faces as hyperedges, enabling the joint generation of topology and features.
+Unlike prior work that focuses solely on topology or relies on flat, non-scalable designs, we present the first unified, scalable approach for jointly modeling hypergraph structure and features.
+FAHNES enhances hierarchical generative modeling with feature integration, a novel node-budget mechanism, and flow-matching training tailored to our strategy.
+3 Feature-aware (Hyper)graph Generation via Next‑scale Prediction
+Notations.
+Throughout this paper, we use calligraphic letters, like $\mathcal{V}$, to represent sets, with their cardinality denoted by $|\mathcal{V}|$.
+Matrices are represented by bold uppercase letters (e.g., $\mathbf{A}$), while vectors are indicated by bold lowercase letters (e.g., $\mathbf{x}$).
+The transpose and point-wise multiplication operations are denoted by $(\cdot)^{\top}$ and $\odot$, respectively.
+$\lceil\cdot\rfloor$ denotes rounding to the nearest integer.
 Basic definitions.
@@
 Each edge $e\in\mathcal{E}$ corresponds to a pair $(u,v)$, indicating a connection between nodes $u$ and $v$.
-A bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ is a special case where the vertex set is split into two disjoint subsets $\mathcal{V}_{L}$ and $\mathcal{V}_{R}$, and edges exist only between the two parts, i.e., $\mathcal{E}\subseteq\mathcal{V}_{L}\times\mathcal{V}_{R}$.
+A bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ is a special case of a graph where the vertex set is split into two disjoint subsets $\mathcal{V}_{L}$ and $\mathcal{V}_{R}$, and edges exist only between the two parts, i.e., $\mathcal{E}\subseteq\mathcal{V}_{L}\times\mathcal{V}_{R}$.
 The full set of nodes is $\mathcal{V}=\mathcal{V}_{L}\cup\mathcal{V}_{R}$.
@@
 A hypergraph $H$ extends the concept of a graph and is specified by a pair $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the vertex set and $\mathcal{E}$ comprises hyperedges, with each $e\in\mathcal{E}$ being a subset of $\mathcal{V}$.
-The key distinction of hypergraphs lies in their ability to connect arbitrary numbers of vertices through a single hyperedge.
-Similar to graphs, hypergraphs can possess node and edge features, which are denoted $\mathbf{F}_{\mathcal{V}}$ and $\mathbf{F}_{\mathcal{E}}$.
+The main distinction of hypergraphs lies in their ability to connect arbitrary numbers of vertices through a single hyperedge.
+Similar to graphs, hypergraphs can possess node and hyperedge features, which are denoted $\mathbf{F}_{\mathcal{V}}$ and $\mathbf{F}_{\mathcal{E}}$.
 We consider two fundamental graph-based representations of hypergraphs: clique and star expansions.
-The clique expansion transforms a hypergraph $H$ into a graph $C=(\mathcal{V}_{c},\mathcal{E}_{c})$, where $\mathcal{E}_{c}=\{(u,v)\mid\exists~{}e\in\mathcal{E}:u,v\in e\}$.
+The clique expansion transforms a hypergraph $H$ into a graph $C=(\mathcal{V},\mathcal{E}_{c})$, where hyperedges are replaced by cliques, i.e., $\mathcal{E}_{c}=\{(u,v)\mid\exists~e\in\mathcal{E}:u,v\in e\}$.
 The star expansion converts a hypergraph $H$ into a bipartite structure $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E}_{b})$, where $\mathcal{V}_{L}=\mathcal{V}$, $\mathcal{V}_{R}=\mathcal{E}$, and $\mathcal{E}_{b}=\{(v,e)\mid v\in\mathcal{V}_{L},e\in\mathcal{V}_{R},v\in e\text{ in }H\}$.
 Left side nodes $\mathcal{V}_{L}$ represent the nodes of the hypergraph, while right side nodes $\mathcal{V}_{R}$ represent hyperedges.
-We provide a visual example of the clique and star expansions in Figure 2.
-Our work aims to develop a model capable of sampling from the underlying distribution of a given featured hypergraph dataset $(H_{1},\dots,H_{N})$, i.e., learning the joint probability of features and topology of this distribution of hypergraphs.
-For detailed mathematical proofs of all propositions and lemmas presented in this work, please consult the supplementary material.
+Our objective is to develop a generative model that can sample from the underlying distribution of a dataset of featured hypergraphs (or graphs) $(H_{1},\dots,H_{N})$, i.e., to learn the joint distribution over both topology and features.
+We present FAHNES in the more general setting of hypergraphs; since graphs are a special case of hypergraphs, the model can be directly adapted to graphs, as discussed in Section 3.6.
+Complete mathematical proofs of all propositions in this paper appear in Appendix A.
 3.1 Overview
-Previous work [6] simplifies hypergraph generation by reducing it to standard graph generation through two projections: i) the clique expansion where every hyperedge becomes a clique that links all its incident nodes, and ii) the star expansion, which produces a bipartite graph by introducing one node per hyperedge and connecting it to the nodes it contains.
-While conceptually elegant, this approach faces significant challenges: i) it struggles to produce the correct number of nodes, especially in large hypergraphs, and ii) it does not support the generation of node and hyperedge features, limiting its applicability to real-world, featured domains.
-To overcome the limitations of prior approaches, we propose two key contributions.
-First, we introduce a node budget mechanism, in which generation begins with a single super-node carrying the full node budget.
-At each generation step, this budget is recursively split among child nodes.
-This approach enables a more localized and structured control over the final node count, compared to previous methods that simply concatenate the desired size to node embeddings.
-Second, we propose hierarchical feature generation, where each expansion step predicts the mean feature of its future child nodes, conditioned on the parent’s feature.
-This hierarchical feature generation scheme is thus a generalization of [20] to hypergraphs, where the prediction at current scale conditions the predictions at next scale.
-To train FAHNES, we adopt the flow-matching framework [21], where the generation process is formulated as learning to reverse a known noising procedure.
-Specifically, the model is trained to recover the true values for node and hyperedge expansions, edge deletions, budget allocations, and features, all of which are subjected to noise.
-Our workflow is illustrated in Figure 3.
+FAHNES follows a hierarchical pipeline as shown in Figure 2: i) during training, multi-scale representations of input (hyper)graphs are produced; ii) a model learns to reconstruct a higher scale from a lower one.
+Our approach adopts a coarsening–expansion framework [10].
+We first perform standard spectrum-preserving coarsening [21] on the clique expansion of the hypergraph to obtain the multi-scale representation.
+The process is then reversed through expansion and refinement, reconstructing both topology and features at progressively finer scales.
+Expansion and refinement are learned on the bipartite representation using the flow-matching framework [22], treating generation as the inverse of the coarsening process.
 Figure 2: Our framework adopts a coarsening-expansion strategy.
@@
 Node features are averaged during merging, and budgets are summed.
-ii) The model is trained to predict which nodes were merged at each scale.
-iii) In the expansion phase, merged nodes are expanded (shown in dark), inheriting their parent’s features, budget, and connectivity.
-The model is trained to (a) identify which edges should be removed (dotted lines), (b) predict how the parent’s budget should be split across the children (instead of predicting absolute values), and (c) refine the features of newly expanded nodes (refinement).
+ii) The model is then trained to predict which nodes were merged at each scale.
+iii) In the expansion phase, merged nodes (shown in dark in the leftmost column) are expanded back (copies shown in dark), inheriting their parent’s features, budget, and connectivity.
+In the refinement phase, the model is trained to (a) identify which edges should be removed (dotted lines), (b) predict how the parent’s budget should be split across the children, and (c) refine the features of newly expanded nodes.
+To address uneven growth across regions, we introduce a node budget mechanism: each cluster is assigned a budget indicating its remaining expansions.
+Budgets are recursively divided among child nodes, starting from a single super-node with the full budget and ending when all clusters have a budget of one.
+This provides more local control over the final node count and improves global consistency compared to prior methods [9, 10], which only append the desired size to node embeddings.
+At each expansion step, the model also predicts the mean feature of future child nodes conditioned on the parent’s feature, extending the scale-wise autoregressive idea of [23] so that predictions at one scale guide those at the next.
 3.2 Budgeted Coarsening
-During coarsening, we keep track of two quantities for every cluster we create.
-First, budgets count how many nodes the cluster already contains.
-Every node and hyperedge begins with a budget of $1$; when several of them merge, their budgets are summed and assigned to the new super-node or super-hyperedge.
-Secondly, features summarize the attributes of the merged vertices.
-We use a weighted average: the feature of a cluster is the budget-weighted mean of the features of its members.
+We adopt the coarsening strategy of [10], representing the hypergraph at multiple resolution levels by merging nodes into clusters.
+Node pairs to be merged are selected using the spectrum-preserving coarsening of [21], applied to the hypergraph’s clique expansion.
+Mergings in the clique expansion are then mirrored in the bipartite representation, with corresponding left side nodes being merged, and copies of the same hyperedge being subsequently collapsed.
+We limit node mergings to at most two per cluster, which consequently bounds hyperedge mergings to at most three [10].
+For each resulting cluster, we track two quantities.
+First, the budget is the total number of nodes contained in the cluster.
+Initially, each node and hyperedge has a budget of $1$; when clusters merge, their budgets are summed and assigned to the new super-cluster.
+The node budget provides a local growth constraint, indicating how many fine-level nodes each coarse cluster should expand into and guiding refinement in regions that grow at different rates.
+Second, the super-cluster’s feature is computed as the weighted mean of the features of the clusters it absorbs, so it represents the average feature of all its contained nodes.
+We denote node (left‑side) budgets and features by $\mathbf{b}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{F}_{L}$, and hyperedge (right‑side) budgets and features by $\mathbf{b}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ and $\mathbf{F}_{R}$.
+The following definition formalizes our coarsening process.
+Definition 1 (Bipartite graph coarsening).
+Let $H$ be an arbitrary hypergraph, $C=(\mathcal{V}_{c},\mathcal{E}_{c})$ its clique expansion, and $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E}_{b},\mathbf{b}_{L},\mathbf{b}_{R},\mathbf{F}_{L},\mathbf{F}_{R})$ its featured bipartite representation.
+Let $\mathcal{P}_{L}=\{\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(n)}\}$ be a partitioning [^2] That is, $\mathcal{V}^{(p)}\subseteq\mathcal{V}_{L}$, $\bigcup_{i=1}^{{n}}\mathcal{V}^{(i)}=\mathcal{V}_{L}$, and $\mathcal{V}^{(i)}\cap\mathcal{V}^{(j)}=\emptyset~\forall~1\leq i,j\leq n$}.
+of the node set $\mathcal{V}_{L}$ such that each set $\mathcal{V}^{(p)}$ induces a connected subgraph in $C$.
+We construct an intermediate coarsening $\tilde{B}(B,\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\mathcal{V}_{R},\bar{\mathcal{E}}_{b},\bar{\mathbf{b}}_{L},\mathbf{b}_{R},\bar{\mathbf{F}}_{L},\mathbf{F}_{R})$ by merging each part $\mathcal{V}^{(p)}$ into a single node $v^{(p)}\in\bar{\mathcal{V}}_{L}$, and by defining:
+$\bar{\mathbf{b}}_{L}[p]=\sum_{v\in\mathcal{V}^{(p)}}\mathbf{b}_{L}[v],\qquad\bar{\mathbf{F}}_{L}[p]=\frac{1}{\bar{\mathbf{b}}_{L}[p]}\sum_{v\in\mathcal{V}^{(p)}}\mathbf{b}_{L}[v]\mathbf{F}_{L}[v],$ (1)
+for every merged node $\mathcal{V}^{(p)}$.
+An edge $e_{\{p,q\}}\in\bar{\mathcal{E}}_{b}$ is added between $v^{(p)}\in\bar{\mathcal{V}}_{L}$ and $v^{(q)}\in\mathcal{V}_{R}$ if there exists an edge $e_{\{i,q\}}\in\mathcal{E}_{b}$ in the original bipartite representation between some $v^{(i)}\in\mathcal{V}^{(p)}$ and $v^{(q)}$.
+To complete the coarsening process, we define an equivalence relation $v_{1}\sim v_{2}\iff\mathcal{N}(v_{1})=\mathcal{N}(v_{2})$ on $\mathcal{V}_{R}$, where $\mathcal{N}(v)$ denotes the set of neighbors of $v$, i.e., we consider right side nodes having the same set of neighbors as equivalent, or in other words we consider hyperedges containing the same set of nodes as equivalent.
+This induces a partitioning $\mathcal{P}_{R}=\{\mathcal{V}_{R}^{(1)},\dots,\mathcal{V}_{R}^{(m)}\}$, allowing us to construct the fully coarsened bipartite representation $\bar{B}(\tilde{B},\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\bar{\mathcal{V}}_{R},\bar{\mathcal{E}}_{b},\bar{\mathbf{b}}_{L},\bar{\mathbf{b}}_{R},\bar{\mathbf{F}}_{L},\bar{\mathbf{F}}_{R})$ by merging each part $\mathcal{V}_{R}^{(p)}$ into a single node $v_{R}^{(p)}\in\bar{\mathcal{V}}_{R}$, similarly to the construction of $\bar{\mathcal{V}}_{L}$, and by defining:
+$\bar{\mathbf{b}}_{R}[p]=\sum_{v\in\mathcal{V}_{R}^{(p)}}\mathbf{b}_{R}[v],\qquad\bar{\mathbf{F}}_{R}[p]=\frac{1}{\bar{\mathbf{b}}_{R}[p]}\sum_{v\in\mathcal{V}_{R}^{(p)}}\mathbf{b}_{R}[v]\mathbf{F}_{R}[v],$ (2)
+for every merged hyperedge $\mathcal{V}_{R}^{(p)}$.
+Remark 2.
+Informally, we cluster nodes in the clique expansion, merge the corresponding left-side nodes of the bipartite graph, and compute their budgets and features as described above.
+Right-side nodes (hyperedges) connected to the same left-side nodes are then merged.
+In our implementation, we select nodes for clustering using spectrum-preserving coarsening [21].
+In the final reduction—when only one node and one hyperedge remain—we replace their features with zero matrices to yield a feature-agnostic initialization for generation.
+Please note that cluster size vectors introduced in [9] correspond to the sizes of all sets $\mathcal{V}_{L}^{(p)}$ and $\mathcal{V}_{R}^{(p)}$, i.e., the number of mergings for each cluster at this specific step, and not the number of nodes absorbed since the initialization of coarsening—which we call budget.
+Further details about the coarsening methodology are provided in Appendix B.
+Some visual examples of coarsening sequences are provided in Appendix C.
+We provide the following theoretical result regarding the feature merging process in FAHNES.
 Proposition 3.
-Setting cluster features as the average of the nodes they contain minimizes the mean squared error (MSE) between the features of the fully expanded hypergraph and those of the original hypergraph.
-In what follows we denote left-side (node) budgets and features by $\mathbf{b}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{F}_{L}$, and right-side (hyperedge) budgets and features by $\mathbf{b}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ and $\mathbf{F}_{R}$.
-Definition 2 (Bipartite graph coarsening).
-Let $H$ be an arbitrary hypergraph, $C=(\mathcal{V}_{c},\mathcal{E}_{c})$ its unfeatured weighted clique expansion, and $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{b}_{L},\mathbf{b}_{R},\mathbf{F}_{L},\mathbf{F}_{R})$ its featured bipartite representation.
-Given a partitioning $\mathcal{P}_{L}=\{\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(n)}\}$ of the node set $\mathcal{V}_{L}$ such that each part $\mathcal{V}^{(p)}$ induces a connected subgraph in $C$, we construct an intermediate coarsening $\tilde{B}(B,\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\mathcal{V}_{R},\bar{\mathcal{E}},\bar{\mathbf{b}},\bar{\mathbf{F}}_{L},\mathbf{F}_{R})$ by merging each part $\mathcal{V}^{(p)}$ into a single node $v^{(p)}\in\bar{\mathcal{V}}_{L}$, and by defining:
-$\left(\bar{\mathbf{b}}_{L}\right)_{v^{(p)}}=\sum_{v\in\mathcal{V}^{(p)}}\left(\mathbf{b}_{L}\right)_{v},\qquad\left(\bar{\mathbf{F}}_{L}\right)_{v^{(p)}}=\frac{1}{\left(\bar{\mathbf{b}}_{L}\right)_{v^{(p)}}}\sum_{v\in\mathcal{V}^{(p)}}\left(\mathbf{b}_{L}\right)_{v}\left(\mathbf{F}_{L}\right)_{v},$ (1)
-for every supernode $\mathcal{V}^{(p)}$.
-An edge $e_{\{p,q\}}\in\bar{\mathcal{E}}$ is added between $v^{(p)}\in\bar{\mathcal{V}}_{L}$ and $v^{(q)}\in\mathcal{V}_{R}$ if there exists an edge $e_{\{i,q\}}\in\mathcal{E}$ in the original bipartite representation between some $v^{(i)}\in\mathcal{V}^{(p)}$ and $v^{(q)}$.
-To complete the coarsening process, we define an equivalence relation $v_{1}\sim v_{2}\iff\mathcal{N}(v_{1})=\mathcal{N}(v_{2})$ on $\mathcal{V}_{R}$, where $\mathcal{N}(v)$ denotes the set of neighbors of $v$.
-This induces a partitioning $\mathcal{P}_{R}=\{\mathcal{V}_{R}^{(1)},\dots,\mathcal{V}_{R}^{(m)}\}$, allowing us to construct the fully coarsened bipartite representation $\bar{B}(\tilde{B},\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\bar{\mathcal{V}}_{R},\bar{\mathcal{E}},\bar{\mathbf{b}},\bar{\mathbf{F}}_{L},\bar{\mathbf{F}}_{R})$ by merging each part $\mathcal{V}_{R}^{(p)}$ into a single node $v_{R}^{(p)}\in\bar{\mathcal{V}}_{R}$, similarly to the construction of $\bar{\mathcal{V}}_{L}$, and by defining:
-$\left(\bar{\mathbf{b}}_{R}\right)_{v^{(p)}}=\sum_{v\in\mathcal{V}_{R}^{(p)}}\left(\mathbf{b}_{R}\right)_{v},\qquad\left(\bar{\mathbf{F}}_{R}\right)_{v^{(p)}}=\frac{1}{\left(\bar{\mathbf{b}}_{R}\right)_{v^{(p)}}}\sum_{v\in\mathcal{V}_{R}^{(p)}}\left(\mathbf{b}_{R}\right)_{v}\left(\mathbf{F}_{R}\right)_{v},$ (2)
-for every superhyperedge $\mathcal{V}_{R}^{(p)}$.
-Remark 3.
-Informally, we first pick the node clusters on the clique expansion, merge those nodes on the left side of the bipartite graph, and compute their budgets and features as above.
-We then merge right-side nodes that appear multiple times, i.e., they connect the same left-side nodes.
-In practice, for the final reduction step, when only one node and one hyperedge remain, we also replace their features by matrices full of zeros.
-This design choice enables a fully agnostic initial step during generation.
+Setting cluster features as the mean of the nodes they contain minimizes the mean squared error between the features of the fully expanded hypergraph and those of the original hypergraph.
 3.3 Budgeted Expansion and Refinement
-During expansion, since our framework only conditions on the total number of nodes in the final hypergraph, we maintain a single node budget vector $\mathbf{b}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ for the left-side (node) partition and discard the right-side (hyperedge) budget.
-Each cluster is split into multiple child nodes, which initially inherit both the parent’s budget and feature vector.
-A subsequent refinement step then determines: i) how the parent’s budget should be divided among its children, and ii) how to update the children’s feature vectors.
-More specifically, the expansion step duplicates vertices and hyperedges, with child nodes inheriting all of their parent’s connections.
-The refinement step then i) removes edges that should not persist at the finer resolution, ii) redistributes the integer budget among the child nodes according to a probability-like split vector, and iii) replaces the coarse features with new refined predictions.
+Once multi-scale representations for the dataset of hypergraphs are produced by the coarsening procedure, the objective is to train a model able to reverse it.
+In this section, we define the inverse of the coarsening procedure, which the model will learn to imitate.
+It comprises two stages: expansion (“upsampling” by duplicating nodes and hyperedges) and refinement (adjusting connectivity, budgets, and features).
+This process exclusively works on the bipartite representation.
+Since our framework only conditions on the total number of nodes in the final hypergraph, we maintain a single node budget vector $\mathbf{b}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ for the node (left-side) partition and discard the hyperedge (right-side) budget.
+To undo mergings, clusters are recursively split into multiple child nodes.
+Each child initially copies the parent’s connections and inherits its feature and budget, before a subsequent refinement step, then: i) redistributes the budget among the children, ii) updates their features, and iii) removes edges that should not persist at the finer scale.
 Formally, the expansion and refinement steps can be described as follows.
 Definition 4 (Bipartite graph expansion).
-Given a bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{b},\mathbf{F}_{L},\mathbf{F}_{R})$ and two cluster size vectors $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$, $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$, denoting the expansion size of each node, let $\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b}^{\text{expanded}},\mathbf{F}_{L}^{\text{expanded}},\mathbf{F}_{R}^{\text{expanded}})$ denote the expansion of $B$, whose node sets, budgets, and features are given by:
-$\begin{split}&\tilde{\mathcal{V}}_{L}=\mathcal{V}_{L}^{(1)}\cup\cdots\cup\mathcal{V}_{L}^{(|\mathcal{V}_{L}|)},\text{~{}where~{}~{}}\mathcal{V}_{L}^{(p)}=\{v_{L}^{(p,i)}\mid 1\leq i\leq\mathbf{v}_{L}[p]\}\text{~{}~{}for~{}~{}}1\leq p\leq|\mathcal{V}_{L}|,\\ &\tilde{\mathcal{V}}_{R}=\mathcal{V}_{R}^{(1)}\cup\cdots\cup\mathcal{V}_{R}^{(|\mathcal{V}_{R}|)},\text{~{}where~{}~{}}\mathcal{V}_{R}^{(p)}=\{v_{R}^{(p,i)}\mid 1\leq i\leq\mathbf{v}_{R}[p]\}\text{~{}~{}for~{}~{}}1\leq p\leq|\mathcal{V}_{R}|,\\ &\left(\mathbf{b}^{\text{expanded}}\right)_{v_{L}^{(p,i)}}=\mathbf{b}_{\mathcal{V}_{L}^{(p)}}\text{~{}~{}for~{}~{}}1\leq i\leq\mathbf{v}_{L}[p],1\leq p\leq|\mathcal{V}_{L}|,\\ &\left(\mathbf{F}_{L}^{\text{expanded}}\right)_{v_{L}^{(p,i)}}=\left(\mathbf{F}_{L}\right)_{\mathcal{V}_{L}^{(p)}}\text{~{}~{}for~{}~{}}1\leq i\leq\mathbf{v}_{L}[p],1\leq p\leq|\mathcal{V}_{L}|,\\ &\left(\mathbf{F}_{R}^{\text{expanded}}\right)_{v_{R}^{(p,i)}}=\left(\mathbf{F}_{R}\right)_{\mathcal{V}_{R}^{(p)}}\text{~{}~{}for~{}~{}}1\leq i\leq\mathbf{v}_{R}[p],1\leq p\leq|\mathcal{V}_{R}|.\end{split}$ (3)
+Given a bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{b},\mathbf{F}_{L},\mathbf{F}_{R})$ and two expansion size vectors $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$, $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$, denoting the number of duplication for nodes (left side) and hyperedges (right side), respectively.
+Let $\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b}^{\text{expanded}},\mathbf{F}_{L}^{\text{expanded}},\mathbf{F}_{R}^{\text{expanded}})$ denote the expansion of $B$, whose node sets, budgets, and features are given by:
+[…]
 The edge set $\tilde{\mathcal{E}}$ includes all the cluster interconnecting edges: $\{e_{\{p,i;q,j\}}\mid e_{\{p,q\}}\in\mathcal{E},v_{L}^{(p,i)}\in\mathcal{V}_{L}^{(p)},v_{R}^{(q,j)}\in\mathcal{V}_{R}^{(q)}\}$.
 Remark 5.
-Expansion thus acts as a clone-and-rewire operation: vertices and hyperedges are duplicated, and each child inherits every incident edge of its parent.
+Expansion thus acts as a clone‑and‑rewire operation: vertices and hyperedges are duplicated, and each child inherits every incident edge of its parent.
 Definition 6 (Bipartite graph refinement).
-Given a bipartite graph $\tilde{B}=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b},\mathbf{F}_{L},\mathbf{F}_{R})$, an edge selection vector $\mathbf{e}\in\{0,1\}^{|\mathcal{E}|}$, a budget split vector $\mathbf{f}\in[0,1]^{|\tilde{\mathcal{V}}_{L}|}$, satisfying $\sum_{v\in\mathcal{V}_{L}^{(p)}}\mathbf{f}_{v}=1$ for all cluster $\mathcal{V}_{L}^{(p)}$ in $\mathcal{V}_{L}$, and two feature refinement vectors $\mathbf{F}_{L}^{\text{refine}}$ and $\mathbf{F}_{R}^{\text{refine}}$ with the same dimensions as $\mathbf{F}_{L}$ and $\mathbf{F}_{R}$, let $B(\tilde{B},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\mathcal{E},\lceil\mathbf{b}\odot\mathbf{f}\rfloor,\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$ denote the refinement of $\tilde{B}$, where $\mathcal{E}\subseteq\tilde{\mathcal{E}}$ such that the $i$-th edge $e_{(i)}\in\mathcal{E}$ if and only if $\mathbf{e}[i]=1$.
+Given a bipartite graph $\tilde{B}=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\tilde{\mathcal{E}},\mathbf{b},\mathbf{F}_{L}^{\text{expanded}},\mathbf{F}_{R}^{\text{expanded}})$, an edge selection vector $\mathbf{e}\in\{0,1\}^{|\tilde{\mathcal{E}}|}$, a budget split vector $\mathbf{f}\in[0,1]^{|\tilde{\mathcal{V}}_{L}|}$, satisfying $\sum_{v\in\mathcal{V}_{L}^{(p)}}\mathbf{f}[v]=1$ for all cluster $\mathcal{V}_{L}^{(p)}$ in $\tilde{\mathcal{V}}_{L}$, and two feature refinement vectors $\mathbf{F}_{L}^{\text{refine}}$ and $\mathbf{F}_{R}^{\text{refine}}$ with the same dimensions as $\mathbf{F}_{L}^{\text{expanded}}$ and $\mathbf{F}_{R}^{\text{expanded}}$, let $B(\tilde{B},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})=(\tilde{\mathcal{V}}_{L},\tilde{\mathcal{V}}_{R},\mathcal{E},\lceil\mathbf{b}\odot\mathbf{f}\rfloor,\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$ denote the refinement of $\tilde{B}$, where $\mathcal{E}\subseteq\tilde{\mathcal{E}}$ such that the $i$-th edge $e_{(i)}\in\mathcal{E}$ if and only if $\mathbf{e}[i]=1$.
 Remark 7.
@@
 In the case of a tie (e.g., when an odd number must be split evenly), the child with the lowest index receives the larger share, and the remaining budget is distributed among the others accordingly.
-The generation of a hypergraph with $N$ nodes proceeds in three main stages:
-  1. 1.
-    Initialization.
-Start from a minimal bipartite graph: $B^{(L)}=(\{1\},\{2\},\{(1,2)\})$, consisting of a single node on each side connected by one edge.
-The left-side node is assigned the full node budget, i.e., $\mathbf{b}=(N)$.
-If node and hyperedge features need to be generated, $\mathbf{F}_{L}$ and $\mathbf{F}_{R}$ are initialized as zero matrices.
-  2. 2.
-[…]
-  3. 3.
-    Hypergraph reconstruction. Once the final bipartite graph is generated, construct the hypergraph by collapsing each right-side node into a hyperedge connecting its adjacent left-side nodes.
 3.4 Probabilistic Modeling
@@
 hypergraph instances.
-Our aim is to approximate the unknown generative process by learning a distribution $p(H)$.
-We model the marginal likelihood of each hypergraph $H$ as a sum over the likelihoods of its bipartite representation’s expansion sequences:
-$p(H)=p(B)=\sum_{\varpi\in\Pi(B)}p(\varpi),$ (4)
-where $\Pi(B)$ denotes the set of valid expansion sequences from a minimal bipartite graph to the full bipartite representation $B$ corresponding to $H$.
-Each intermediate $B^{(l-1)}$ is generated by refining its predecessor, in accordance with Definitions 4 and 6.
+Our objective is to approximate the unknown generative process by learning a distribution $p(H)$.
+We model the marginal likelihood of each hypergraph $H$ as a sum over the likelihoods of its bipartite representation’s expansion sequences $p(H)=p(B)=\sum_{\varpi\in\Pi(B)}p(\varpi)$, where $\Pi(B)$ denotes the set of valid expansion sequences from a minimal bipartite graph to the full bipartite representation $B$ corresponding to $H$.
+Each intermediate $B^{(l-1)}$ is generated by expanding and refining its predecessor, in accordance with Definitions 4 and 6.
+To simplify notations, we drop the superscript for $\mathbf{F}^{\text{refine}}$ and simply write $\mathbf{F}$.
 Assuming a Markovian generative structure, the likelihood of a specific expansion sequence $\varpi$ is factorized as:
-$p(\varpi)=\underbrace{p(B^{(L)})}_{1}\prod_{l=L}^{1}p(B^{(l-1)}|B^{(l)})=\prod_{l=L}^{1}p(\mathbf{e}^{(l-1)}|\tilde{B}^{(l-1)})p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)}).$ (5)
-To simplify the modeling process and avoid learning two separate distributions $p(\mathbf{e}^{(l)}|\tilde{B}^{(l)})$ and $p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)})$, we rearrange terms as follows:
-$p(\varpi)=p(\mathbf{e}^{0}|\tilde{B}^{0})p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)})\left[\prod_{l=L-1}^{1}p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)}|\tilde{B}^{(l)})\right],$ (6)
-where $p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)},\mathbf{b}^{(L)},\mathbf{F}_{L}^{(L)},\mathbf{F}_{R}^{(L)})=p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)},\mathbf{b}^{(L)},\mathbf{F}_{L}^{(L)},\mathbf{F}_{R}^{(L)}|B^{(L)})$.
-We assume that the variables $\mathbf{v}_{L}^{(l)}$, $\mathbf{v}_{R}^{(l)}$, $\mathbf{b}^{(l)}$, $\mathbf{F}_{L}^{(l)}$, and $\mathbf{F}_{R}^{(l)}$ are conditionally independent of $\tilde{B}^{(l)}$ when conditioned on $B^{(l)}$:
-$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)},\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)}).$ (7)
+$p(\varpi)=\overbrace{p(B^{(L)})}^{1}\prod_{l=L}^{1}p(B^{(l-1)}|B^{(l)})=\prod_{l=L}^{1}p(\mathbf{e}^{(l-1)},\mathbf{f}^{(l-1)},\mathbf{F}_{L}^{(l-1)},\mathbf{F}_{R}^{(l-1)}|\tilde{B}^{(l-1)})p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)}).$
+To simplify the modeling process and avoid learning two separate distributions $p(\mathbf{e}^{(l)},\mathbf{f}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|\tilde{B}^{(l)})$ and $p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)})$, we rearrange terms as follows:
+$p(\varpi)=p(\mathbf{e}^{(0)},\mathbf{f}^{(0)},\mathbf{F}_{L}^{(0)},\mathbf{F}_{R}^{(0)}|\tilde{B}^{0})p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)})\left[\prod_{l=L-1}^{1}p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)},\mathbf{f}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|\tilde{B}^{(l)})\right],$
+where $p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)})=p(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)}|B^{(L)})$.
+We assume that the variables $\mathbf{v}_{L}^{(l)}$ and $\mathbf{v}_{R}^{(l)}$ are conditionally independent of $\tilde{B}^{(l)}$ when conditioned on $B^{(l)}$:
+$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)},\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)}).$ (6)
 This allows us to write the combined likelihood as:
-$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)}|\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{e}^{(l)},\mathbf{b}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|B^{(l)}).$ (8)
-3.5 Flow-matching
-We use a flow-matching framework [21] with an endpoint parameterization similar to [22], or alternatively a denoising diffusion model [23] with linear interpolation between the prior and target distribution.
-The objective is to align two probability distributions $p_{0}$ and $p_{1}$ by evolving the points from the initial distribution $p_{0}$ to the target distribution $p_{1}$.
-This is achieved by defining a flow through a time-dependent vector field $\mathbf{f}(\mathbf{x},t)$, governed by the following ordinary differential equation (ODE) $\frac{\partial\mathbf{x}}{\partial t}=\mathbf{f}(\mathbf{x},t),\mathbf{x}_{0}\sim p_{0}$, where $\mathbf{f}(\mathbf{x},t)$ is learned such that, after integrating the ODE, samples from $p_{1}$ can be generated.
-With endpoint parametrization, the model learns to predict the endpoint $\mathbf{x}_{1}$ of the trajectory.
-The flow $\mathbf{f}(\mathbf{x},t)$ can then be recovered from the predicted endpoint $\hat{\mathbf{x}}_{1}(\mathbf{x},1)$ using the relation:
-$\mathbf{f}(\mathbf{x}_{t},t)=\frac{\hat{\mathbf{x}}_{1}(\mathbf{x},1)-\mathbf{x}_{t}}{1-t}.$ (9)
-To train the model, the loss function minimizes the squared difference between the predicted endpoint $\hat{\mathbf{x}}_{1}(x_{t},t)$ and the actual endpoint $\mathbf{x}_{1}$:
-$\mathcal{L}=\mathbb{E}_{t\sim\mathcal{U}(0,1),\>\mathbf{x}_{0}\sim p_{0},\>\mathbf{x}_{1}\sim p_{1}}\left[\left(\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{1}\right)^{2}\right],\text{~{}~{}where~{}~{}}\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0}.$ (10)
-For the expansion process and edge selection, we set $p_{0}$ as the Gaussian distribution.
-For budget splits, as their support is constrained to the simplex, we use flow matching similarly to [22], where $p_{0}$ is a Dirichlet distribution across all clusters.
-The model predictions are projected on the simplex using a Von Neumann projection [24].
-Features are generated using the process introduced in [20], i.e., $p_{0}$ is the Gaussian distribution, and the feature of the parent cluster is given as conditioning during generation using a FiLM layer [25].
-More details can be found in the supplementary material.
+$p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)}|B^{(l)})p(\mathbf{e}^{(l)},\mathbf{f}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|\tilde{B}^{(l)})=p(\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)},\mathbf{e}^{(l)},\mathbf{f}^{(l)},\mathbf{F}_{L}^{(l)},\mathbf{F}_{R}^{(l)}|\tilde{B}^{(l)}),$ (7)
+which corresponds to the combined expansion and refinement step that inverts coarsening at depth $l$.
+During training, the model approximates the right-hand side of (7), learning to generate, for each expansion and refinement step, $\mathbf{v}_{L}$, $\mathbf{v}_{R}$, $\mathbf{e}$, $\mathbf{f}$, and $\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}}$.
+Once trained, the model generates a hypergraph with $N$ nodes through successive expansion and refinement steps:
+  1. 1.
+    Initialization.
+Start from a minimal bipartite graph: $B^{(L)}=(\{1\},\{2\},\{(1,2)\},(N))$, consisting of a single node on each side connected by one edge.
+The left-side node is assigned the full node budget, i.e., $\mathbf{b}=[N]$.
+If node and hyperedge features need to be generated, $\mathbf{F}_{L}$ and $\mathbf{F}_{R}$ are initialized as zero matrices.
+  2. 2.
+    Expansion and refinement.
+Iteratively expand and refine the current bipartite representation to add details until the desired size is attained: $B^{(l)}\xrightarrow{\text{expand}}\tilde{B}^{(l-1)}\xrightarrow{\text{refine}}B^{(l-1)}$.
+  3. 3.
+    Hypergraph reconstruction. Once the final bipartite graph is generated, construct the hypergraph by collapsing each right-side node into a hyperedge connecting its adjacent left-side nodes.
 3.5 Minibatch OT-Coupling
-Minibatch OT-coupling [12, 13] has shown promising results in image generation by encouraging smoother and more linear generative trajectories, which can improve sampling efficiency and model convergence.
-Instead of sampling independently from the prior and target distributions, OT-coupling aligns samples within each minibatch via an optimal transport plan.
-Specifically, samples are first drawn from the prior (typically Gaussian), then their corresponding target samples are drawn, and finally, the prior samples are optimally matched (reindexed) to the targets to minimize transport cost.
-We extend this idea to our setting by applying OT-coupling to the children of an expanded node, treating these children as a minibatch.
-For each expansion, we solve a local OT problem to align the initial noised samples of children with their corresponding targets.
-A crucial constraint is to avoid swapping samples between non-equivalent nodes—i.e., only nodes that are structurally and feature-wise indistinguishable up to noise can be swapped.
-Swapping non-equivalent nodes would alter the learned data distribution and degrade model fidelity.
-To keep computational costs manageable, we restrict OT-coupling to groups of children within a single expanded cluster, satisfying the equivalence condition.
-Since our expansion step produces two children per cluster, OT-coupling simplifies to evaluating just two permutations per cluster, making it extremely lightweight and easily parallelizable using standard tensor operations.
-We summarize our theoretical guarantee as follows:
+Graphs and hypergraphs lack a natural node ordering, making it challenging to align predictions with their targets.
+After an expansion, the model may produce a permutation of the correct predictions; yet, without alignment, the computed loss will be large despite the prediction being correct.
+This mismatch introduces significant noise into the learning signal.
+To address this, we generalize minibatch OT-coupling [12, 13] from image generation to our setting.
+In its original form, OT-coupling aligns samples from the prior and target distributions within a minibatch via an optimal transport plan, reindexing prior samples to minimize the matching cost.
+Formally, given prior samples $\mathbf{X}\in\mathbb{R}^{B\times d}$ and targets $\mathbf{Y}\in\mathbb{R}^{B\times d}$, OT-coupling finds:
+$\mathbf{P}^{\ast}=\operatorname*{\operatorname{arg\,min}}_{\mathbf{P}\in\Pi_{B}}\|\mathbf{X}\mathbf{P}-\mathbf{Y}\|^{2},$ (8)
+where $\mathbf{P}$ is a permutation matrix and $\Pi_{B}$ denotes the set of all $B\times B$ permutations.
+In our context, for a given bipartite graph with $n$ nodes, adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$, expanded budgets $\mathbf{b}\in\mathbb{N}^{n}$ and expanded features $\mathbf{F}\in\mathbb{R}^{n\times d}$ where $d$ is the dimension of node or hyperedge features, $\mathbf{X}$ and $\mathbf{Y}$ correspond to prior samples and targets, respectively, and $\mathbf{P}$ is implemented as a node-swapping operation.
+To avoid bias, swaps are only allowed between equivalent nodes—those structurally and feature-wise indistinguishable up to noise.
+Swapping non-equivalent nodes would alter the learned distribution and degrade model performance.
+This leads to the following problem:
+$\mathbf{P}^{\ast}=\operatorname*{\operatorname{arg\,min}}_{\mathbf{P}\in\Pi_{n}}\|\mathbf{P}\mathbf{X}-\mathbf{Y}\|^{2}\quad\text{s.t.}\quad\mathbf{P}^{\top}\mathbf{A}\mathbf{P}=\mathbf{A},\;\mathbf{P}\mathbf{b}=\mathbf{b},\;\mathbf{P}\mathbf{F}=\mathbf{F}.$ (9)
+To reduce overhead, we only consider permuting children from a single cluster expansion, where equivalence holds naturally.
+In our implementation, with exactly two or three children per cluster, only two or six permutations are possible, making the operation lightweight and easily parallelizable with standard tensor operations.
 Proposition 8 (Informal).
-Minibatch OT-coupling in our framework preserves the target distribution while improving the straightness and consistency of the learned generative flow.
+Minibatch OT-coupling in our framework preserves the target distribution and produces shorter target paths.
+3.6 Generalization to Graphs
+Our framework can be directly extended to generate standard graphs by adapting the coarsen–expand procedure of [9].
+As in the hypergraph setting, each node begins with a budget of one, which is summed when nodes are merged, while features are aggregated through a weighted average with weights proportional to node budgets.
+During expansion, clusters propagate their budgets and features to child nodes, and the model predicts how to split the budget among them while refining their features.
+Graph generation, therefore, starts from a single node with a budget equal to the desired size and proceeds through successive expansion and refinement steps.
+To ensure stable training and alignment between predicted and target structures, minibatch OT-coupling is applied to groups of expanded nodes exactly as in the hypergraph case.
 4 Experiments and Results
-In this section, we detail our experimental setup, covering datasets, evaluation metrics, results, and ablation studies.
-The datasets are divided into featureless and featured hypergraphs.
-For the featureless hypergraphs, we compare FAHNES against HyperPA [26], a Variational Autoencoder (VAE) [27], a Generative Adversarial Network (GAN) [28], and a standard 2D diffusion model [23] trained on incidence matrix images, where hyperedge membership is represented by white pixels and absence by black pixels.
-For the featured hypergraphs, we compare FAHNES against the graph-based models DiGress [7], DisCo [9], Cometh [29], and DeFoG [30] in a molecular dataset.
-For 3D meshes, we also compare FAHNES with a sequential disjoint generation baseline (see Figure 1), where we first generate the hypergraph topology and then the features.
-Extended numerical results and several visualizations of our generated hypergraphs are available in the supplementary material.
-The graph-based approaches DiGress, DisCo, Cometh, and DeFoG are only tested on a simple graph dataset because they cannot generate hypergraphs.
-Although one might attempt to generate hypergraphs using these existing graph generators over the star or clique expansions, this strategy is impractical for several reasons.
-First, generating the clique expansion of a hypergraph and then recovering the original hyperedges is an NP-hard problem, as it requires enumerating all cliques.
-Second, in the bipartite representation of a hypergraph, it is non-trivial to determine which partition represents nodes and which represents hyperedges—a distinction that is critical for correct reconstruction.
-Empirically, we also find that graph-based generative models often fail to produce valid bipartite structures.
-For instance, we observed that only 30% of outputs from both models [11, 7] generated valid bipartite graphs.
+In this section, we detail our experimental setup, covering datasets, evaluation metrics, implementation choices, baselines, results, and ablation studies.
+Full experimental details are provided in Appendix D.
+Visualizations of the generated samples are provided in Appendix I.
+4.1 Datasets and Experimental Setup
 Datasets.
 We evaluate our method on five featureless datasets: Stochastic Block Model (SBM) [24], Ego [25], Tree [26], ModelNet40 bookshelf, and ModelNet40 piano [27].
-We also evaluate FAHNES on three featured hypergraphs datasets, including two sets of 3D meshes and one molecular dataset: Manifold40 bench, Manifold40 airplane [^1] The full set of numerical results for the Manifold40 datasets is provided in the supplementary material.
-[35], and QM9 [36, 37] with implicit hydrogens.
-In the case of bench and airplane, node features are 3D positions, and hyperedges do not have features.
-For QM9, both nodes and hyperedges have features, which are one-hot vectors coding for atom types and bond types, respectively.
-Although molecules can be represented as traditional graphs, modeling them with hypergraphs opens the door to representing higher-order interactions between atoms.
-However, in this work, we treat molecules strictly as graphs, with hyperedges corresponding to simple pairwise bonds.
-Future work could group functional groups’ atoms together in hyperedges.
-Further details of the datasets are provided in the supplementary material.
+We also evaluate FAHNES on two featured hypergraphs datasets, comprising two sets of 3D meshes: Manifold40 bench and Manifold40 airplane [28].
+Additionally, we adapt our method to graphs using the methodology outlined in Section 3.6 and evaluate it on two point cloud datasets sampled on the meshes of the two previous mesh datasets.
+For 3D mesh and point cloud datasets, node features are 3D positions, and edges or hyperedges do not have features.
+Table 1: Comparison between FAHNES and other baselines for the SBM, Ego, and Tree hypergraphs.
+Method SBM Hypergraphs Ego Hypergraphs Tree Hypergraphs
+Valid SBM $\uparrow$ Node Num $\downarrow$ Spectral $\downarrow$ Valid Ego $\uparrow$ Node Num $\downarrow$ Spectral $\downarrow$ Valid Tree $\uparrow$ Node Num $\downarrow$ Spectral $\downarrow$
+HyperPA $2.5$ $\underline{0.075}$ $0.273$ $0.0$ $35.830$ $0.237$ $0.0$ $2.350$ $0.159$
+VAE $0.0$ $0.375$ $0.024$ $0.0$ $47.580$ $0.133$ $0.0$ $9.700$ $0.124$
+GAN $0.0$ $1.200$ $0.059$ $0.0$ $60.350$ $0.230$ $0.0$ $6.000$ $0.089$
+Diffusion $0.0$ $0.150$ $0.031$ $0.0$ $\underline{4.475}$ $0.190$ $0.0$ $2.225$ $0.127$
+HYGENE $\underline{65.0}$ $0.525$ $\underline{0.010}$ $\underline{90.0}$ $12.550$ $\mathbf{0.004}$ $\underline{77.5}$ $\mathbf{0.000}$ $\underline{0.012}$
+FAHNES $\mathbf{87.8\vbox{\hbox{$\bm{\pm 3.1}$}}}$ $\mathbf{0.029\vbox{\hbox{$\bm{\pm 0.009}$}}}$ $\mathbf{0.006\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $\mathbf{99.5\vbox{\hbox{$\bm{\pm 1.1}$}}}$ $\mathbf{0.128\vbox{\hbox{$\bm{\pm 0.171}$}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.003}$}}}$ $\mathbf{89.7\vbox{\hbox{$\bm{\pm 6.0}$}}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{0.003\vbox{\hbox{$\bm{\pm 0.002}$}}}$
+Table 2: Evaluation on ModelNet40.
+Method ModelNet40 Bookshelf ModelNet40 Piano
+Node Num $\downarrow$ Spectral $\downarrow$ Node Num $\downarrow$ Spectral $\downarrow$
+HyperPA $8.025$ $\underline{0.048}$ $0.825$ $\underline{0.067}$
+VAE $47.450$ $0.190$ $75.350$ $0.396$
+GAN $\mathbf{0.000}$ $0.476$ $\mathbf{0.000}$ $0.697$
+Diffusion $\mathbf{0.000}$ $0.079$ $\underline{0.050}$ $0.113$
+HYGENE $69.730$ $0.068$ $42.520$ $0.117$
+[…]
+Table 3: Evaluation on the graph point cloud datasets.
+Method Airplane Point Clouds Bench Point Clouds
+Chamfer Dist $\downarrow$ Spectral $\downarrow$ Ratio $\downarrow$ Chamfer Dist $\downarrow$ Spectral $\downarrow$ Ratio $\downarrow$
+DiGress OOM OOM OOM OOM OOM OOM
+DeFoG OOM OOM OOM OOM OOM OOM
+FAHNES $\mathbf{0.094\vbox{\hbox{$\bm{\pm 0.006}$}}}$ $\mathbf{0.005\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $\mathbf{67.311\vbox{\hbox{$\bm{\pm 44.903}$}}}$ $\mathbf{0.130\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{84.423\vbox{\hbox{$\bm{\pm 0.000}$}}}$
 Metrics.
 For featureless hypergraphs, we follow the evaluation criteria in [10].
-These include: i) structural comparison metrics such as Node Num (difference in node counts), Node Deg (Wasserstein distance between node degree distributions), and Edge Size (Wasserstein distance between hyperedge size distributions); ii) topological analysis with Spectral (maximum mean discrepancy between the spectral distributions).
-In scenarios where datasets enforce structural constraints, we report Valid—the proportion of generated samples satisfying those constraints.
+These include: i) structural comparison metrics such as Node Num (difference in node counts); ii) topological analysis with Spectral (maximum mean discrepancy between the spectral distributions).
+In scenarios where datasets enforce structural constraints, we report Valid—the percentage of generated samples satisfying those constraints.
 For all metrics except Valid, lower values indicate improved performance, while higher values are preferable for Valid.
-In the case of 3D meshes, we use the nearest training sample Chamfer distance, which computes the Chamfer distance between point clouds sampled from a generated sample and all training samples and outputs the minimum distance.
-For QM9, we use the same metrics as in [30]: i) overall validity of molecules with Valid (fraction of molecules that can be sanitized using RDKit [38]); ii) diversity of samples with Unique (fraction of unique SMILES generated); and iii) chemical similarity with the learned distribution with FCD (Fréchet inception distance between the activations of ChemNet on the train samples and generated samples [39]).
+In the case of 3D meshes, we use ChamDist (the nearest training sample Chamfer distance), which computes the Chamfer distance between point clouds sampled from a generated sample and all training samples and outputs the minimum distance.
+For graph point cloud datasets, we report the nearest training sample Chamfer distance, the maximum mean discrepancy (MMD) between the spectral distributions of the generated samples and training set, and the ratio of various metrics computed for the generated samples and test set (see Appendix D for details).
+Detailed numerical results with additional metrics can be found in Appendix H.
 Implementation details.
-We adopt the Local PPGN architecture [11, 6].
-During generation, we use inpainting to enforce constraints: i) budget splits are fixed to 1 for unexpanded or terminal clusters, ii) equal splits are enforced for expanded clusters with a budget of 2, iii) clusters of size one cannot expand, and iv) features for non-expanded clusters are copied.
-Table 1: Comparison between FAHNES and other baselines for the SBM, Ego, and Tree hypergraphs.
-Model SBM Hypergraphs Ego Hypergraphs Tree Hypergraphs
-Valid SBM $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Ego $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Tree $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
-HyperPA [26] 2.5% 0.075 4.062 0.407 0.273 0% 35.83 2.590 0.423 0.237 0% 2.350 0.315 0.284 0.159
-VAE [27] 0% 0.375 1.280 1.059 0.024 0% 47.58 0.803 1.458 0.133 0% 9.700 0.072 0.480 0.124
-GAN [28] 0% 1.200 2.106 1.203 0.059 0% 60.35 0.917 1.665 0.230 0% 6.000 0.151 0.459 0.089
-Diffusion [23] 0% 0.150 1.717 1.390 0.031 0% 4.475 3.984 2.985 0.190 0% 2.225 1.718 1.922 0.127
-HYGENE [6] 65% 0.525 0.321 0.002 0.010 90% 12.55 0.063 0.220 0.004 77.5% 0.000 0.059 0.108 0.012
+Our implementation uses a custom flow-matching framework [22] together with a local PPGN architecture [9].
+We also leverage our budget component using graph inpainting techniques, enabling the model to learn only the necessary parts of the distribution: i) budget splits are fixed to 1 for unexpanded clusters or those with a budget of $1$, ii) equal splits are enforced for expanded clusters with a budget of $2$, iii) clusters of size one are not allowed to expand, and iv) the features of non-expanded clusters are copied unchanged.
+Extensive implementation details are provided in Appendix E.
+The training and sampling procedures are explained in Appendix F.
+We also analyze the computational complexity in Appendix G.
+Baselines.
+For the featureless hypergraphs, we compare FAHNES against HYGENE [10], HyperPA [29], a Variational Autoencoder (VAE) [30], a Generative Adversarial Network (GAN) [31], and a standard 2D diffusion model [32] trained on incidence matrix images, where hyperedge membership is represented by white pixels and absence by black pixels.
+For 3D meshes, we compare FAHNES with a sequential disjoint generation baseline, in which the hypergraph topology is generated first, followed by the feature generation.
+For the graph point cloud datasets, we compare FAHNES against DeFoG [33] and DiGress [6], two state-of-the-art flat models for graph generation.
+4.2 Results and Discussion
+Comparison with the baselines.
+Tables 1 and 3 show the comparisons for the featureless hypergraphs.
+We see that our node-budget and OT-coupling components improve the generation quality.
+Regarding the Manifold40 featured hypergraph dataset [^3] The full set of numerical results for the Manifold40 datasets is provided in Appendix H., FAHNES obtains a Chamfer distance of 0.073 and 0.049 for bench and airplane, respectively, while the sequential baseline reaches 0.143 and 0.117, respectively.
+This shows the better modeling capability of FAHNES in jointly generating topology and features in hypergraphs instead of a simple two-step approach.
+About the graph point cloud datasets, Table 3 shows that our hierarchical method is the only one able to perform at this scale, which shows its superior scalability.
+Detailed numerical results with more metrics are shown in Appendix H.
+Table 4: Ablation studies on the node budget (Budg.) and minibatch OT-coupling (Coup.) for SBM, Ego, and Tree hypergraphs.
+Budg.
+Coup.
+SBM Tree Ego
+Valid $\uparrow$
+NodeNum $\downarrow$
+Spectral $\downarrow$
+Valid $\uparrow$
+NodeNum $\downarrow$
+Spectral $\downarrow$
+Valid $\uparrow$
+NodeNum $\downarrow$
+Spectral $\downarrow$
 […]
-FAHNES 81.4% 0.010 0.603 0.005 0.005 100% 0.162 0.171 0.129 0.007 82.8% 0.000 0.043 0.046 0.002
-Table 2: Evaluation on the ModelNet40 hypergraphs.
-Model ModelNet40 Bookshelf ModelNet40 Piano
-Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
-HyperPA [26] 8.025 7.562 0.044 0.048 0.825 9.254 0.023 0.067
-VAE [27] 47.45 6.190 1.520 0.190 75.35 8.060 1.686 0.396
-GAN [28] 0.000 397.2 46.30 0.456 0.000 409.0 86.38 0.697
-Diffusion [23] 0.000 20.36 2.346 0.079 0.050 20.90 4.192 0.113
-HYGENE [6] 69.73 1.050 0.034 0.068 42.52 6.290 0.027 0.117
 […]
-FAHNES 0.875 5.600 0.087 0.019 0.125 1.878 0.028 0.036
-Table 3: Evaluation on QM9.
-Model QM9
-Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
-DiGress [7] 99.0% 96.2
-DisCo [9] 99.3%
-Cometh [29] 99.6% 96.8 0.25
-DeFoG [30] 99.3% 96.3 0.12
+$\underline{86.7\vbox{\hbox{$\pm$6.7}}}$ $\underline{0.039\vbox{\hbox{$\pm$0.014}}}$ $\mathbf{0.006\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $89.3\vbox{\hbox{$\pm$3.6}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $0.005\vbox{\hbox{$\pm$0.011}}$ $\underline{99.9\vbox{\hbox{$\pm$0.4}}}$ $\mathbf{0.073\vbox{\hbox{$\bm{\pm 0.050}$}}}$ $0.007\vbox{\hbox{$\pm$0.012}}$
+$84.6\vbox{\hbox{$\pm$4.6}}$ $049\vbox{\hbox{$\pm$0.045}} 0.0$$007\vbox{\hbox{$\pm$0.007}}$ $\mathbf{95.6\vbox{\hbox{$\bm{\pm 3.7}$}}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $0.005\vbox{\hbox{$\pm$0.003}}$ $99.5\vbox{\hbox{$\pm$1.1}}$ $\underline{0.117\vbox{\hbox{$\pm$0.155}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.003}$}}}$
+Table 5: Ablation studies on the node budget (Budg.) and minibatch OT-coupling (Coup.) for ModelNet and ManifoldNet datasets.
+Budg.
+Coup.
+ModelNet Bookshelf ModelNet Piano ManifoldNet Airplane ManifoldNet Bench
+NodeNum $\downarrow$
+Spectral $\downarrow$
+Node Num $\downarrow$
+Spectral $\downarrow$ ChamDist $\downarrow$
+Node Num $\downarrow$
+Spectral $\downarrow$ ChamDist $\downarrow$
+NodeNum $\downarrow$
+Spectral $\downarrow$
 […]
-FAHNES 77.8% 94.3 3.86
-Comparison with the baselines.
-Tables 1 and 3 show the comparisons in the featureless hypergraphs.
-We observe that FAHNES significantly improves previous methods in many metrics.
-This is especially evident in the SBM and Ego hypergraphs, where the percentage of valid hypergraphs increases by 16.4% and 10%, respectively.
-Similarly, the spectral similarity has been improved in the ModelNet40 and Tree hypergraphs.
-Table 4: Ablation studies on the node budget (Budg.) and minibatch OT-coupling (Coup.) for SBM, Ego, and Tree hypergraphs.
-Budg. Coup. SBM Hypergraphs Ego Hypergraphs Tree Hypergraphs
-Valid SBM $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Ego $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Valid Tree $\uparrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
-77.7% 0.020 0.889 0.041 0.006 99.0% 2.383 0.239 0.139 0.006 92.2% 0.005 0.066 0.161 0.011
-79.9% 0.123 0.815 0.023 0.006 98.5% 1.721 0.223 0.184 0.010 83.8% 0.000 0.045 0.083 0.006
-71.6% 0.039 1.413 0.005 0.013 100% 0.459 0.129 0.279 0.010 74.5% 0.005 0.039 0.050 0.003
-81.4% 0.010 0.603 0.005 0.005 100% 0.162 0.171 0.129 0.007 82.8% 0.000 0.043 0.046 0.002
-Regarding the Manifold40 dataset, FAHNES obtains a Chamfer distance of 0.073 and 0.049 for bench and airplane, respectively, while the sequential baseline reaches 0.143 and 0.117, respectively.
-This shows the better modeling capacity of FAHNES to jointly generate topology and features in hypergraphs instead of a simple two-step approach.
-Similarly, Table 3 shows the results in the QM9 dataset, where we observe FAHNES presents promising results, even though our model is not explicitly tailored for the problem of molecule generation.
-Extended results on featured hypergraphs and visual comparisons of the different approaches are provided in the supplementary material.
-Table 5: Ablation studies on the node budget (Budg.) and minibatch OT-coupling (Coup.) for ModelNet40 datasets.
-Budg. Coup. ModelNet40 Bookshelf ModelNet40 Piano
-Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$
-8.6 5.951 0.043 0.027 8.125 3.546 0.075 0.026
-7.225 4.183 0.109 0.037 2.375 5.529 0.160 0.028
-1.525 5.232 0.069 0.013 11.275 4.127 0.072 0.027
-0.875 5.600 0.087 0.019 0.125 1.878 0.028 0.036
-Ablation studies.
-Tables 4 and 5 present some ablation studies of FAHNES.
-More precisely, we analyze the importance of the budget and minibatch OT-coupling.
-In general terms, we observe that using budgets instead of concatenating the target size to each node embedding has an important effect on the Node Num, Node Deg, and Edge Size metrics.
-Similarly, the minibatch OT-coupling has a positive effect on the validity of the spectral similarity of the generated hypergraphs.
-We also observe additional interesting properties tied to the budget: i) training is more stable, with smoother curves for the metrics on the validation and test sets, and ii) we observe less variance in the quality of generated samples, which translates into a higher correlation between the metrics of the validation set and those of the test set.
-We hypothesize that the constraints introduced by the node budgets maintain the predictions in the high probability regions.
-Limitations.
-While our node budget mechanism helps mitigate the issue of missing nodes, it does not fully solve it.
-On molecular generation tasks, FAHNES is less competitive due to the maturity of one-shot models in this domain: in hierarchical settings, structural errors accumulate over sequential predictions.
-In addition, FAHNES assumes continuous features, making it less effective for categorical data like atom or bond types—scenarios where discrete diffusion is more appropriate.
-Similarly, specialized techniques such as marginal probabilities [7], designed specifically for molecule generation, cannot be readily integrated into our setting.
-Furthermore, our framework prioritizes scalability, which limits the use of high-capacity backbones such as graph transformers [40].
-Future work could explore discrete diffusion for topology and adopt a mixed discrete-continuous formulation [22] to better handle structured, symbolic domains.
+$940\vbox{\hbox{$\pm$0.917}} 0.0$$032\vbox{\hbox{$\pm$0.013}}$ $3.622\vbox{\hbox{$\pm$1.822}}$ $055\vbox{\hbox{$\pm$0.040}} 0.0$$079\vbox{\hbox{$\pm$0.019}}$ $426\vbox{\hbox{$\pm$0.820}} 0.0$$14\vbox{\hbox{$\pm$0.007}} 0.0$$090\vbox{\hbox{$\pm$0.003}}$ $240\vbox{\hbox{$\pm$0.265}} 0.0$$018\vbox{\hbox{$\pm$0.009}}$
+$\underline{0.265\vbox{\hbox{$\pm$0.496}}}$ $\mathbf{0.014\vbox{\hbox{$\bm{\pm 0.007}$}}}$ $\underline{3.155\vbox{\hbox{$\pm$3.637}}}$ $\mathbf{0.030\vbox{\hbox{$\bm{\pm 0.048}$}}}$ $\underline{0.050\vbox{\hbox{$\pm$0.005}}}$ $\underline{0.052\vbox{\hbox{$\pm$0.065}}}$ $\underline{0.012\vbox{\hbox{$\pm$0.005}}}$ $\underline{0.085\vbox{\hbox{$\pm$0.056}}}$ $\mathbf{0.020\vbox{\hbox{$\bm{\pm 0.032}$}}}$ $\mathbf{0.013\vbox{\hbox{$\bm{\pm 0.003}$}}}$
+[…]
+4.3 Ablation studies
+Tables 4 and 5 present ablation studies of FAHNES for the node-budget and OT-coupling components.
+We observe that using budgets instead of concatenating the target size to each node embedding, like in [9, 10], improves generation quality.
+OT-coupling has a more nuanced effect, clearly improving quality on some datasets, like SBM or ManifoldNet Airplane, while not changing much for others.
+Those two components are also essential for feature generation, as lacking one of them results in much worse results, as seen by the large increase in ChamDist when one component is ablated.
+4.4 Limitations
+While our node budget mechanism helps mitigate the issue of missing nodes, it does not fully resolve it, and our method still struggles when generating very large hypergraphs, such as ModelNet Bookshelf and Piano, or very large graphs, such as point cloud datasets.
+These cases highlight that scalability remains challenging when both the number of nodes and the structural complexity grow substantially.
+Moreover, our framework currently assumes relatively simple feature distributions (e.g., 3D coordinates), and extending it to domains with richer or heterogeneous node and hyperedge attributes (such as categorical or multi-modal features) may require additional modeling components.
+Finally, although FAHNES improves stability through minibatch OT-coupling, training remains computationally expensive for large-scale datasets, which may limit applicability in resource-constrained settings.
 5 Conclusion
-We presented FAHNES, the first hierarchical framework for jointly generating hypergraph topology and features.
-By combining coarse-to-fine structural modeling with feature-aware generation, our method enables scalable synthesis of complex hypergraphs, surpassing the limitations of flat or disjoint approaches.
-Key innovations include a recursive node budget mechanism and the use of minibatch OT-coupling, which together improve sample quality and training convergence.
-Experiments across synthetic, molecular, and geometric datasets demonstrate strong performance in both topology and feature generation, often outperforming existing baselines.
-While FAHNES is not yet competitive with domain-specific pipelines—particularly for molecule generation—it opens promising directions for structured generative modeling.
-Appendix A Implementation Details
-A.1 Model Architecture
-Our method represents the expansion numbers for left and right nodes, along with edge presence, as attributes of the bipartite graph.
-To model the distribution $p(\mathbf{v}^{(l)}_{L},\mathbf{v}^{(l)}_{R},\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}}\mid\tilde{B}^{(l)})$, we adopt an endpoint-parameterized flow-matching framework [21].
-Within this framework, the attributes—namely, the expansion vectors and edge indicators—are corrupted with noise, and a denoising network is trained to reconstruct the original values.
-The denoising network is structured as follows:
-  1. 1.
-    Positional encoding: Node positions within the graph are encoded using SignNet [41]. These encodings are replicated according to the respective expansion numbers.
-  2. 2.
-    Attribute embedding: Five separate linear layers are used to embed the bipartite graph attributes: left node features, right node features, edge features, node-specific features, and hyperedge-specific features.
-FiLM conditioning [25] is applied to incorporate contextual information into node and hyperedge features.
-Node budgets are embedded using sinusoidal positional encodings [42].
-  3. 3.
-    Feature concatenation:
-    • For each left and right node, embeddings are concatenated with positional encodings and the desired reduction fraction. Left nodes also receive the node budget embedding.
-    • If node features are present, they are appended to the left nodes. Likewise, hyperedge features are appended to the right nodes when available.
-    • For edges, embeddings include edge features, concatenated positional encodings of the incident nodes, and the reduction fraction.
-  4. 4.
-    Graph processing: The attribute-enriched bipartite graph is processed through a stack of sparse PPGN layers, following the architecture from [11].
-  5. 5.
-    Output prediction: The final graph representations are passed through three linear projection heads to generate outputs.
-    • Left node head: Predicts expansion values, budget splits, and refined node features.
-    • Right node head: Predicts hyperedge expansions and refined hyperedge features.
-    • Edge head: Predicts edge existence.
-A.2 Flow-matching Framework
-We employ a flow-matching generative modeling framework [21], with endpoint parameterization following [22], equivalent to denoising diffusion models [23] using linear interpolation between the prior $p_{0}$ and target $p_{1}$ distributions.
-The goal is to align two distributions by transporting samples from $p_{0}$ to $p_{1}$ through a learned time-dependent vector field $\mathbf{f}(\mathbf{x},t)$, governed by the ODE:
-$\frac{\partial\mathbf{x}}{\partial t}=\mathbf{f}(\mathbf{x},t),\quad\mathbf{x}_{0}\sim p_{0}.$ (11)
-Here, $\mathbf{f}(\mathbf{x},t)$ is trained such that integrating this ODE produces samples from $p_{1}$.
-Endpoint parameterization. Instead of directly modeling the flow field, we learn the terminal point $\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)$ of the trajectory. The flow can be recovered using:
-$\mathbf{f}(\mathbf{x}_{t},t)=\frac{\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{t}}{1-t}.$ (12)
-Training objective. The model is trained to minimize the expected squared error between the predicted and true endpoints:
-$\mathcal{L}=\mathbb{E}_{t\sim\mathcal{U}(0,1),\>\mathbf{x}_{0}\sim p_{0},\>\mathbf{x}_{1}\sim p_{1}}\left[\left\|\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{1}\right\|^{2}\right],$ (13)
-where $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$ is a linear interpolation between samples from the prior and target distributions.
-Modeling setup. We use different prior distributions depending on the task:
-  • Node and edge predictions: Prior samples $p_{0}$ are drawn from a Gaussian distribution.
-Targets are either $-1$ or $1$, indicating binary decisions (e.g., whether a node is expanded or an edge is retained).
-  • Hyperedge expansion: Similar to node and edge predictions, with targets $-1$, 00, or $1$, encoding the number of expansions (none, one, or two).
-  • Budget fractions: Prior samples $p_{0}$ are drawn from a Dirichlet distribution with concentration parameter $\alpha=1.5$, linearly mapped to $[-1,1]$ via $12x-$ 1.
-The target is the budget fraction of each child node, linearly mapped to $[-1,1]$ via $12x-$ 1.
-If a cluster is not expanded, the corresponding budget becomes $1$.
-Parent cluster budgets are encoded using sinusoidal positional encodings [42] (dimension 32, base frequency $10^{-4}$).
-  • Feature generation: Following [20], we draw prior features from a Gaussian and predict true node features, conditioned on the parent node’s feature using a FiLM layer [25].
-Simplex projection via Von Neumann method.
-When modeling budget fractions, predictions must lie on the probability simplex.
-To ensure this, we project the model’s outputs using the Von Neumann projection [24], which finds the closest point (in Euclidean distance) on the simplex:
-$\Delta^{K}=\left\{\mathbf{x}\in\mathbb{R}^{K}\mid x_{i}\geq 0,\sum_{i=1}^{K}x_{i}=1\right\}.$ (14)
-  1. i)
-    Sort $\mathbf{z}\in\mathbb{R}^{K}$ into a descending vector $\mathbf{u}$, such that $u_{1}\geq u_{2}\geq\dots\geq u_{K}$.
-  2. ii)
-    Find the smallest index $\rho\in\{1,\dots,K\}$ such that:
-    $u_{\rho}-\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right)>0.$ (15)
-  3. iii)
-    Compute the threshold:
-    $\tau=\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right).$ (16)
-  4. iv)
-    The projection is then:
-    $\mathbf{x}^{*}=\max(\mathbf{z}-\tau,0).$ (17)
-A.3 Additional Details
-Perturbed expansion.
-Building on [11, 6], we augment Definitions 4 and 6—which are sufficient for reversing coarsening steps—with additional randomness to enhance generative quality.
-This modification is especially beneficial in low-data regimes where overfitting is a concern.
-Specifically, we introduce a probabilistic mechanism that supplements the set of edges $\tilde{\mathcal{E}}$ by randomly adding edges between node pairs on opposite sides of the bipartite graph that are within a fixed distance in $B$.
-The following definition extends the expansion process (Definition 4) to include this stochastic component.
-Definition 9 (Perturbed hypergraph expansion).
-Let $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ be a bipartite graph, and let $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ denote the left and right cluster size vectors.
-For a given radius $r\in\mathbb{N}$ and probability $0\leq p\leq 1$, we construct $\tilde{B}$ as in Definition 4.
-Additionally, for each pair of distinct nodes $\mathbf{v}_{L}{(p)}\in\tilde{\mathcal{V}_{L}}$ and $\mathbf{v}_{R}{(q)}\in\tilde{\mathcal{V}}_{R}$ that are within a distance of at most $12r+$ 1 in $B$, we independently add an edge $e_{\{p_{i},q_{j}\}}$ to $\tilde{\mathcal{E}}$ with probability $p$.
-Spectral conditioning.
-In line with [43, 11], we incorporate spectral information—specifically, the principal eigenvalues and eigenvectors of the normalized Laplacian—as a form of conditioning during the generative process.
-This technique has been shown to improve the quality of generated graphs.
-To generate $B^{(l)}$ from its coarser form $B^{(l+1)}$, we leverage the approximate spectral invariance under coarsening.
-We compute the $k$ smallest non-zero eigenvalues and their corresponding eigenvectors from the normalized Laplacian matrix $\mathbf{\mathcal{L}}^{(l+1)}$ of $B^{(l+1)}$.
-These eigenvectors are processed using SignNet [41] to produce node embeddings for $B^{(l+1)}$.
-These embeddings are then propagated to the expanded nodes of $B^{(l)}$, helping to preserve structural coherence and facilitate cluster identification.
-The hyperparameter $k$ controls the number of spectral components used.
-Minibatch OT-coupling.
-Minibatch OT-coupling [12, 13] has demonstrated effectiveness in image generation tasks by encouraging models to learn more linear and efficient generative trajectories.
-Rather than sampling from the prior and target distributions independently, the two distributions are sampled jointly through the following procedure.
-First, samples are drawn from the prior distribution (typically Gaussian), followed by sampling corresponding targets (e.g., images).
-Then, the prior samples are reindexed to minimize the total distance between each input and its associated output.
-This coupling yields faster convergence and smoother generative flows, thereby reducing the number of sampling steps required during inference.
-We extend this idea to our setting by treating the children of an expanded node as a minibatch and applying OT-coupling to the prior samples associated with these nodes.
-Importantly, we avoid mixing prior samples between “non-equivalent nodes”, i.e., nodes for which swapping the targets results in a different target hypergraph, as doing so changes the learned distribution.
-To prevent this, we restrict OT-coupling to groups of nodes that only differ with respect to their starting noise, i.e., nodes that are equivalent both topologically and in terms of features.
-To manage computational complexity, we adopt the reasonable assumption that such groups of equivalent nodes correspond to the children of an expanded cluster.
-This assumption holds in practice when the joint topology-feature distribution is sufficiently complex.
-As a result, our method achieves faster convergence and requires fewer sampling steps at inference time, with only a very negligible impact on training efficiency.
-In our framework, expanded clusters yield two child nodes, so minibatch OT-coupling reduces to evaluating two permutations per cluster—an operation that can be efficiently parallelized using tensor operations.
-Algorithm 1 summarizes the process.
-Algorithm 1 Minibatch OT-coupling for coarsening/expansion strategy
-1:Expanded bipartite representation $B$ with clusters $\{V^{(p)}\}$, each $V^{(p)}$ containing 1 or 2 nodes; target samples $\{x_{i}\}$
-2:Reindexed noise samples $\{\tilde{z}_{i}\}$
-3:Sample noise $\{z_{i}\}$ for each node in $B$
-4:for all clusters $V^{(p)}\in B$ do
-5:     if $|V^{(p)}|=1$ then
-6:         No reassignment needed for singleton cluster
-7:     else if $|V^{(p)}|=2$ then
-8:         Let $i,j$ be the indices of the two nodes in $V^{(p)}$
-9:         Let $x_{i},x_{j}$ be their corresponding targets
-10:         Compute normal order cost:
-$C_{\text{normal}}\leftarrow\|z_{i}-x_{i}\|^{2}+\|z_{j}-x_{j}\|^{2}$
-11:         Compute swapped order cost:
-$C_{\text{swap}}\leftarrow\|z_{j}-x_{i}\|^{2}+\|z_{i}-x_{j}\|^{2}$
-12:         if $C_{\text{swap}}<C_{\text{normal}}$ then
-13:              Swap $z_{i}\leftrightarrow z_{j}$
-14:         end if
-15:     end if
-16:end for
-17:return $\{\tilde{z}_{i}\}$ (reassigned noise samples)
-Appendix B Proofs
-B.1 Averaging Node Features for Clusters’ Features
+We presented FAHNES, the first scalable hierarchical framework for jointly generating graph and hypergraph topology together with features.
+By integrating coarse-to-fine structural modeling with feature-aware generation, FAHNES overcomes the limitations of flat or disjoint approaches and enables the generation of complex structures at larger scales.
+Key innovations include a node budget mechanism, which provides fine-grained control over local growth, and minibatch OT-coupling, which improves alignment and stability during training.
+Extensive experiments on synthetic, 3D mesh, and point cloud datasets show that FAHNES achieves strong performance in both topology and feature generation while scaling to graph and hypergraph sizes that are out of reach for prior models.
+Our framework opens new directions for generative modeling of structured data, including applications in 3D geometry and circuit modeling, as well as extensions to richer feature modalities.
+Appendix A Proofs
+A.1 Averaging Node Features for Clusters’ Features
 Proposition 9.
@@
 which is the arithmetic mean of the feature vectors in the cluster $\mathcal{C}$. ∎
-B.2 Minibatch OT-coupling
+A.2 Minibatch OT-coupling
 Let $B$ be a bipartite graph.
@@
 Since $f$ is defined on graphs and graphs are invariant under isomorphism, we have:
-$f(\mathbf{x}_{0})=f(\sigma(\mathbf{x}_{0})).$
+$f(\mathbf{x}_{0})=f(\sigma(\mathbf{x}_{0})).$ (14)
 Therefore:
@@
 $=\mathbb{E}_{q(\mathbf{x}_{0})}[f(\mathbf{x}_{0})].$ (18)
+where 16 comes from the transfer formula.
 Thus, the marginal distribution over $\mathbf{x}_{0}$ remains unchanged. The same argument applies symmetrically for $\mathbf{x}_{1}$ by using $\sigma^{-1}$, concluding the proof. ∎
-Appendix C Complexity Analysis
-In this section, we investigate the asymptotic complexity of our proposed algorithm, which extends the methodology introduced by [11] and [6].
-To construct a hypergraph comprising $n$ nodes, $m$ hyperedges, and $k$ incidences, the algorithm sequentially produces a series of bipartite graphs $B^{(L)}=(\{1\},\{2\},\{(1,2)\}),B^{(L-1)},\dots,B^{(0)}=B$, where the final graph $B$ corresponds to the bipartite representation of the generated hypergraph.
-We use $n$, $m$, and $k$ to denote, respectively, the number of nodes, hyperedges, and incidences in the hypergraph, and as the number of left-side nodes, right-side nodes, and edges in the corresponding bipartite graph.
-[…]
-Next, we assess the computational cost of generating a single expansion step.
-At level $l=L$, this consists of creating a pair of connected nodes, initializing features as matrices of zeros, initializing budget as the targeted node count, and predicting the expansion vectors $\mathbf{v}_{L}$ and $\mathbf{v}_{R}$—a process with constant complexity $\mathcal{O}(1)$.
-For levels $0\leq l<L$, given $B^{(l+1)}$ and expansion vectors $\mathbf{v}^{(l+1)}_{L}$, $\mathbf{v}^{(l+1)}_{R}$, the algorithm constructs the expanded bipartite graph $\tilde{B}(B^{(l+1)},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)})$ in $\mathcal{O}(n+m)$ time.
-It then samples $\mathbf{v}_{L}^{(l)}$, $\mathbf{v}_{R}^{(l)}$, $\mathbf{e}^{(l)}$, $\mathbf{f}$, $\mathbf{F}_{L}^{\text{refine}}$, and $\mathbf{F}_{R}^{\text{refine}}$, and constructs the refined graph $B^{(l)}=B(\tilde{B}^{(l)},\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$.
-Letting $v^{L}_{\max}$ and $v^{R}_{\max}$ be the maximum cluster sizes, the incidence count in $\tilde{B}^{(l)}$ is bounded by $k_{l}\leq k_{l+1}v^{L}_{\max}v^{R}_{\max}$.
-The sampling process queries a denoising model a constant number of times per step.
-The complexity is thus governed by the architecture.
-In our case, since bipartite graphs are triangle-free, the Local PPGN model [11] has linear complexity $\mathcal{O}(n+m+k)$.
-Embedding computation for $B^{(l)}$ similarly costs $\mathcal{O}(n+m+k)$.
-This includes calculating the top $K$ eigenvalues/eigenvectors of the Laplacian via the method from [44], with complexity $\mathcal{O}\left(K(n_{l+1}+m_{l+1}+k_{l+1})\right)$, and embedding via SignNet, also linear in graph size due to fixed $K$.
-The final transformation from the bipartite graph to a hypergraph—by collapsing right-side nodes into hyperedges—has a cost of $\mathcal{O}(m+k)$.
-Under these assumptions, the total complexity to generate a hypergraph $H$ with $n$ nodes, $m$ hyperedges, and $k$ incidences is $\mathcal{O}(n+m+k)$.
-Appendix D Coarsening Sequence Sampling
-[…]
+Proposition 12 (Shorter paths).
+Under the OT-coupling of Algorithm 2, the loss-induced paths for a given minibatch are, on average, shorter than without OT-coupling.
+Formally, let $\pi$ be the permutation returned by Algorithm 2 for starting points $\mathbf{x}^{0}$ and target points $\mathbf{x}^{1}$.
+Then
+$\sum_{i}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}\;\geq\;\sum_{i}\|\mathbf{x}^{0}_{\pi(i)}-\mathbf{x}^{1}_{i}\|^{2}.$ (19)
+Proof.
+We can rewrite the left-hand side of (19) as a sum over clusters:
+$\sum_{i}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}$ $=\sum_{\text{clusters }\mathcal{C}}\;\sum_{i\in\mathcal{C}}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}.$ (20)
+By construction, Algorithm 2 minimizes $\sum_{i\in\mathcal{C}}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}$ within each cluster $\mathcal{C}$.
+Therefore,
+$\sum_{\text{clusters }\mathcal{C}}\sum_{i\in\mathcal{C}}\|\mathbf{x}^{0}_{i}-\mathbf{x}^{1}_{i}\|^{2}$ $\geq\sum_{\text{clusters }\mathcal{C}}\sum_{i\in\mathcal{C}}\|\mathbf{x}^{0}_{\pi(i)}-\mathbf{x}^{1}_{i}\|^{2}$ (21)
+$=\sum_{i}\|\mathbf{x}^{0}_{\pi(i)}-\mathbf{x}^{1}_{i}\|^{2},$ (22)
+which proves the claim. ∎
+Appendix B Coarsening Sequence Sampling
+This section outlines our methodology for sampling a coarsening sequence $\pi\in\Pi_{F}(H)$ for a given hypergraph $H$.
+The full procedure is detailed in Algorithm 1.
+At each coarsening step $l$, let $H^{(l)}$ denote the current hypergraph, $B^{(l)}$ its bipartite representation, and $C^{(l)}$ its weighted clique expansion.
+We begin by sampling a target reduction fraction $\text{red\_frac}\sim\mathcal{U}([\rho_{\text{min}},\rho_{\text{max}}])$.
+We then evaluate all possible contraction sets $F(C^{(l-1)})$ using a cost function $c$, where lower cost indicates higher preference.
+We employ a greedy randomized strategy that processes contraction sets in order of increasing cost.
+For each set:
   • The set is stochastically rejected with probability $1-\lambda$.
@@
 i.e., when the number of nodes on the left side (corresponding to the original hypergraph nodes) has been reduced by the sampled fraction.
-This framework is flexible, allowing a variety of cost functions $c$, contraction families $F$, reduction fraction ranges $[\rho_{\text{min}},\rho_{\text{max}}]$, and randomization parameters $\lambda$.
+This framework is flexible, allowing a variety of cost functions $c$, contraction families $F$, reduction fraction ranges $\rho_{\text{min}},\rho_{\text{max}}]$, and randomization parameters $\lambda$.
 Practical considerations.
@@
   • Cost function: Local Variation Cost [21] with a preserving eigenspace size of $k=8$.
-  • Reduction fraction range: $[\rho_{\text{min}},\rho_{\text{max}}]=[0.1,0.3]$.
+  • Reduction fraction range: $\rho_{\text{min}},\rho_{\text{max}}]=[0.1,0.3]$.
 Algorithm 1 Hypergraph coarsening sequence sampling: Randomized iterative coarsening of a hypergraph.
@@
 Accepted contractions update the hypergraph structure, node budgets, and features.
-1:contraction family $F$, cost function $c$, reduction fraction range $[\rho_{\text{min}},\rho_{\text{max}}]$, randomization parameter $\lambda$
+1:contraction family $F$, cost function $c$, reduction fraction range $\rho_{\text{min}},\rho_{\text{max}}]$, randomization parameter $\lambda$
 2:hypergraph $H$ with $n$ nodes and $m$ hyperedges; node features $\mathbf{F}_{L}$; hyperedge features $\mathbf{F}_{R}$
@@
 7:  $C^{(0)}\leftarrow\text{WeightedCliqueExpansion}(H^{(0)})$
-8:     $\mathbf{b}_{L}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{n}$triangleright Initial node budgets
-9:     $\mathbf{b}_{R}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{m}$triangleright Initial hyperedge budgets
+8:  $\mathbf{b}_{L}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{n}$ $\triangleright$ Initial node budgets
+9:  $\mathbf{b}_{R}^{(0)}\leftarrow(1,\dots,1)\in\mathbb{R}^{m}$ $\triangleright$ Initial hyperedge budgets
 10:  $\pi\leftarrow(B^{(0)},\mathbf{b}_{L}^{(0)},\mathbf{F}_{L},\mathbf{F}_{R})$
@@
 13:   $l\leftarrow l+1$
-[…]
-15:         $f\leftarrow c(\cdot,C^{(l-1)},(\mathcal{P}^{(l-1)},\ldots,\mathcal{P}^{(0)}))$triangleright Cost function
-[…]
+14:   $\text{red\_frac}\sim\text{Uniform}([\rho_{\text{min}},\rho_{\text{max}}])$ $\triangleright$ Sample reduction fraction
+15:   $f\leftarrow c(\cdot,C^{(l-1)},(\mathcal{P}^{(l-1)},\ldots,\mathcal{P}^{(0)}))$ $\triangleright$ Cost function
+16:   $\text{accepted\_contractions}\leftarrow\emptyset$
 17:   for $S\in\text{SortedByCost}(F(C^{(l-1)}))$ do
@@
 21:        $B_{\text{temp}}\leftarrow\text{CoarsenBipartite}(B^{(l-1)},S)$
+22:        if  $\forall\text{ right cluster }R\in B_{\text{temp}}:|R|\leq 3$ then
+23:         $\text{accepted\_contractions}\leftarrow\text{accepted\_contractions}\cup\{S\}$
+24:         $C^{(l)}\leftarrow C_{\text{temp}}$, $B^{(l)}\leftarrow B_{\text{temp}}$
 […]
-[…]
-24:                           $C^{(l)}\leftarrow C_{\text{temp}}$, $B^{(l)}\leftarrow B_{\text{temp}}$
-25$\triangleright$ Update budgets and features for the new cluster
-26:                           Let $\text{Let }S=\{v_{1},\ldots,v_{k}\},\text{ and the new node be }v^{*}$
+26:         $\text{Let }S=\{v_{1},\ldots,v_{k}\},\text{ and the new node be }v^{*}$
 27:         $\mathbf{b}_{L}^{(l)}[v^{*}]\leftarrow\sum_{i=1}^{k}\mathbf{b}_{L}^{(l-1)}[v_{i}]$
@@
 39:end function
-Appendix E Training and Sampling Procedures
-In this section, we present the complete training and inference procedures, detailed in Algorithms 4 and 5. Both pipelines rely on node embeddings produced by Algorithm 3.
-Algorithm 3 Node embedding computation: Here we describe the way the left and right side node embeddings are computed for a given bipartite representation of a hypergraph.
-Embeddings are computed for the input bipartite representation and then replicated according to the cluster size vectors.
-1:number of spectral features $k$
-2:bipartite representation $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$, spectral feature model SignNetθ, cluster size vector $\mathbf{v}_{L}$ and $\mathbf{v}_{R}$
-3:node embeddings computed for all nodes in $\mathcal{V}_{L}$ and $\mathcal{V}_{R}$ and replicated according to $\mathbf{v}_{L}$ and $\mathbf{v}_{R}$
-4:function Embeddings($B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$, SignNetθ, $\mathbf{v}_{L}$, $\mathbf{v}_{R}$)
-5:     if $k=0$ then
-6:         $\mathbf{H}=[h^{(1)},\ldots,h^{(|\mathcal{V}|)}]\overset{i.i.d.}{\sim}\mathcal{N}(0,I)$triangleright Sample random embeddings
-7:     else
-8:         if $k<|\mathcal{V}|$ then
-9:              $[\lambda_{1},\ldots,\lambda_{k}],[u_{1},\ldots,u_{k}]\leftarrow\text{EIG}(B)$triangleright Compute $k$ spectral features
-10:         else
-11:              $[\lambda_{1},\ldots,\lambda_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1}],[u_{1},\ldots,u_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1}]\leftarrow\text{EIG}(B)$triangleright Compute $|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1$ spectral features
-12:              $[\lambda_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|},\ldots,\lambda_{k}],[u_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|},\ldots,u_{k}]\leftarrow[0,\ldots,0],[0,\ldots,0]$triangleright Pad with zeros
-13:         end if
-14:         $\mathbf{H}=[h^{(1)},\ldots,h^{(|\mathcal{V}_{L}|+|\mathcal{V}_{R}|)}]\leftarrow\text{SignNet}_{\theta}([\lambda_{1},\ldots,\lambda_{k}],[u_{1},\ldots,u_{k}],B)$
-15:     end if
-16:     $\tilde{B}=(\mathcal{V}_{L}^{(1)}\cup\cdots\cup\mathcal{V}_{L}^{(p_{l})},\mathcal{V}_{R}^{(1)}\cup\cdots\cup\mathcal{V}_{R}^{(p_{r})},\tilde{\mathcal{E}})\leftarrow\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R})$triangleright Expand as per Definition 4
-17:     set $\tilde{B}$ s.t.
-for all $p_{L}\in[|\mathcal{V}_{L}|]$ and all $p_{R}\in[|\mathcal{V}_{R}|]$: for all $\mathbf{v}_{L}^{(p_{i})}\in\mathcal{V}_{L}^{(p_{l})}$, $\tilde{\mathbf{H}}[p_{i}]=\mathbf{H}[p_{l}]$ and for all $\mathbf{v}_{R}^{(p_{i})}\in\mathcal{V}_{R}^{(p_{r})}$, $\tilde{\mathbf{H}}[p_{i}]=\mathbf{H}[p_{r}]$triangleright Replicate embeddings
-18:     return $\tilde{\mathbf{H}}$
-19:end function
-Algorithm 4 End-to-end training procedure: This describes the entire training procedure for our model.
-1:number of spectral features $k$ for node embeddings
-2:dataset $\mathcal{D}=\{H_{1},\ldots,H_{N}\}$, denoising model GNNθ, spectral feature model SignNetθ
-3:trained model parameters $\theta$
-4:function Train($\mathcal{D}$, GNNθ, SignNetθ)
-5:     while not converged do
-6:         $H\sim\text{Uniform}(\mathcal{D})$triangleright Sample graph
-7:         $(B^{(0)},\ldots,B^{(L)})\leftarrow\text{RndRedSeq}(H)$triangleright Sample coarsening sequence by Algorithm 2
-8:         $l\sim\text{Uniform}(\{0,\ldots,L\})$triangleright Sample level
-9:         if $l=0$ then
-10:              $\mathbf{v}_{L}^{(0)}\leftarrow 1$, $\mathbf{v}_{R}^{(0)}\leftarrow 1$
-11:         else
-12:              set $\mathbf{v}_{L}^{(l)}$ and $\mathbf{v}_{R}^{(l)}$ such that the node sets of $\tilde{B}(B^{(l)},\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)})$ equals that of $B^{(l-1)}$
-13:         end if
-14:         if $l=L$ then
-15:              $B^{(l+1)}\leftarrow B^{(l)}=(\{1\},\{2\},\{(1,2)\},\mathbf{b}=size(H),\mathbf{F}_{L}=0,\mathbf{F}_{R}=0)$
-16:              $\mathbf{v}_{L}^{(l+1)}\leftarrow 1$
-17:              $\mathbf{v}_{R}^{(l+1)}\leftarrow 1$
-18:              $\mathbf{e}^{(l)}\leftarrow 1$
-19:         end if
-20:         set $\mathbf{e}^{(l)}$, $\mathbf{f}$, $\mathbf{F}_{L}^{\text{refine}}$ and $\mathbf{F}_{R}^{\text{refine}}$ such that $B(\tilde{B}(B^{(l+1)},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)}),\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})=B^{(L)}$
-21:         $\mathbf{H}^{(l)}\leftarrow\text{Embeddings}(B^{(l+1)},\text{SignNet}_{\theta},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)})$triangleright Compute node embeddings
-22:         $\hat{\rho}\leftarrow 1-(n^{(l)}/n^{(l-1)})$, with $n^{(l)}$ and $n^{(l-1)}$ being the size of the left side of $B^{(l)}$ and $B^{(l-1)}$
-23:         $D_{\theta}\leftarrow\text{GNN}_{\theta}(\cdot,\cdot,\tilde{B}^{(l)},\mathbf{H}^{(l)},n^{(0)},\rho)$, where $n^{(0)}$ is the size of the left side of $B^{(0)}$
-24:         take gradient descent step on $\nabla_{\theta}\text{DiffusionLoss}(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)},\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}},D_{\theta})$
-25:     end while
-26:     return $\theta$
-27:end function
-Algorithm 5 End-to-end sampling procedure with deterministic expansion size: This describes the sampling procedure.
-Note that this assumes that the maximum cluster sizes are 2 and 3, which is the case when using edges of the clique representation as the contraction set family for model training.
-1:reduction fraction range $[\rho_{\text{min}},\rho_{\text{max}}]$
-2:target hypergraph size $N$, denoising model GNNθ, spectral feature model SignNetθ
-3:sampled hypergraph $H=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=N$
-4:function Sample($N$, GNNθ, SignNetθ)
-5:     $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})\leftarrow(\{1\},\{2\},\{(1,2)\},N,0,0)$triangleright Start with a minimal bipartite graph
-6:     $\mathbf{v}_{L}\leftarrow[1]$, $\mathbf{v}_{R}\leftarrow[1]$triangleright Initial cluster size vectors
-7:     while $|\mathcal{V}_{L}|<N$ do
-8:         $\mathbf{H}\leftarrow\text{Embeddings}(B,\text{SignNet}_{\theta},\mathbf{v}_{L},\mathbf{v}_{R})$triangleright Compute node embeddings
-9:         $n\leftarrow\|\mathbf{v}_{L}\|_{1}$
-10:         $\rho\sim\text{Uniform}([\rho_{\text{min}},\rho_{\text{max}}])$triangleright random reduction fraction
-11:         set $n^{+}$ s.t. $n^{+}=\lceil\rho(n+n^{+})\rceil$triangleright number of left side nodes to add
-12:         $n^{+}\leftarrow\min(n^{+},N-n)$triangleright ensure not to exceed target size
-13:         $\hat{\rho}\leftarrow 1-(n/(n+n^{+}))$triangleright actual reduction fraction
-14:         $D_{\theta}\leftarrow\text{GNN}_{\theta}(\cdot,\cdot,\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R}),\mathbf{H},N,\hat{\rho})$
-15:         $(\mathbf{v}_{L})_{0},(\mathbf{v}_{R})_{0},(\mathbf{e})_{0},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}}\leftarrow\text{Sample}(D_{\theta})$triangleright Sample features
-16:         set $\mathbf{v}_{L}$ s.t. for $i\in[n]$: $\mathbf{v}_{L}[i]=2$ if |$|\{j\in[n]\mid(\mathbf{v}_{L})_{0}[j]\geq(\mathbf{v}_{L})_{0}[i]\}|\geq n^{+}$ and $v[i]=1$ otherwise
-17:         set $\mathbf{v}_{R}$ s.t.
-for $i\in[|(\mathbf{v}_{R})_{0}|]$: $\mathbf{v}_{R}[i]=1$ if $(\mathbf{v}_{R})_{0}<1.66$, $\mathbf{v}_{R}[i]=2$ if $(\mathbf{v}_{R})_{0}<2.33$ and $\mathbf{v}_{R}[i]=3$ otherwise
-18:         set $\mathbf{e}$ s.t. for $i\in[|(\mathbf{e})_{0}|]$: $\mathbf{e}[i]=1$ if $(\mathbf{e})_{0}>0.5$ and $\mathbf{e}[i]=0$ otherwise
-19:         $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})\leftarrow B(\tilde{B},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$triangleright Refine as per Definition 6
-20:     end while
-21:     build $H$ from its bipartite representation $B$
-22:     return $H$
-23:end function
-Appendix F Experimental Details
+Appendix C Examples of Coarsening Sequences
+Figure 3: Examples of coarsening sequence for various meshes. Thick lines represent 2-edges.
+Appendix D Experimental Details
+In this section, we detail all three types of experiments – unfeatured hypergraphs, 3D meshes and graph point clouds – individually, detailing their datasets, baselines, metrics and specific hyperparameters.
 For all experiments, we use embeddings with 32 dimensions for edge selection vectors and node and hyperedge expansion numbers.
@@
 We always use 25 sampling steps.
-All experiments are run for 24 hours on a single L40S.
+All experiments are run for 1M steps on a single L40S.
 We use 8 CPU workers.
-F.1 Experimental Details: Unfeatured Hypergraphs
+D.1 Unfeatured Hypergraphs
 Datasets. Our experiments utilize five datasets: three synthetic and two real-world, consistent with those described in [10]:
@@
 Our model has 4M parameters.
-Detailed numerical results.
-Table 6: Detailed numerical results for ablation studies.
-SBM Hypergraphs
-Node Budget Minibatch OT Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
-77.7% 0.020 0.889 0.041 0.006 5.118 0.015 0.002
-79.9% 0.123 0.815 0.023 0.006 6.3 0.015 0.003
-71.6% 0.039 1.413 0.005 0.013 9.862 0.006 0.004
-81.4% 0.010 0.603 0.005 0.005 3.962 0.009 0.002
-Ego Hypergraphs
-Node Budget Minibatch OT Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
-99.0% 2.383 0.239 0.139 0.006 5.720 0.007 4e-6
-98.5% 1.721 0.223 0.184 0.010 3.784 6.7e-4 1.8e-5
-100% 0.459 0.129 0.279 0.010 3.010 0 0
-100% 0.162 0.171 0.129 0.007 6.061 0 0
-Tree Hypergraphs
-Node Budget Minibatch OT Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
-92.2% 0.005 0.066 0.161 0.011 0.406 0.022 0.006
-83.8% 0 0.045 0.083 0.006 0.515 0.025 0.027
-74.5% 0.005 0.039 0.050 0.003 0.528 0.042 0.026
-82.8% 0 0.043 0.046 0.002 0.258 0.040 0.027
-ModelNet Bookshelf
-Node Budget Minibatch OT NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
-8.615 5.951 0.043 0.027 102.547 0.113 0.003
-7.225 4.183 0.109 0.037 63.778 0.197 0.003
-1.525 5.232 0.069 0.013 79.016 0.087 0.003
-0.875 5.600 0.087 0.019 32.723 0.113 0.002
-ModelNet Piano
-Node Budget Minibatch OT NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
-8.125 3.546 0.075 0.026 42.991 0.102 0.002
-2.375 5.529 0.160 0.028 50.349 0.097 0.002
-11.275 4.127 0.072 0.027 31.312 0.168 0.002
-0.125 1.878 0.028 0.036 43.326 0.072 0.001
-F.2 3D Meshes
+D.2 3D Meshes
 Datasets.
@@
 Metrics.
-We use the same metrics as for hypergraphs, which are: NodeNumDiff, NodeDegree, EdgeSize, and Spectral.
-To this we add a metric NearChamDist which computes the Chamfer distance between a point cloud sampled from the surface of the generated mesh and equivalent point clouds sampled from the validation/test set meshes.
+We use the same metrics as for unfeatured hypergraphs.
+To this we add a metric ChamDist which computes the minimal Chamfer distance between a point cloud sampled from the surface of the generated mesh and equivalent point clouds sampled from all validation/test set meshes.
 Baselines. We compare against a simple sequential baseline:
@@
 Our model has 6M parameters.
-Detailed numerical results.
-Table 7: Evaluation on Manifold40 Bench and Airplane datasets.
-Method Manifold40 Bench Manifold40 Airplane
-Cham Dist $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Centr Harm $\downarrow$ Centr Clos $\downarrow$ Centr Betw $\downarrow$ Cham Dist $\downarrow$ Node Num $\downarrow$ Node Deg $\downarrow$ Edge Size $\downarrow$ Spectral $\downarrow$ Centr Harm $\downarrow$ Centr Clos $\downarrow$ Centr Betw $\downarrow$
-Sequential 0.143 0.367 0.801 0.004 0.007 4.964 0.087 0.006 0.117 0.078 0.332 0.011 0.015 3.131 0.035 0.007
+D.3 Graph Point Clouds
+Datasets.
+We reuse the same datasets as for 3D meshes.
+Point clouds comprising 1024 nodes are sampled on the surface of each mesh, then each node is connected to its 3 nearest neighbors.
+Only the largest connected component is kept.
+Metrics. Evaluation Metrics. We adopt the same evaluation protocol as [35] to assess the quality of generated graphs, reporting the following metrics:
+  • NumDiff: MMD between node counts distributions of generated and reference graphs.
+  • Deg: MMD between degree distributions of generated and reference graphs.
+  • Clustering: MMD between clustering coefficient distributions.
+  • Orbit: MMD between graphlet orbit count distributions.
+  • Spectral: MMD between Laplacian spectra.
+  • Wavelet: MMD between wavelet coefficient distributions.
+  • Ratio: Average ratio of generated-to-reference values across the above metrics, used as a global indicator of statistical similarity.
+To this we add a metric ChamDist which computes the minimal Chamfer distance between the generated mesh and all validation/test set point clouds.
+Baselines.
+We compare against two state-of-the-art flat generative models: Defog [33], DiGress [6].
+As these two methods are tailored for discrete data, we use a mixed-discrete continuous framework for continuous features.
+Specific hyperparameters.
+We use $\lambda=0.3$, our Local PPGN layers have a dimension of 128, and the hidden dimension for our MLP is 256.
+We use perturbed hypergraph expansion with a radius of 2 and dropout of 0.5.
+Our model has 4M parameters.
+Appendix E Implementation Details
+E.1 Flow-matching Framework
+We employ a flow-matching generative modeling framework [22], with endpoint parameterization following [36], equivalent to denoising diffusion models [32] using linear interpolation between the prior $p_{0}$ and target $p_{1}$ distributions.
+The goal is to align two distributions by transporting samples from $p_{0}$ to $p_{1}$ through a learned time-dependent vector field $\mathbf{f}(\mathbf{x},t)$, governed by the ODE:
+$\frac{\partial\mathbf{x}}{\partial t}=\mathbf{f}(\mathbf{x},t),\quad\mathbf{x}_{0}\sim p_{0}.$ (23)
+Here, $\mathbf{f}(\mathbf{x},t)$ is trained such that integrating this ODE produces samples from $p_{1}$.
+Endpoint parameterization. Instead of directly modeling the flow field, we learn the terminal point $\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)$ of the trajectory. The flow can be recovered using:
+$\mathbf{f}(\mathbf{x}_{t},t)=\frac{\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{t}}{1-t}.$ (24)
+Training objective. The model is trained to minimize the expected squared error between the predicted and true endpoints:
+$\mathcal{L}=\mathbb{E}_{t,\>\mathbf{x}_{0}\sim p_{0},\>\mathbf{x}_{1}\sim p_{1}}\left[\left\|\hat{\mathbf{x}}_{1}(\mathbf{x}_{t},1)-\mathbf{x}_{1}\right\|^{2}\right],$ (25)
+where $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\mathbf{x}_{1}$ is a linear interpolation between samples from the prior and target distributions.
+Following [37], during training $t$ is sampled using a log-normal distribution, which has been shown to improve performance.
+Prior distributions. We use different prior distributions depending on the task:
+  • Node and edge predictions: Prior samples $p_{0}$ are drawn from a Gaussian distribution.
+Targets are either 1-1 or $1$, indicating binary decisions (e.g., whether a node is expanded or an edge is retained).
+  • Hyperedge expansion: Similar to node and edge predictions, with targets 1-1, 0, or $1$, encoding the number of expansions (none, one, or two).
+  • Budget fractions: Prior samples $p_{0}$ are drawn from a Dirichlet distribution with concentration parameter $\alpha=1.5$, linearly mapped to [1,1][-1,1] via 2x12x-1.
+The target is the budget fraction of each child node, linearly mapped to [1,1][-1,1] via 2x12x-1.
+If a cluster is not expanded, the corresponding budget becomes $1$.
+Parent cluster budgets are encoded using sinusoidal positional encodings [38] (dimension 32, base frequency $10^{-4}$).
+  • Feature generation: Following [23], we draw prior features from a Gaussian and predict true node features, conditioned on the parent node’s feature using a FiLM layer [39].
+Simplex projection via Von Neumann method.
+When modeling budget fractions, predictions must lie on the probability simplex.
+To ensure this, we project the model’s outputs using the Von Neumann projection [40], which finds the closest point (in Euclidean distance) on the simplex:
+$\Delta^{K}=\left\{\mathbf{x}\in\mathbb{R}^{K}\mid x_{i}\geq 0,\sum_{i=1}^{K}x_{i}=1\right\}.$ (26)
+  1. i)
+    Sort $\mathbf{z}\in\mathbb{R}^{K}$ into a descending vector $\mathbf{u}$, such that $u_{1}\geq u_{2}\geq\dots\geq u_{K}$.
+  2. ii)
+    Find the smallest index $\rho\in\{1,\dots,K\}$ such that:
+    $u_{\rho}-\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right)>0.$ (27)
+  3. iii)
+    Compute the threshold:
+    $\tau=\frac{1}{\rho}\left(\sum_{j=1}^{\rho}u_{j}-1\right).$ (28)
+  4. iv)
+    The projection is then:
+    $\mathbf{x}^{*}=\max(\mathbf{z}-\tau,0).$ (29)
+Graph inpainting.
+During generation, we use inpainting to enforce constraints: i) budget splits are fixed to 1 for unexpanded clusters or those with a budget of $1$, ii) equal splits are enforced for expanded clusters with a budget of 2, iii) clusters of size one are not allowed to expand; and iv) features of non-expanded clusters are copied unchanged.
+E.2 Model Architecture
+Our method represents the expansion numbers for left and right nodes, along with edge presence, as attributes of the bipartite graph.
+To model the distribution $p(\mathbf{v}^{(l)}_{L},\mathbf{v}^{(l)}_{R},\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}}\mid\tilde{B}^{(l)})$, we adopt an endpoint-parameterized flow-matching framework [22].
+Within this framework, the attributes—namely, the expansion vectors and edge indicators—are corrupted with noise, and a denoising network is trained to reconstruct the original values.
+The denoising network is structured as follows:
+  1. 1.
+    Positional encoding: Node positions within the graph are encoded using SignNet [41]. These encodings are replicated according to the respective expansion numbers.
+  2. 2.
+    Attribute embedding: Five separate linear layers are used to embed the bipartite graph attributes: left node features, right node features, edge features, node-specific features, and hyperedge-specific features.
+FiLM conditioning [39] is applied to incorporate contextual information into node and hyperedge features.
+Node budgets are embedded using sinusoidal positional encodings [38].
+  3. 3.
+    Feature concatenation:
+    • For each left and right node, embeddings are concatenated with positional encodings and the desired reduction fraction. Left nodes also receive the node budget embedding.
+    • If node features are present, they are appended to the left nodes. Likewise, hyperedge features are appended to the right nodes when available.
+    • For edges, embeddings include edge features, concatenated positional encodings of the incident nodes, and the reduction fraction.
+  4. 4.
+    Graph processing: The attribute-enriched bipartite graph is processed through a stack of sparse PPGN layers, following the architecture from [9].
+  5. 5.
+    Output prediction: The final graph representations are passed through three linear projection heads to generate outputs.
+    • Left node head: Predicts expansion values, budget splits, and refined node features.
+    • Right node head: Predicts hyperedge expansions and refined hyperedge features.
+    • Edge head: Predicts edge existence.
+E.3 Additional Details
+Perturbed expansion.
+Building on [9, 10], we augment Definitions 4 and 6—which are sufficient for reversing coarsening steps—with additional randomness to enhance generative quality.
+This modification is especially beneficial in low-data regimes where overfitting is a concern.
+Specifically, we introduce a probabilistic mechanism that supplements the set of edges $\tilde{\mathcal{E}}$ by randomly adding edges between node pairs on opposite sides of the bipartite graph that are within a fixed distance in $B$.
+The following definition extends the expansion process (Definition 4) to include this stochastic component.
+Definition 13 (Perturbed hypergraph expansion).
+Let $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ be a bipartite graph, and let $\mathbf{v}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{v}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ denote the left and right cluster size vectors.
+For a given radius $r\in\mathbb{N}$ and probability $0\leq p\leq 1$, we construct $\tilde{B}$ as in Definition 4.
+Additionally, for each pair of distinct nodes $\mathbf{v}_{L}{(p)}\in\tilde{\mathcal{V}_{L}}$ and $\mathbf{v}_{R}{(q)}\in\tilde{\mathcal{V}}_{R}$ that are within a distance of at most $2r+1$ in $B$, we independently add an edge $e_{\{p_{i},q_{j}\}}$ to $\tilde{\mathcal{E}}$ with probability $p$.
+Spectral conditioning.
+In line with [35, 9], we incorporate spectral information—specifically, the principal eigenvalues and eigenvectors of the normalized Laplacian—as a form of conditioning during the generative process.
+This technique has been shown to improve the quality of generated graphs.
+To generate $B^{(l)}$ from its coarser form $B^{(l+1)}$, we leverage the approximate spectral invariance under coarsening.
+We compute the $k$ smallest non-zero eigenvalues and their corresponding eigenvectors from the normalized Laplacian matrix $\mathbf{\mathcal{L}}^{(l+1)}$ of $B^{(l+1)}$.
+These eigenvectors are processed using SignNet [41] to produce node embeddings for $B^{(l+1)}$.
+These embeddings are then propagated to the expanded nodes of $B^{(l)}$, helping to preserve structural coherence and facilitate cluster identification.
+The hyperparameter $k$ controls the number of spectral components used.
+Minibatch OT-coupling.
+Minibatch OT-coupling [12, 13] accelerates training by jointly sampling priors and targets and reindexing priors to minimize input–output distance, leading to smoother flows and fewer inference steps.
+As detailed in Section 3.5, we adapt this idea by treating each side of the bipartite graph as a minibatch and applying OT-coupling within them.
+To preserve the distribution, we restrict permutations to groups of equivalent nodes (identical topology and features, differing only in noise).
+For efficiency, we assume equivalence only among nodes from the same cluster expansion.
+Since each cluster produces two or three children, coupling reduces to evaluating two or six possible permutations per cluster, which can be efficiently parallelized with tensors (Algorithm 2).
+Algorithm 2 Minibatch OT-coupling for coarsening/expansion strategy
+1:Expanded bipartite representation $B$ with clusters $\{V^{(p)}\}$, each $V^{(p)}$ containing 1 or 2 nodes; target samples $\{x_{i}\}$
+2:Reindexed noise samples $\{\tilde{z}_{i}\}$
+3:Sample noise $\{z_{i}\}$ for each node in $B$
+4:for all clusters $V^{(p)}\in B$ do
+5:  if $|V^{(p)}|=1$ then
+6:   No reassignment needed for singleton cluster
+7:  else if $|V^{(p)}|=2$ then
+8:   Let $i,j$ be the indices of the two nodes in $V^{(p)}$
+9:   Let $x_{i},x_{j}$ be their corresponding targets
+10:   Compute normal order cost:
+$C_{\text{normal}}\leftarrow\|z_{i}-x_{i}\|^{2}+\|z_{j}-x_{j}\|^{2}$
+11:   Compute swapped order cost:
+$C_{\text{swap}}\leftarrow\|z_{j}-x_{i}\|^{2}+\|z_{i}-x_{j}\|^{2}$
+12:   if $C_{\text{swap}}<C_{\text{normal}}$ then
+13:     Swap $z_{i}\leftrightarrow z_{j}$
+14:   end if
+15:  end if
+16:end for
+17:return $\{\tilde{z}_{i}\}$ (reassigned noise samples)
+Appendix F Training and Sampling Procedures
+In this section, we present the complete training and inference procedures, detailed in Algorithms 4 and 5. Both pipelines rely on node embeddings produced by Algorithm 3.
+Algorithm 3 Node embedding computation: Here we describe the way the left and right side node embeddings are computed for a given bipartite representation of a hypergraph.
+Embeddings are computed for the input bipartite representation and then replicated according to the cluster size vectors.
+1:number of spectral features $k$
+2:bipartite representation $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$, spectral feature model SignNetθ, cluster size vector $\mathbf{v}_{L}$ and $\mathbf{v}_{R}$
+3:node embeddings computed for all nodes in $\mathcal{V}_{L}$ and $\mathcal{V}_{R}$ and replicated according to $\mathbf{v}_{L}$ and $\mathbf{v}_{R}$
 […]
-FAHNES 0.073 0.067 0.581 0.008 0.014 5.28 0.047 0.004 0.049 0 0.304 0.033 0.015 1.837 0.231 0.004
-F.3 Molecules
-Datasets. We use the QM9 dataset with implicit hydrogens. We generate 20,000 molecules for the metrics.
-Metrics. We use the same metrics as DeFoG [30], that is:
-  • Valid: Fraction of generated samples that can be sanitized using RDKit.
-  • Novel: Fraction of generated samples whose SMILES are not present in the training dataset.
-  • Unique: Fraction of generated samples whose SMILES are unique.
-  • FCD: Fréchet ChemNet Distance [39] between the generated samples and test set.
-Baselines.
-Specific hyperparameters.
-We use $\lambda=0.1$, our Local PPGN layers have a dimension of 128, and the hidden dimension for our MLP is 256.
-We do not use perturbed hypergraph expansion.
-Our model has 4M parameters.
-Appendix G Validation Curves
-Figure 4: Valid SBM with and without budgets.
-Figure 5: Valid Ego with and without budgets.
-Figure 6: Valid Tree with and without budgets.
-Figure 7: NodeNumDiff for Bookshelf with and without budgets.
-Appendix H Examples of Coarsening Sequences
-Figure 8: Examples of coarsening sequence for various meshes. Broad lines represent 2-edges.
+5:  if $k=0$ then
+6:   $\mathbf{H}=[h^{(1)},\ldots,h^{(|\mathcal{V}|)}]\overset{i.i.d.}{\sim}\mathcal{N}(0,I)$ $\triangleright$ Sample random embeddings
+7:  else
+8:   if $k<|\mathcal{V}|$ then
+9:     $\lambda_{1},\ldots,\lambda_{k}],[u_{1},\ldots,u_{k}]\leftarrow\text{EIG}(B) \triangleright$ Compute $k$ spectral features
+10:   else
+11:     $\lambda_{1},\ldots,\lambda_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1}],[u_{1},\ldots,u_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1}]\leftarrow\text{EIG}(B) \triangleright$ Compute $|\mathcal{V}_{L}|+|\mathcal{V}_{R}|-1$ spectral features
+12:     $\lambda_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|},\ldots,\lambda_{k}],[u_{|\mathcal{V}_{L}|+|\mathcal{V}_{R}|},\ldots,u_{k}]\leftarrow[0,\ldots,0],[0,\ldots,0] \triangleright$ Pad with zeros
+13:   end if
+14:   $\mathbf{H}=[h^{(1)},\ldots,h^{(|\mathcal{V}_{L}|+|\mathcal{V}_{R}|)}]\leftarrow\text{SignNet}_{\theta}([\lambda_{1},\ldots,\lambda_{k}],[u_{1},\ldots,u_{k}],B)$
+15:  end if
+16:  $\tilde{B}=(\mathcal{V}_{L}^{(1)}\cup\cdots\cup\mathcal{V}_{L}^{(p_{l})},\mathcal{V}_{R}^{(1)}\cup\cdots\cup\mathcal{V}_{R}^{(p_{r})},\tilde{\mathcal{E}})\leftarrow\tilde{B}(B,\mathbf{v}_{L},\mathbf{v}_{R})$ $\triangleright$ Expand as per Definition 4
+17:  set $\tilde{B}$ s.t.
+for all $p_{L}\in[|\mathcal{V}_{L}|]$ and all $p_{R}\in[|\mathcal{V}_{R}|]$: for all $\mathbf{v}_{L}^{(p_{i})}\in\mathcal{V}_{L}^{(p_{l})}$, $\tilde{\mathbf{H}}[p_{i}]=\mathbf{H}[p_{l}]$ and for all $\mathbf{v}_{R}^{(p_{i})}\in\mathcal{V}_{R}^{(p_{r})}$, $\tilde{\mathbf{H}}[p_{i}]=\mathbf{H}[p_{r}]$ $\triangleright$ Replicate embeddings
+18:  return $\tilde{\mathbf{H}}$
+19:end function
+Algorithm 4 End-to-end training procedure: This describes the entire training procedure for our model.
+1:number of spectral features $k$ for node embeddings
+2:dataset $\mathcal{D}=\{H_{1},\ldots,H_{N}\}$, denoising model GNNθ, spectral feature model SignNetθ
+3:trained model parameters $\theta$
+4:function Train($\mathcal{D}$, GNNθ, SignNetθ)
+5:  while not converged do
+6:   $H\sim\text{Uniform}(\mathcal{D})$ $\triangleright$ Sample graph
+7:   $(B^{(0)},\ldots,B^{(L)})\leftarrow\text{RndRedSeq}(H)$ $\triangleright$ Sample coarsening sequence by Algorithm 1
+8:   $l\sim\text{Uniform}(\{0,\ldots,L\})$ $\triangleright$ Sample level
+9:   if $l=0$ then
+10:     $\mathbf{v}_{L}^{(0)}\leftarrow 1$, $\mathbf{v}_{R}^{(0)}\leftarrow 1$
+11:   else
+12:     set $\mathbf{v}_{L}^{(l)}$ and $\mathbf{v}_{R}^{(l)}$ such that the node sets of $\tilde{B}(B^{(l)},\mathbf{v}_{L}^{(l)},\mathbf{v}_{R}^{(l)})$ equals that of $B^{(l-1)}$
+13:   end if
+14:   if $l=L$ then
+15:     $B^{(l+1)}\leftarrow B^{(l)}=(\{1\},\{2\},\{(1,2)\},\mathbf{b}=size(H),\mathbf{F}_{L}=0,\mathbf{F}_{R}=0)$
+16:     $\mathbf{v}_{L}^{(l+1)}\leftarrow 1$
+17:     $\mathbf{v}_{R}^{(l+1)}\leftarrow 1$
+18:     $\mathbf{e}^{(l)}\leftarrow 1$
+19:   end if
+20:   set $\mathbf{e}^{(l)}$, $\mathbf{f}$, $\mathbf{F}_{L}^{\text{refine}}$ and $\mathbf{F}_{R}^{\text{refine}}$ such that $B(\tilde{B}(B^{(l+1)},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)}),\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})=B^{(L)}$
+21:   $\mathbf{H}^{(l)}\leftarrow\text{Embeddings}(B^{(l+1)},\text{SignNet}_{\theta},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)})$ $\triangleright$ Compute node embeddings
+22:   $\hat{\rho}\leftarrow 1-(n^{(l)}/n^{(l-1)})$, with $n^{(l)}$ and $n^{(l-1)}$ being the size of the left side of $B^{(l)}$ and $B^{(l-1)}$
+23:   $D_{\theta}\leftarrow\text{GNN}_{\theta}(\cdot,\cdot,\tilde{B}^{(l)},\mathbf{H}^{(l)},n^{(0)},\rho)$, where $n^{(0)}$ is the size of the left side of $B^{(0)}$
+24:   take gradient descent step on $\nabla_{\theta}\text{DiffusionLoss}(\mathbf{v}_{L}^{(L)},\mathbf{v}_{R}^{(L)},\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}},D_{\theta})$
+25:  end while
+26:  return $\theta$
+27:end function
+Algorithm 5 End-to-end sampling procedure with deterministic expansion size: This describes the sampling procedure.
+Note that this assumes that the maximum cluster sizes are 2 and 3, which is the case when using edges of the clique representation as the contraction set family for model training.
+1:reduction fraction range $\rho_{\text{min}},\rho_{\text{max}}]$
+2:target hypergraph size NN, denoising model GNNθ, spectral feature model SignNetθ
+3:sampled hypergraph $H=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=N$
+4:function Sample(NN, GNNθ, SignNetθ)
+5:  $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})\leftarrow(\{1\},\{2\},\{(1,2)\},N,0,0)$ $\triangleright$ Start with a minimal bipartite graph
+6:  $\mathbf{v}_{L}\leftarrow[1]$, $\mathbf{v}_{R}\leftarrow[1]$ $\triangleright$ Initial cluster size vectors
+7:  while $|\mathcal{V}_{L}|<N$ do
+8:   $\mathbf{H}\leftarrow\text{Embeddings}(B,\text{SignNet}_{\theta},\mathbf{v}_{L},\mathbf{v}_{R})$ $\triangleright$ Compute node embeddings
+9:   $n\leftarrow\|\mathbf{v}_{L}\|_{1}$
+10:   $\rho\sim\text{Uniform}([\rho_{\text{min}},\rho_{\text{max}}])$ $\triangleright$ random reduction fraction
+11:   set $n^{+}$ s.t. $n^{+}=\lceil\rho(n+n^{+})\rceil$ $\triangleright$ number of left side nodes to add
+12:   $n^{+}\leftarrow\min(n^{+},N-n)$ $\triangleright$ ensure not to exceed target size
+13:   $\hat{\rho}\leftarrow 1-(n/(n+n^{+}))$ $\triangleright$ actual reduction fraction
+[…]
+15:   $(\mathbf{v}_{L})_{0},(\mathbf{v}_{R})_{0},(\mathbf{e})_{0},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}}\leftarrow\text{Sample}(D_{\theta})$ $\triangleright$ Sample features
+16:   set $\mathbf{v}_{L}$ s.t. for $i\in[n]$: $\mathbf{v}_{L}[i]=2$ if $|\{j\in[n]\mid(\mathbf{v}_{L})_{0}[j]\geq(\mathbf{v}_{L})_{0}[i]\}|\geq n^{+}$ and $v[i]=1$ otherwise
+17:   set $\mathbf{v}_{R}$ s.t.
+for $i\in[|(\mathbf{v}_{R})_{0}|]$: $\mathbf{v}_{R}[i]=1$ if $(\mathbf{v}_{R})_{0}<1.66$, $\mathbf{v}_{R}[i]=2$ if $(\mathbf{v}_{R})_{0}<2.33$ and $\mathbf{v}_{R}[i]=3$ otherwise
+18:   set $\mathbf{e}$ s.t. for $i\in[|(\mathbf{e})_{0}|]$: $\mathbf{e}[i]=1$ if $(\mathbf{e})_{0}>0.5$ and $\mathbf{e}[i]=0$ otherwise
+19:   $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})\leftarrow B(\tilde{B},\mathbf{e},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$ $\triangleright$ Refine as per Definition 6
+20:  end while
+21:  build $H$ from its bipartite representation $B$
+22:  return $H$
+23:end function
+Appendix G Complexity Analysis
+In this section, we investigate the asymptotic complexity of our proposed algorithm, which extends the methodology introduced by [9] and [10].
+To construct a hypergraph comprising $n$ nodes, $m$ hyperedges, and $k$ incidences, the algorithm sequentially produces a series of bipartite graphs $B^{(L)}=(\{1\},\{2\},\{(1,2)\}),B^{(L-1)},\dots,B^{(0)}=B$, where the final graph $B$ corresponds to the bipartite representation of the generated hypergraph.
+We use $n$, $m$, and $k$ to denote, respectively, the number of nodes, hyperedges, and incidences in the hypergraph, and as the number of left-side nodes, right-side nodes, and edges in the corresponding bipartite graph.
+For each level $0\leq l<L$ of the sequence, the number of left-side nodes in $B^{(l)}$, denoted $n_{l}$, satisfies $n_{l}\geq(1+\epsilon)n_{l-1}$ for some $\epsilon>0$ (e.g., $\epsilon=\texttt{reduction\_frac}/(1-\texttt{reduction\_frac})$).
+This implies an upper bound on the number of steps in the expansion sequence: $\lceil\log_{1+\epsilon}n\rceil\in\mathcal{O}(\log n)$.
+Since the expansion process only increases node counts, all $B_{l}$ graphs contain fewer than $n$ left-side and $m$ right-side nodes.
+The number of edges, however, may temporarily exceed $k$, as the intermediate bipartite graphs may include additional edges removed in later refinements.
+Still, because the coarsening during training consistently reduces incidences, the model is expected to learn accurate edge refinement and avoid such accumulation.
+Consequently, we assume $k_{l}\leq k$ and $m_{l}\leq m$ for all $0\leq l\leq L$.
+Next, we assess the computational cost of generating a single expansion step.
+At level $l=L$, this consists of creating a pair of connected nodes, initializing features as matrices of zeros, initializing budget as the targeted node count, and predicting the expansion vectors $\mathbf{v}_{L}$ and $\mathbf{v}_{R}$—a process with constant complexity $\mathcal{O}(1)$.
+For levels $0\leq l<L$, given $B^{(l+1)}$ and expansion vectors $\mathbf{v}^{(l+1)}_{L}$, $\mathbf{v}^{(l+1)}_{R}$, the algorithm constructs the expanded bipartite graph $\tilde{B}(B^{(l+1)},\mathbf{v}_{L}^{(l+1)},\mathbf{v}_{R}^{(l+1)})$ in $\mathcal{O}(n+m)$ time.
+It then samples $\mathbf{v}_{L}^{(l)}$, $\mathbf{v}_{R}^{(l)}$, $\mathbf{e}^{(l)}$, $\mathbf{f}$, $\mathbf{F}_{L}^{\text{refine}}$, and $\mathbf{F}_{R}^{\text{refine}}$, and constructs the refined graph $B^{(l)}=B(\tilde{B}^{(l)},\mathbf{e}^{(l)},\mathbf{f},\mathbf{F}_{L}^{\text{refine}},\mathbf{F}_{R}^{\text{refine}})$.
+Letting $v^{L}_{\max}$ and $v^{R}_{\max}$ be the maximum cluster sizes, the incidence count in $\tilde{B}^{(l)}$ is bounded by $k_{l}\leq k_{l+1}v^{L}_{\max}v^{R}_{\max}$.
+The sampling process queries a denoising model a constant number of times per step.
+The complexity is thus governed by the architecture.
+In our case, since bipartite graphs are triangle-free, the Local PPGN model [9] has linear complexity $\mathcal{O}(n+m+k)$.
+Embedding computation for $B^{(l)}$ similarly costs $\mathcal{O}(n+m+k)$.
+This includes calculating the top $K$ eigenvalues/eigenvectors of the Laplacian via the method from [42], with complexity $\mathcal{O}\left(K(n_{l+1}+m_{l+1}+k_{l+1})\right)$, and embedding via SignNet, also linear in graph size due to fixed $K$.
+The final transformation from the bipartite graph to a hypergraph—by collapsing right-side nodes into hyperedges—has a cost of $\mathcal{O}(m+k)$.
+Under these assumptions, the total complexity to generate a hypergraph $H$ with $n$ nodes, $m$ hyperedges, and $k$ incidences is $\mathcal{O}(n+m+k)$.
+Appendix H Detailed Numerical Results
+In this section, we present detailed numerical results for all datasets and metrics described in Appendix D.
+Reported values of the form $a\pm b$ indicate the mean $a$ and twice the standard deviation $b$ computed over 5 runs.
+The best and second-best results are highlighted in bold and underlined, respectively.
+H.1 Comparisons with the Baselines
+SBM Hypergraphs
+Method Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+HyperPA [29] $2.5$ $\underline{0.075}$ $4.062$ $0.407$ $0.273$ $77.840$ $0.074$ $0.008$
+VAE [30] $0.0$ $0.375$ $1.280$ $1.059$ $0.024$ $6.543$ $\mathbf{0.007}$ $0.006$
+GAN [31] $0.0$ $1.200$ $2.106$ $1.203$ $0.059$ $10.700$ $0.076$ $0.012$
+Diffusion [32] $0.0$ $0.150$ $1.717$ $1.390$ $0.031$ $13.940$ $0.040$ $0.004$
+HYGENE [10] $\underline{65.0}$ $0.525$ $\mathbf{0.321}$ $\mathbf{0.002}$ $\underline{0.010}$ $\mathbf{2.990}$ $0.016$ $\mathbf{0.000}$
+FAHNES $\mathbf{87.8\vbox{\hbox{$\bm{\pm 3.1}$}}}$ $\mathbf{0.029\vbox{\hbox{$\bm{\pm 0.009}$}}}$ $\underline{0.846\vbox{\hbox{$\pm 0.457$}}}$ $\underline{0.005\vbox{\hbox{$\pm 0.003$}}}$ $\mathbf{0.006\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $\underline{6.410\vbox{\hbox{$\pm 3.124$}}}$ $\underline{0.009\vbox{\hbox{$\pm 0.006$}}}$ $\underline{0.003\vbox{\hbox{$\pm 0.001$}}}$
+Ego Hypergraphs
+Method Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+HyperPA [29] $0.0$ $35.830$ $2.590$ $0.423$ $0.237$ $143.000$ $0.354$ $0.002$
+VAE [30] $0.0$ $47.580$ $0.803$ $1.458$ $0.133$ $38.950$ $0.558$ $0.019$
+GAN [31] $0.0$ $60.350$ $0.917$ $1.665$ $0.230$ $41.800$ $0.612$ $0.015$
+Diffusion [32] $0.0$ $\underline{4.475}$ $3.984$ $2.985$ $0.190$ $6.911$ $0.407$ $0.009$
+HYGENE [10] $\underline{90.0}$ $12.550$ $\mathbf{0.063}$ $\underline{0.220}$ $\mathbf{0.004}$ $\underline{5.790}$ $\underline{0.025}$ $\mathbf{0.000}$
+FAHNES $\mathbf{99.5\vbox{\hbox{$\bm{\pm 1.1}$}}}$ $\mathbf{0.128\vbox{\hbox{$\bm{\pm 0.171}$}}}$ $\underline{0.124\vbox{\hbox{$\pm 0.086$}}}$ $\mathbf{0.155\vbox{\hbox{$\bm{\pm 0.067}$}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.003}$}}}$ $\mathbf{2.703\vbox{\hbox{$\bm{\pm 3.468}$}}}$ $\mathbf{0.003\vbox{\hbox{$\bm{\pm 0.005}$}}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$
+Tree Hypergraphs
+Method Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+HyperPA [29] $0.0$ $2.350$ $0.315$ $0.284$ $0.159$ $5.941$ $0.477$ $0.168$
+VAE [30] $0.0$ $9.700$ $0.072$ $0.480$ $0.124$ $3.869$ $0.280$ $0.139$
+GAN [31] $0.0$ $6.000$ $0.151$ $0.469$ $0.089$ $2.198$ $0.201$ $0.124$
+Diffusion [32] $0.0$ $2.225$ $1.718$ $1.922$ $0.127$ $8.565$ $0.353$ $0.139$
+HYGENE [10] $\underline{77.5}$ $\mathbf{0.000}$ $\underline{0.059}$ $\underline{0.108}$ $\underline{0.012}$ $\underline{1.099}$ $\underline{0.041}$ $\underline{0.016}$
+FAHNES $\mathbf{89.7\vbox{\hbox{$\bm{\pm 6.0}$}}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{0.022\vbox{\hbox{$\bm{\pm 0.022}$}}}$ $\mathbf{0.030\vbox{\hbox{$\bm{\pm 0.034}$}}}$ $\mathbf{0.003\vbox{\hbox{$\bm{\pm 0.002}$}}}$ $\mathbf{0.171\vbox{\hbox{$\bm{\pm 0.106}$}}}$ $\mathbf{0.014\vbox{\hbox{$\bm{\pm 0.006}$}}}$ $\mathbf{0.014\vbox{\hbox{$\bm{\pm 0.004}$}}}$
+ModelNet Bookshelf
+Method NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+HyperPA [29] $8.025$ $7.562$ $0.044$ $\underline{0.048}$ $877.500$ $0.211$ $0.005$
+VAE [30] $47.450$ $6.190$ $1.520$ $0.190$ $113.600$ $\underline{0.145}$ $\underline{0.003}$
+GAN [31] $\mathbf{0.000}$ $397.200$ $46.300$ $0.476$ $670.100$ $0.707$ $0.007$
+Diffusion [32] $\mathbf{0.000}$ $20.360$ $2.346$ $0.079$ $264.100$ $0.239$ $0.006$
+HYGENE [10] $69.730$ $\mathbf{1.050}$ $\underline{0.034}$ $0.068$ $\mathbf{27.400}$ $0.204$ $0.004$
+FAHNES $0.135\vbox{\hbox{$\pm 0.276$}}$ $\underline{2.980\vbox{\hbox{$\pm 2.107$}}}$ $\mathbf{0.020\vbox{\hbox{$\bm{\pm 0.025}$}}}$ $\mathbf{0.024\vbox{\hbox{$\bm{\pm 0.015}$}}}$ $\underline{46.614\vbox{\hbox{$\pm 58.803$}}}$ $\mathbf{0.086\vbox{\hbox{$\bm{\pm 0.030}$}}}$ $\mathbf{0.001\vbox{\hbox{$\bm{\pm 0.001}$}}}$
+ModelNet Piano
+Method NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+HyperPA [29] $0.825$ $9.254$ $\mathbf{0.023}$ $\underline{0.067}$ $\mathbf{77.840}$ $\underline{0.236}$ $0.004$
+VAE [30] $75.350$ $8.060$ $1.686$ $0.396$ $184.300$ $0.241$ $0.003$
+GAN [31] $\mathbf{0.000}$ $409.000$ $86.380$ $0.697$ $622.200$ $0.738$ $0.005$
+Diffusion [32] $\underline{0.050}$ $20.900$ $4.192$ $0.113$ $289.300$ $0.303$ $0.004$
+HYGENE [10] $42.520$ $\underline{6.290}$ $\underline{0.027}$ $0.117$ $155.000$ $0.285$ $\mathbf{0.002}$
+[…]
+Table 6: Detailed numerical results for unfeatured hypergraphs.
+ManifoldNet Airplane
+Method ChamDist NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+Sequential $0.143$ $0.367$ $0.801$ $\mathbf{0.004}$ $\mathbf{0.007}$ $4.964$ $0.087$ $0.006$
+FAHNES $\mathbf{0.048\vbox{\hbox{$\bm{\pm 0.003}$}}}$ $\mathbf{0.017\vbox{\hbox{$\bm{\pm 0.070}$}}}$ $\mathbf{0.218\vbox{\hbox{$\bm{\pm 0.085}$}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.007}$}}}$ $0.010\vbox{\hbox{$\pm 0.003$}}$ $\mathbf{2.428\vbox{\hbox{$\bm{\pm 1.455}$}}}$ $\mathbf{0.032\vbox{\hbox{$\bm{\pm 0.007}$}}}$ $\mathbf{0.003\vbox{\hbox{$\bm{\pm 0.001}$}}}$
+ManifoldNet Bench
+Method ChamDist NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+Sequential $0.117$ $0.078$ $\mathbf{0.332}$ $0.011$ $\mathbf{0.015}$ $\mathbf{3.131}$ $\mathbf{0.035}$ $0.007$
+[…]
+Table 7: Detailed numerical results for 3D meshes.
+Airplane Point Clouds
+Method ChamDist $\downarrow$ NumDiff $\downarrow$ Wavelet $\downarrow$ Orbit $\downarrow$ Clustering $\downarrow$ Deg $\downarrow$ Spectral $\downarrow$ Ratio $\downarrow$
+DiGress [6] OOM OOM OOM OOM OOM OOM OOM OOM
+DeFoG [33] OOM OOM OOM OOM OOM OOM OOM OOM
+FAHNES $\mathbf{0.094\vbox{\hbox{$\bm{\pm 0.006}$}}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.001}$}}}$ $\mathbf{0.074\vbox{\hbox{$\bm{\pm 0.104}$}}}$ $\mathbf{0.264\vbox{\hbox{$\bm{\pm 0.254}$}}}$ $\mathbf{0.002\vbox{\hbox{$\bm{\pm 0.002}$}}}$ $\mathbf{0.005\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $\mathbf{67.311\vbox{\hbox{$\bm{\pm 44.903}$}}}$
+Bench Point Clouds
+Method ChamDist $\downarrow$ NumDiff $\downarrow$ Wavelet $\downarrow$ Orbit $\downarrow$ Clustering $\downarrow$ Deg $\downarrow$ Spectral $\downarrow$ Ratio $\downarrow$
+DiGress [6] OOM OOM OOM OOM OOM OOM OOM OOM
+DeFoG [33] OOM OOM OOM OOM OOM OOM OOM OOM
+FAHNES $\mathbf{0.130\vbox{\hbox{$\bm{\pm 0.001}$}}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{0.003\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{0.013\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $\mathbf{0.229\vbox{\hbox{$\bm{\pm 0.071}$}}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.000}$}}}$ $\mathbf{73.279\vbox{\hbox{$\bm{\pm 22.287}$}}}$
+Table 8: Detailed numerical results for graph point cloud datasets.
+H.2 Ablation Studies
+SBM Hypergraphs
+Node Budget Minibatch OT Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+$\mathbf{87.8\vbox{\hbox{$\bm{\pm 3.1}$}}}$ $\mathbf{0.029\vbox{\hbox{$\bm{\pm 0.009}$}}}$ $\mathbf{0.846\vbox{\hbox{$\bm{\pm 0.457}$}}}$ $\mathbf{0.005\vbox{\hbox{$\bm{\pm 0.003}$}}}$ $\mathbf{0.006\vbox{\hbox{$\bm{\pm 0.004}$}}}$ $\underline{6.410\vbox{\hbox{$\pm$3.124}}}$ $\mathbf{0.009\vbox{\hbox{$\bm{\pm 0.006}$}}}$ $\mathbf{0.003\vbox{\hbox{$\bm{\pm 0.001}$}}}$
+[…]
+[…]
+$84.6\vbox{\hbox{$\pm$4.6}}$ $0.049\vbox{\hbox{$\pm$0.045}}$ $916\vbox{\hbox{$\pm$0.749}} 0.0$$47\vbox{\hbox{$\pm$0.070}} 0.0$$007\vbox{\hbox{$\pm$0.007}}$ $6.665\vbox{\hbox{$\pm$5.199}}$ $012\vbox{\hbox{$\pm$0.014}} 0.0$$004\vbox{\hbox{$\pm$0.004}}$
+Ego Hypergraphs
+Node Budget Minibatch OT Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+$99.5\vbox{\hbox{$\pm$1.1}}$ $0.128\vbox{\hbox{$\pm$0.171}}$ $\underline{0.124\vbox{\hbox{$\pm$0.086}}}$ $\underline{0.155\vbox{\hbox{$\pm$0.067}}}$ $\mathbf{0.004\vbox{\hbox{$\bm{\pm 0.003}$}}}$ $\mathbf{2.703\vbox{\hbox{$\bm{\pm 3.468}$}}}$ $03\vbox{\hbox{$\pm$0.005}} 0.000\mathbf${0.000}
+$\mathbf{100.0\vbox{\hbox{$\bm{\pm 0.0}$}}}$ $0.118\vbox{\hbox{$\pm$0.158}}$ $0.140\vbox{\hbox{$\pm$0.104}}$ $\mathbf{0.133\vbox{\hbox{$\bm{\pm 0.106}$}}}$ $0.005\vbox{\hbox{$\pm$0.003}}$ $3.426\vbox{\hbox{$\pm$2.629}}$ $\mathbf{0.000}$ $\mathbf{0.000}$
+$\underline{99.9\vbox{\hbox{$\pm$0.4}}}$ $\mathbf{0.073\vbox{\hbox{$\bm{\pm 0.050}$}}}$ $0.237\vbox{\hbox{$\pm$0.465}}$ $193\vbox{\hbox{$\pm$0.175}} 0.0$$007\vbox{\hbox{$\pm$0.012}}$ $4.521\vbox{\hbox{$\pm$5.629}}$ $\mathbf{0.000\vbox{\hbox{$\bm{\pm 0.002}$}}}$ $\mathbf{0.000}$
+[…]
+Tree Hypergraphs
+Node Budget Minibatch OT Valid $\uparrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+$89.7\vbox{\hbox{$\pm$6.0}}$ $\mathbf{0.000}$ $\underline{0.022\vbox{\hbox{$\pm$0.022}}}$ $\mathbf{0.030\vbox{\hbox{$\bm{\pm 0.034}$}}}$ $\mathbf{0.003\vbox{\hbox{$\bm{\pm 0.002}$}}}$ $\underline{0.171\vbox{\hbox{$\pm$0.106}}}$ $\underline{0.014\vbox{\hbox{$\pm$0.006}}}$ $0.014\vbox{\hbox{$\pm$0.004}}$
+[…]
+[…]
+$\mathbf{95.6\vbox{\hbox{$\bm{\pm 3.7}$}}}$ $\mathbf{0.000}$ $042\vbox{\hbox{$\pm$0.042}} 0.0$$77\vbox{\hbox{$\pm$0.091}} 0.0$$005\vbox{\hbox{$\pm$0.003}}$ $217\vbox{\hbox{$\pm$0.169}} 0.0$$013\mathbf$$\vbox{\hbox{$\bm{\pm 0.013}$}}$} $\mathbf{0.008\vbox{\hbox{$\bm{\pm 0.011}$}}}$
+ModelNet Bookshelf
+Node Budget Minibatch OT NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+$\mathbf{0.135\vbox{\hbox{$\bm{\pm 0.276}$}}}$ $\underline{2.980\vbox{\hbox{$\pm$2.107}}}$ $\mathbf{0.020\vbox{\hbox{$\bm{\pm 0.025}$}}}$ $\underline{0.024\vbox{\hbox{$\pm$0.015}}}$ $\mathbf{46.614\vbox{\hbox{$\bm{\pm 58.803}$}}}$ $\mathbf{0.086\vbox{\hbox{$\bm{\pm 0.030}$}}}$ $\mathbf{0.001\vbox{\hbox{$\bm{\pm 0.001}$}}}$
+$0.940\vbox{\hbox{$\pm$0.917}}$ $3.040\vbox{\hbox{$\pm$1.446}}$ $089\vbox{\hbox{$\pm$0.099}} 0.0$$032\vbox{\hbox{$\pm$0.013}}$ $59.300\vbox{\hbox{$\pm$108.101}}$ $137\vbox{\hbox{$\pm$0.042}} 0.0$$003\vbox{\hbox{$\pm$0.001}}$
+$\underline{0.265\vbox{\hbox{$\pm$0.496}}}$ $4.400\vbox{\hbox{$\pm$1.635}}$ $\underline{0.025\vbox{\hbox{$\pm$0.021}}}$ $\mathbf{0.014\vbox{\hbox{$\bm{\pm 0.007}$}}}$ $\underline{54.558\vbox{\hbox{$\pm$30.276}}}$ $\underline{0.102\vbox{\hbox{$\pm$0.020}}}$ $\mathbf{0.001\vbox{\hbox{$\bm{\pm 0.001}$}}}$
+$1.325\vbox{\hbox{$\pm$1.631}}$ $\mathbf{2.820\vbox{\hbox{$\bm{\pm 0.958}$}}}$ $055\vbox{\hbox{$\pm$0.077}} 0.0$$031\vbox{\hbox{$\pm$0.009}}$ $79.889\vbox{\hbox{$\pm$98.721}}$ $135\vbox{\hbox{$\pm$0.012}} 0.0$$003\vbox{\hbox{$\pm$0.001}}$
+ModelNet Piano
+Node Budget Minibatch OT NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+$\mathbf{0.846\vbox{\hbox{$\bm{\pm 1.009}$}}}$ $\mathbf{3.265\vbox{\hbox{$\bm{\pm 1.954}$}}}$ $\mathbf{0.042\vbox{\hbox{$\bm{\pm 0.056}$}}}$ $0.040\vbox{\hbox{$\pm$0.026}}$ $\underline{119.158\vbox{\hbox{$\pm$81.023}}}$ $\mathbf{0.123\vbox{\hbox{$\bm{\pm 0.119}$}}}$ $\mathbf{0.002\vbox{\hbox{$\bm{\pm 0.002}$}}}$
+[…]
+[…]
+[…]
+Table 9: Detailed numerical results for ablation studies on unfeatured hypergraphs.
+ManifoldNet Bench
+Node Budget Minibatch OT ChamDist $\downarrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+[…]
+$0.090\vbox{\hbox{$\pm$0.003}}$ $0.240\vbox{\hbox{$\pm$0.265}}$ $1.089\vbox{\hbox{$\pm$0.346}}$ $013\vbox{\hbox{$\pm$0.015}} 0.0$$018\vbox{\hbox{$\pm$0.009}}$ $6.507\vbox{\hbox{$\pm$2.771}}$ $\mathbf{0.014\vbox{\hbox{$\bm{\pm 0.005}$}}}$ $\underline{0.005\vbox{\hbox{$\pm$0.002}}}$
+[…]
+$0.098\vbox{\hbox{$\pm$0.024}}$ $0.267\vbox{\hbox{$\pm$0.319}}$ $1.242\vbox{\hbox{$\pm$0.582}}$ $013\vbox{\hbox{$\pm$0.011}} 0.0$$022\vbox{\hbox{$\pm$0.014}}$ $7.619\vbox{\hbox{$\pm$4.165}}$ $\underline{0.015\vbox{\hbox{$\pm$0.016}}}$ $0.006\vbox{\hbox{$\pm$0.004}}$
+ManifoldNet Airplane
+Node Budget Minibatch OT ChamDist $\downarrow$ NumDiff $\downarrow$ Deg $\downarrow$ EdgeSize $\downarrow$ Spectral $\downarrow$ Harmonic $\downarrow$ Closeness $\downarrow$ Betweenness $\downarrow$
+[…]
+[…]
+$\underline{0.050\vbox{\hbox{$\pm$0.005}}}$ $\underline{0.052\vbox{\hbox{$\pm$0.065}}}$ $\underline{0.235\vbox{\hbox{$\pm$0.063}}}$ $\underline{0.014\vbox{\hbox{$\pm$0.027}}}$ $\underline{0.012\vbox{\hbox{$\pm$0.005}}}$ $\mathbf{1.604\vbox{\hbox{$\bm{\pm 1.400}$}}}$ $\mathbf{0.022\vbox{\hbox{$\bm{\pm 0.007}$}}}$ $\mathbf{0.003\vbox{\hbox{$\bm{\pm 0.001}$}}}$
+$0.100\vbox{\hbox{$\pm$0.023}}$ $0.304\vbox{\hbox{$\pm$0.437}}$ $864\vbox{\hbox{$\pm$0.358}} 0.0$$20\vbox{\hbox{$\pm$0.016}} 0.0$$019\vbox{\hbox{$\pm$0.009}}$ $4.184\vbox{\hbox{$\pm$2.176}}$ $\underline{0.025\vbox{\hbox{$\pm$0.009}}}$ $0.005\vbox{\hbox{$\pm$0.002}}$
+Table 10: Detailed numerical results for ablation studies on 3D meshes.
 Appendix I Comparison between Training and Generated Samples
@@
 (vii) Airplane 3D meshes.
-(viii) Generated QM9 molecules.
+Train samples
+Generated samples
+(iii) Airplane point clouds.
+Train samples
+Generated samples
+(iii) Bench point clouds.